A Complete Bibliography of Publications in ACM SIGSOFT Software Engineering Notes: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 August 2018 Version 1.00

Title word cross-reference

#5 [Sch85b]. M [CMP89]. Δ [KsKGC89]. N [BKL87]. -grammars [KsKGC89]. -version [BKL87]. 0-262-12112-3 [Con86]. 1 [MCB82, Neu89]. 14 [Neu89]. 14-1 [Neu89]. 1838A [MOPT88]. 1979 [Kil83]. 1980 [Smo81]. 1981 [Ger82, Smo84]. 1982 [Miy82]. 1983 [Neu84, Rei85, Smo83]. 1985 [Neu85, Smi85]. 2 [GC88, Jor89]. 2167 [Coa87]. 2nd [CKCM89]. 4 [Blu87]. 5th [Whe81]. 6 [Wal83]. 7th [Ano80b]. 85 [Mus85, Sch85b]. 8th [Wei85]. A1 [Ben85]. ABEL [DO85]. Abstract [Vit84, DMP82]. abstraction [BEH82, Dav85, Zah83, Con86]. abstractions [Sha89]. Abstracts [Ano84, GHW86, NTB^{+}86, Sof80, Sof81, Sof82a, Sof82b, Sof83a, Sof83b, Sof83c, Sof84a, Sof84b, Sof84c, Sof85a, Sof85b, Sof85c]. access [BK85a, Pro85]. accordian [Bro84]. ACM [RBB^{+}82, Sof85d]. acquisition [NMS89]. ACT [MCB82]. ACT/1 [MCB82]. acting [Rob89a]. Action [KSJ89, Dub89]. active [RSS88]. activities [Sto85, Yel81]. activities/positions [Yel81]. Activity [Che89]. actual [Zav84c]. Ada [Sof85d, Dun82, Boy87, CHR86, Dru82b, FWH84, Hoo85, JM83, Lad88, MN85, Ncu88, Ram89, Rey85, Weg80]. Adaptable [Ano83]. Ada(R) [ST84]. Additional [Tan81]. adequacy [Wei89]. Advanced [SBB^{+}81, Sut89, Fri83, ST84, Kil89]. advice [Sal89]. affinities [MB89]. AFFIRM [Ger81a, Mus85]. AFFIRM-85 [Mus85]. Affording [Lub86a]. AI [Di 85, Par86c, ABF88, Poh86]. AI-SM [Par86c]. aid [TMO89, TSP83]. Aided [LC84, KR89]. AIDS [Mus85]. Aladdin [Pak89]. Aladdin/lamp [Pak89]. algebraic [Ant89, Ard86, vI89]. algorithm [Neu85, Ung84]. Algorithms [Sch86b, Brä89]. all-du-paths [BS89b]. also [Cha83]. alternate [Blu82]. Alternating [DiV81]. alternative [JG86]. always [Law85b]. American [Dij81]. Amoeba [BVT89]. among [Ano80b]. Analogical [MH89]. analogy [Sal89]. Analysis [MLJ89, Ara89, AWD89, BS89a, Blu89, BH88, Bot86, CRZ88, Fai88, Gla80a, HMH80, KM89, Lei89, Pet89, Pre81, Ric86a, SM89, SCC^{+}83, TSP83, TP86, YTFB89, Ole85]. analyst [Bal83]. analyzer [OO89b]. anecdotage [Tav82]. anecdotes [Ano80b]. annotated [Coo82, RS82, WB87]. Announcement [Sno86]. answer [Poh86]. antidotes [Neu85]. anxiety [Neu85]. Application [AF88, Chr89a, ABF88, CR83, CR84, CGGS^{+}82, Fos83, TP86]. Applications [Lan81, DRAP82, KS89, LW85, WY89, Zuc89b]. Applying [BB89, IS86, KT84, Not89]. apprentice [RW89]. Approach [Leh86, BT79, BW83, BK84, CH80, CCAJ86, Fri83, JvB83, KN85, Kat89, Keu82, KSJ89, Liv88, MSK89, MH89, Mor89, Rit87, Sca89a, SM89, Sno84, Spi83, Tar88, Wil89, YES86]. Approaches [ROT89, Smo82, Coa87, Gre86, LGM^{+}86]. approved [Buc86a]. APSE [Dru82a, MOPT88]. Arcadia [TBC^{+}88]. architecture [Phi89, TBC^{+}88]. architectures [Brä89]. Arcturus [ST84]. area [MRY86]. areas [BT79]. arising [Cra85]. arrays [Den80, GMW80]. art [Ara89, Neu82, Yac82]. Article [Bin89]. artificial [AF88, Par86c]. aspect [Hit89]. aspects [Buc86a, Par85e]. assessing [CCAJ86]. Assessment [Zav88, Ger89, Ric86a, WB87]. assistance [OO83]. assistant [CY84, Dar89, HL88, PF89]. assisted [GH83]. assumptions [RW82]. assurance [Lam85, LGM^{+}86, Liv88, Loy88, Yia84]. Assuring [CS89]. ASTM [Sch86a]. Attribute [NP88, Gil81b]. attributes [Gil86a]. attribution [ACR^{+}88]. Augmented [Hen89]. Automated [Ost89, AWD89, Ben81, Bob87]. Automatic [BHO89, Bar82, Fos80, Gor89]. Automating [KM89]. automation [Huf89, Sti89a]. avenue [FB81]. avenues [Ger81c]. avionics [Lev87]. axioms [DMP82].

B [Rei85, Gel85]. B-ware [Gel85]. B. [BK83]. Babb [Ano80a]. back [Ham80b]. backtracking [AS89, Mil82]. badly [Bro84, Rob84]. Balloting [Buc88]. Barbara [Con86]. base [DHPT89, RVAR85, Rom89, Lun89]. base/OPEN [Lun89]. based [Cle88, FJM^{+}84, GM84, Gre86, HH82, HM84, HL88, Kai89a, KS85, LBC^{+}83, MM82, Nar89, NMS89, ROT89, RC89, Sca89a, STJ83, Sol84, Str88, TD85, VF83, Yue89]. bases [CCR83]. BASIC [CY84]. BASIC-PLUS [CY84]. basis [BGW82, Ger85]. battle [Hor85]. be [Ger81c, Smi85]. becoming [Tar88]. behavior [CSB82, Win85]. behavioral [Wil89]. belittled [Loy85]. benefit [Gla80a]. benefits [McF86]. Beneich [CS86]. better [MOT81]. between [Not88]. Beware [Har85]. Beyond [Di 85, Win85, RR85, Tai89]. Bibliography [AS89, Coo82, WB87]. Bit [DiV81]. bites [Lev87]. Blit [Car83b]. blocks [Jos84]. Boehm [BK83]. Book [CS86, Con86, Kil89, Gri82]. books [Ger82, Miy82, Smo81]. both [Cha83]. bottom [McF86]. bottom-up [McF86]. branch [Tai89, YM89]. breakins [Rei86]. Breton [CS86]. brick [Buc86b]. Brown [Rob84]. budget [Gan83]. BUG [Gar81, vV89]. building [Goo81a, Gor89]. business [KS89].

C [Smo81, Smo84]. CA [Smi85]. calculus [KN85]. calendar [Jan84]. Canada [Bon81]. capabilities [MLJ89]. capital [Weg82]. capital-intensive [Weg82]. Capturing [RLL89, Zuc89b]. case [DG88, Dun82, HB89, Lei89, War85, BS89a, Kar89, Zuc89a]. cases [BS89b]. Catastrophe [Pra87]. cause [AG83]. cause-effect [AG83]. CDL [LS89]. Centaur [BCD^{+}88]. Cesar [OO89b]. chaff [Gri80]. Change [LCDK89, You89a, Flo88, GGH^{+}82, LL89, Phi89, WCW89]. Change/configuration [You89a]. Characteristics [HW87]. characterization [Raj82]. characterize [CSB82]. charts [PMG85, Sha82]. checklist [Win86]. Cheheyl [Den82]. CHIPSY [EHW89]. classes [Win85]. Classics [Miy82]. Clocksin [Smo84]. clustering [New89]. CMA [PF89]. CMS [Mur88b]. Cobol [HMH80]. Code [Bin89, Lub86b, Ric86b, WPM89]. Coding [Per86a, Joh87]. cognitively [Sol84]. cognitively-based [Sol84]. Combining [Gor89, Win85]. Comment [CR84, Fos83, Fos80, Ham80a, Per80, Yac82]. Comments [Bow86, Gri82, Joh83, Joh87, Wal81]. Commercial [GMS89]. common [MOPT88, Rin87]. communication [BCK^{+}82, CH80]. communication-oriented [CH80]. Comparing [FWH84, Wei89]. Comparison [Ard86, BKL87, Ham89, Man87, RZ89, Tan80a, Tan81]. Comparisons [Sch81a, And82, IMOL83, Pec80]. COMPCON [Bal82]. compilation [EHW89, Fri83, You89b]. compiler [Fer83, OO83]. compiling [Fri84]. complex [Rit87]. complexity [CN87, Tan80a, Tan81, WB87]. Components [Wil85, LM89]. compose [Gar85]. composing [Zim88]. composition [Gar85]. computation [WY89]. Computer [KR89, LC84, Zah88b, Ano80b, BK85a, Ger89, GLR^{+}82, Lyo83, Neu82, Neu84, Neu85, Neu86a, Neu86b, Neu86c, Neu86d, Pak89, Str88, Web89, Whi86]. Computer-Aided [LC84]. computers [Neu87c]. Computing [Hor85, Neu82, Yac82, Den82]. concentration [New89]. concept [KT84, MJ82]. concepts [Bol89, IS86, Liv87]. concurrency [Chr88, Pix85, YTFB89]. Concurrent [Aoy87, AWD89, BDM^{+}83, KsKGC89, WY89, Web83b]. Conditional [DP85]. Conference [Kra87, Som87, Bal82, Bre87, Whe81]. Configuration [Fal89, KMS89, Pak89, Sim89, TS89, Win88, AB89, BVT89, Baz85a, Dar89, EGK84, GMS89, HK88, Jor89, Kai89b, KK88, MSK89, Mil89, Mor89, PF89, Tho89a, You89a]. configurations [ML88]. configured [CS89]. Conflict [MCP89]. confounding [CGGS^{+}82]. connection [Ano80b]. conservative [Hil89]. considerations [Fer89]. considered [MJ82]. consistent [BKL87, WBMN88]. constraint [OO89b]. construct [Mok86]. Constructing [FFNS89]. construction [CCR83, KMS89]. constructive [Con81]. constructs [Put85a]. Contemporary [HCI^{+}81]. context [Tul89]. continued [MB85]. contracts [Fra82]. contradictions [Gel85]. contribution [Zel84a]. Control [Mir89a, AB89, Cle89, DeT84, EGK84, Gor89, Hag87, Mur88b, Pak89, Ros81, Sch89, Wes89, Yia84, Zel83, Win88]. control-flow [Zel83]. controller [FP89]. controlling [SAP89, WCW89]. coordination [Che89]. coordinators [Mar89]. cornerstone [Gil81b]. Correction [Ano80a, Zav84a]. correctness [CS89, MCL85a, RL89]. COS [Lan81]. COS/NFE [Lan81]. cost [GGH^{+}82, Gil86b]. cost-of-change [GGH^{+}82]. costing [Mar88]. countess [JM83]. coupling [Off89]. course [Fra82]. CP [Wal83]. CP-6 [Wal83]. craft [Ger82]. Crispin [Neu84]. Criteria [IMOL83, BS89a, Ben85, Wei89]. criterion [BS89b]. critical [Loy85]. Cross [SP89]. CSCW’86 [Kra87]. CSM [Sch85b]. current [Mar88]. Cycle [Str88, Blu82, Gla82a, MJ82, MRY86, Sch85a, Zah83, vdBEF^{+}82]. cycles [Hal82].

Daniel [Kil83, Kil89]. DAPSE [Boy87]. Darwin [MB84]. Data [EMS89, Blu89, BEH82, CCR83, DHPT89, Fos84a, HS89, Mit82, PF89, RVAR85, TD85, Wei89]. Data-flow [EMS89]. database [Kil89, LCDK89, LH88, PL83, SGCZ87, Tho89a]. dataspace [RC89]. Dave [Rob84]. David [Ger82, Ham80a, Mar84, Mey83]. DBO [Gil84]. deadline [Cos84]. debate [Blu82]. debug [ST83, Bea83]. Debugger [Wal83, BDM^{+}83, Bea83, Car83a, Car83b, JK83, Kis83, Zel83, Gar84]. Debugging [Gra83, Ham83, AS89, BW83, Bat83, BH83, Car83a, CR83, CR84, Fer83, Fos83, Fri83, GH83, Gro83, Hil83, LBC^{+}83, Mül83, NLY81, OO83, PC89, PL83, STJ83, TSP83, Web83b]. decomposition [For89]. deeper [Huf89]. defence [Hal82]. defense [Par85e, Wei85]. definable [Mag84b]. defining [BT79]. definition [Di 85, DHPT89, Hum89]. Degree [BD83b]. delayed [Gar81]. Delivery [Gil85]. DELTA [Wal83]. Denis [CS86]. dependence [OO84]. dependencies [PC89]. derivation [DHPT89, MH89, Red88, Rom89]. described [Ost89]. Describing [Rob89a]. Description [GRD84, Kat89, KMS89]. descriptions [CS89]. descriptive [Bol89]. Design [Bal83, Coa87, Fer89, For89, Gil87, Lun89, AF89, Ant89, ABB^{+}82, BVT89, BS89a, Bat83, Ben86, Bot86, Bro84, CCR83, Cra85, CGGS^{+}82, DGGK87, Ear86, EMS89, GGH^{+}82, Gil84, Gil86b, Jam84, Kin82, KSJ89, LST84, LS89, Mac85, MG88, McF86, PRS89, Put85b, Raj82, Ric86a, Rin87, Rob89b, Rob84, Sch81b, SCC^{+}83, TMO89, Tri88, WPM89, Wes88, Whi85, WMAJ89, Zah88a, Har88]. Designing [Mil82, Gol87, Sol84]. designs [Sta82, Zel84b]. detailing [Koo89b]. Detecting [Fea89]. developers [Mar89]. Developing [Ram89, KT84, LH89, RS82, Str88]. Development [BDM^{+}83, CDF^{+}84, Aoy87, Ber85, Boe86, CK84, Chr88, Chr89a, DP89, DH84, Dut84, Fal89, FB81, FFNS89, FM82, FWH84, FP89, Gel85, Gil81a, Gre86, Hen85, Hen86, Hen87, Hil89, HCI^{+}81, Hua87, HL88, Kai89a, Keu82, Kok89, Lad88, Leh86, Lew88, Liv87, Loy85, Mac89, Man87, Mas85, MR88, Mur88a, Neu82, Not88, OO84, Par86a, PMG85, PRS89, Pre81, RR85, Rei84, Sal89, SHD84, Sno84, SGCZ87, SBB^{+}81, Sti89b, SB89, Tan80b, Tar88, TMO89, WBMN88, War85, Web84, Wes89, Win86, Yia84, YTFB89, vdBEF^{+}82, Con86]. development-in-the-large [RR85]. development/programming [Hen85, Hen86]. developments [JG86]. diagram [Lem86]. diagrams [Hen89]. dialogic [FFNS89]. dice [Fri84]. dictionaries [LM89]. dictionary [Neu84]. did [Bro84, Rob84]. didn’t [Nel85]. different [GMS89, Win85]. Dijkstra [SL88]. dik [IS86]. dimension [Bot86]. dimensional [Mas85]. Dining [Kok87a]. direct [IS86]. directed [HL85, New89, Zel84a]. directions [Par86a]. discipline [Ara89, Dru82a, Zim88]. disciplined [Leh86, RW82]. Discovering [Par85a, SAP89]. discussion [Gar81, Ham80c, Wei85]. display [MOT81]. displays [AL85]. disruptive [Ung84]. Distributed [Gro83, BW83, Boy87, DGGK87, DeT84, Fri84, GM85, KSJ89, LC84, Pak89, SHG89, WBMN88, Web84]. DMDOS [IS86]. do [Min86, Tai89]. Document [DGKLM84]. DOD [MOPT88, Coa87]. DOD-STD-1838A [MOPT88]. DoD-STD-2167 [Coa87]. Does [Lan85a, Dru82a]. dogmas [McL85b]. Dolkes [Rei85]. Domain [Ara89, Gre86, Rob89c, SM89]. Donald [Neu84]. Don’t [Ham80b]. dormant [Agg85]. down [Buc86b, Hil89, Zah88a]. Dr. [Gri82]. draft [BW83, Car83b, GH83, STJ83]. dream [Mil82]. driven [Fer89, RVAR85]. Duplicate [Chr89b]. dust [Lev87]. DYMOS [CL83]. dynamic [Bot86, CL83].

E. [Miy82, SL88]. each [vV89]. early [Kar89, SHD84]. Economics [BK83]. Edge [New89]. editing [HL85, Zel84a]. editor [FJM^{+}84, Zel84a]. editors [GM84, HM84]. Education [Spi83, Wer87]. effect [AG83, BCK^{+}82, Dru82b, Off89]. effects [YM89]. efficiency [Lok89, NLY81]. effort [Cra85, Kok88]. eighteen [Rid84a]. elaborations [Dub89]. Eldred [Smo81]. electronic [DdG89]. elements [Chr89b]. Eliminating [Gri80]. embedded [Tar88, Whi86]. Embellishments [Den80]. emerging [Dun85]. emphasis [Zav86b]. Empirical [Sol86]. enable [Dun85]. Enactable [Krz89]. enactment [Ost89]. encrypted [SSDG81]. enemy [Ham81]. engineer [Gil86b, Har87]. Engineering [BD83b, BK83, Cro85, LC84, LF83, PPT88, Spi83, Agr81, Ano84, ABF88, Ara89, BT79, BCE^{+}82, Bre87, Buc86a, Buc88, Cos84, Dru82b, Fai88, Flo88, Gil81b, Gil84, Goo81b, GHW86, Ham80b, Ham80c, Ham81, Hum89, Koe88, Kok87b, LCDK89, ML88, Miy85, Miy87, Moo85, Mül83, Mus83, NTB^{+}86, Pec80, Pra87, RBB^{+}82, Sim89, Sti89a, Tho89a, Wer87, Wes88, Whe81, Zah88b, Zav86a, CS86, Miy82]. engineers [Fra82]. enhanced [MSR85]. enhancement [Bal86, Boe86]. Enhancing [Pix85]. Enlarging [Loy89]. Environment [CDF^{+}84, Str88, CH80, CRZ88, Cle88, DMS84, DG88, EK85, Fri84, GF84, GM84, GBM85, GC88, Gus89, KS85, Kis83, LC84, Leh86, Lew88, LH88, MOT81, MB84, MR88, OCG^{+}81, OO84, RSS88, SHD84, Sim89, SC85, Sno84, SGCZ87, Sof85d, ST84, TBC^{+}88, War85, Web84, Weg80, Wer83, Wes89, YES86, YTFB89, Per89a]. Environments [RW86, Som87, ADE^{+}82, ARSV89, BSSS84, BCE^{+}82, Bre87, Fer89, HK88, Hen85, Hen86, Hen87, Hen88, HW87, HCI^{+}81, Lee83, LPRS88, Mag84a, Mit82, Mül83, NMS89, Not88, Par85d, PRS89, Pre81, RVAR85, Ree87, Sol84, Sti89b, WBMN88, ZAA^{+}81, Dow86b, PPT88]. EPOS [Lem86]. equational [KN85]. ergo [LPRS88, NP88]. Errata [Neu89]. error [CR83, CR84, Fos83, Fos85, Rin87]. error-sensitive [CR83, CR84, Fos83]. ERTS [Cur86]. ESML [BJKW88]. Esprit [Tho88]. essay [BK83]. Essential [LT87]. Estimating [BS89b, Gil86a]. Ethnotech [Kil83]. Ethnotechnical [Kil83]. Evaluating [BCK^{+}82]. Evaluation [ABB^{+}82, Mar88, vdBEF^{+}82, ADE^{+}82, Bow86, CRZ88, Joh83, RBB^{+}82, Rze83].

Evans [Rei85]. event [Mir89b]. EVENT-ANALYSIS [Ole85]. EVES [CS85]. Evolution [Baz85a, Cha83, Eji83, DG88, Gus89, KMS89, Leh83, MB84, Min86, Rob89b, Sca89a]. Evolutionary [Gil81a, Gil85, Hek87, Krz89]. evolutions [Fea89]. evolving [Per86b]. example [Pro85, Ros81, TMO89]. executable [Dav82, GM82, Leh82, Phi89, Smo82, Sta82, Tho89b]. Executing [BH88]. execution [CMP89, CSB82, Kin81, Mik83]. executions [BD83a, RG89]. exercise [Fos80]. exercises [LF83]. existing [MN85]. expansion [CHR86]. expenses [Zav84c]. Experience [GLR^{+}82, Hek87, Moo81, Bal83, BK84, EGK84, Fri84]. Experiences [Jor89, Zuc89b]. experiment [Ben81, Buc88, Goo81b, Kem89, WWFT88]. Experimental [RZ89, Bar82, Kis83]. Experiments [AWD89, CGGS^{+}82, RBB^{+}82]. expertation [Par87a]. Experter [Par87a]. exploitation [FB81]. Exploiting [AL85]. Exploring [Par86b, Par86a]. expression [Pix85]. expressions [BH83, Fos84a]. Extended [Lew88, Vit84, BK85a, BJKW88]. extensions [Pet89]. extraction [MB89].

F [Smo84]. FACC [Yel81]. Facilitating [WCW89]. fact [Off89]. factorial [CGGS^{+}82]. factors [SB89]. factory [Sca89b]. failure [Ben86]. fairness [CM88b, Dij88, SL88]. family [GM84]. fashioned [Ano85]. fast [Mit82]. fault [STJ83, Web89]. fault-tolerance [Web89]. FDM [Lan81, Pix85]. feasibility [Wei85]. features [FM82, SP89]. Feb [Smi85]. FGMSD [Blu87]. FGMSD-80-4 [Blu87]. fiction [Off89]. field [BT79, Clu89]. files [Agg85, Ric85]. find [vV89]. Finkel [Neu84]. first [Gar81, Ham80c, KN85, Yue89, CM88a]. first-order [KN85]. fiscal [Zav84c]. five [OCG^{+}81]. flight [Gar81]. flow [Den80, EMS89, GMW80, GLR^{+}82, HS89, TD85, Zel83]. FM89 [Ger89]. focus [Red88]. foolish [Par85c]. form [Ara89]. Formal [Ano80c, Hag87, Web89, BHO89, Ber85, EK85, Ger89, LH89, PRS89, Pro85, Rom89, SHD84, Sto85, TMO89]. Formalisms [Woo88, Kel89]. formalization [AG83]. formalizing [Lan85b]. Formally [DdG89, Ram89]. forms [KK88]. FORTRAN [Jan84]. Foster [CR84, Pol80]. Found [Zav86a]. foundation [SCC^{+}83]. Foundational [EP86]. Foundations [TBC^{+}88, Kem85]. four [Man87]. fourth [Bot86, Cha84]. fourth-generation [Cha84]. fractional [CGGS^{+}82]. framework [ABF88, FFNS89, Koe88, PRS89, Rom89, Str88, Vit84]. Freedman [Kil83]. friendly [Gol80]. full [Bro84, Not89]. Functional [Gar85, Tho89b, BK85b, Kat89, Mac89, Win85]. functions [HW87]. Future [Liv87, Sof85d, Par86a].

G [Den82]. gaining [Ham80b]. GAO [Blu87]. Gasser [Den82]. gathering [You89a]. general [GMMP89]. Generalized [BH83, KT84]. generated [RLL89]. Generation [Ung84, BHO89, Cha84, GBM85, HB89, MM82, WPM89, Bin89]. generator [RT84]. Geoffrey [Neu84]. gepreek [Par81]. Gerald [Kil83]. Gerhart [Gri82]. get [Gla82a]. Glas [Per80]. Glass [Gla81]. GNOME [GM84]. goal [Dub89]. goal-oriented [Dub89]. goals [Rob89c]. good [GJ84]. Goodfellow [Neu84]. goose [Bon81]. governed [MR88]. grammars [KsKGC89]. Graph [ACR^{+}88, HK88, OO84]. graphic [KS89]. Graphical [Jon89, Rei84, SM86, Tri88, DGGK87, Miy87, Rob89b, Uni87]. graphs [AG83, New89, Tan80b, Yue89]. Gries [Ger82, Ham80a, Mar84, Mey83]. ground [Clu89]. group [Baz85b, Gal87, GY87, Mus83, You88]. groupware [Mar89]. guide [Neu84]. guidebook [Bow86]. guidelines [Mar89].

Guttag [Con86]. Guy [Neu84]. Gypsy [CSG^{+}88, SC85, You89b].

H [Den82]. hacker [Neu84]. Hall [Ger82]. handbook [Kil83]. Hardware [GH83]. harmful [MJ82]. Harper [Neu84]. hazardous [Cha83]. HDM [MSR85]. health [Cha83]. heard [Gar81]. held [Smi85]. help [Lan85a]. Hereford [Rat87]. Hermann [Smo81]. heuristics [WO85]. HIBOL [Mit82]. hierarchical [Iiv87, Kat89, Mus85]. High [Fer83, Ger81b, OO83, ZAA^{+}81, BW83, BH83, Car83a, GH83, Per80, Ung84, Zel83]. High-level [Ger81b, OO83, BW83, Car83a, Per80, Zel83]. higher [BK85b, Lub86a, Sha89]. higher-level [Sha89]. higher-order [BK85b]. highlevel [Gar84]. Highlights [Rit87]. Hill [Con86]. history [War85]. Hoare [vH85]. Holland [Smo81]. Holt [CS86]. Horning [Lee85]. hospitals [GLR^{+}82]. Host [Tar88, KT84]. Host/target [Tar88, KT84]. Huff [Den82]. Human [SB89, BCE^{+}82]. Hyperbook [Kem89]. hypercube [CMP89].

iconic [BL88]. ICSE [Wei85]. identification [HH82]. identifying [Fos89]. Ideographic [GRD84]. IDEOSY [GRD84]. IDL [Sno86]. IEEE [Bal82, Sch85b]. If [Ham80c]. II [KA83, MRY86, Neu81]. III [Smi85, BT85, Par86a]. Illustrating [RB89, SA88]. image [Cha83]. impact [Gom83, NLY81, WCW89]. Implementation [FM82, JK83, KA83, Web83a, Whi85, vdBEF^{+}82]. Implementing [Lin84, RC89]. implications [Neu82, PC89]. importance [Gla80b, Per80]. improving [HM84]. Ina [Ber85]. INATEST [EK85]. income [Zav84c]. Incremental [Sch89, ABF88, Fri83, Fri84]. individual [Gla80b, Per80]. industrial [Agr81, Koe88]. industrial-strength [Koe88]. industry [Sto85]. infeasible [YM89]. information [BH82, Den80, GMW80, GLR^{+}82, Man87, MCB82, Rom89, WS82, WMAJ89]. Initial [TS82, RW89]. initiative [Wei85]. Inscape [Per89a]. inside/outside [Tar81]. insights [Pec80]. instances [Chr88, Chr89b]. Institute [Fai88]. Instruments [Hay81, DdG89]. Integrated [Lee83, SA88, YTFB89, GC88, Gus89, Koe88, LPRS88, Leh86, ML88, RSS88, Wer83, Wes89, Whi85]. Integrating [RAO89, Rob89c]. integration [BL88]. intelligence [AF88, Par86c]. intelligent [DeT84, HL88, MOT81, Par85d]. intensive [Weg82]. inter [BCK^{+}82]. inter-programmer [BCK^{+}82]. Interacting [RSS88]. interaction [BL88]. interactions [For89]. Interactive [BSSS84, CDF^{+}84, GRD84, Mik83, ST83, Web83b, Bea83, EK85, WS82, Zel83]. Interchanged [Neu89]. Interface [Koo89a, BL88, Mag84b, MOPT88, SSDG81]. interference [Fea89, RB89]. interfering [RB89]. international [Whe81, Dow86b]. Interprocedual [HS89]. interval [RG89]. Introduction [SHD84, Bot86, Coo82, Lyo83, Neu81]. Introductory [Wal81, GM84]. inverted [MSK89]. Investigation [BCE^{+}82, Whi86]. Invitational [LGM^{+}86]. IPSE [Hit89]. IPSEs [LT87]. ISBN [Con86]. ISDE [CDF^{+}84]. ISEF [Koe88]. isolated [RR85]. issue [Neu89]. Issues [PPT88, CGGS^{+}82, JK83, KA83, Web83a, Whi86, WMAJ89]. ISTAR [Leh86]. Italy [BD83b]. iteratively [Gil86b].

J [Bin89, Den82]. Jackson [Bal83, Kin82, PMG85, Sha82]. James [Rei85]. Jan [Sch85b]. January [Bin89]. Jim [Lee85]. job [Cle89]. job-control [Cle89]. John [Con86, Ger82, Neu85, Rei85]. join [Kok88]. Jo(R) [Ber85]. JSD [Kok87a, MG88].

K. [Pol80]. Kernel [Uni87, GC88]. Knowledge [LBC^{+}83, STJ83, Wes88, Cle88, Gre86, NMS89, Sca89a]. Knowledge-based [LBC^{+}83, STJ83, Cle88, Gre86, NMS89, Sca89a]. Kopetz [Smo81]. KVM [Lan81].

L [Gla81, Neu84, Per80]. Laboratory [LF83]. lamp [Pak89]. Language [Bab88, BDM^{+}83, BK85b, Brä89, BJKW88, Cra85, DRAP82, EMS89, Fer83, FJM^{+}84, FM82, GM85, GM82, HM84, KS89, LS89, Lok89, Mit82, RC89, Weg80, YES86, ZAA^{+}81]. language-based [FJM^{+}84, HM84]. languages [Cha84, Gio80, Gum82, Lyo83, SHG89, Sol84]. languages/environments/methodologies [Sol84]. large [Dun84, Fal89, Hek87, HN89, Lew88, Lub86b, RR85, WMAJ89, Win85]. Larger [Sha89]. Lastig [Par81]. late [Law85b]. law [MR88]. law-governed [MR88]. LCD [LH89]. LCD-reification [LH89]. lead [Dru82a]. Leaves [Bal83]. LEGO [Cur86]. Lessons [Rei86, Lan85b]. Let [Zav86b]. Letter [Cur86, Gla81, Gri82, Kil81, Kim81, Lee85, Mey83, Pol80]. level [BW83, BH83, Car83a, Fer83, GH83, Ger81b, Ham83, JK83, OO83, Per80, Sha89, VF83, ZAA^{+}81, Zel83]. Levels [Zah83, BCK^{+}82]. lexical [MB89]. Life [MJ82, Str88, Blu82, Gla82a, Hal82, LF83, MRY86, Sch85a, Zah83, vdBEF^{+}82]. life-cycle [Gla82a, Sch85a]. Lifecycle [SA85, Not89]. LIFESPAN [Pir86]. like [Gio80, Nel85, Yue89]. limited [Kis83]. Limits [Koo89b, Huf89]. Lipow [Smo81]. Liskov [Con86]. Lisp [Wer83]. Lispedit [Mik83]. list [Mir89b]. Literary [Smo83]. localization [STJ83]. Logic [KS82, Ano85, Dub89, Fos84a, Koo89a, Per89a, RG89, Yue89]. logic-based [Yue89]. Logical [YES86]. logics [Gum82]. look [Ham80b, LST84, NMS89, Par83a]. Looking [Mir89b]. Lovelace [JM83]. LUCID [GM85].

M [Den82, Kil83, Neu84]. machine [Ard86]. Machines [Lyo83, Smo83]. magic [Rid84a]. magnetic [Dun85]. magnetic-tape [Dun85]. Maintaining [Gus89]. Maintenance [CM88a, CKCM89, CH80, Cha83, Cha84, Eji83, Leh83, Par84, Par85a, Par85b, Par85c, Par86b, Par86a, Par86c, Par87b, PC89]. make [Lem86]. Management [Agg85, PPT88, Spi83, Baz85a, BH82, CSG^{+}88, Fal89, GMS89, HK88, Hor85, Jor89, KK88, LH88, Liv87, Lun89, MCP89, MSK89, Mil89, Mor89, Pak89, PF89, Rei89, Sim89, Tho89a, TS89, WWFT88, You89a, Rei85]. manager [BVT89, Ber85, FP89]. manager/controller [FP89]. managing [Ric85]. manhandled [Rei82]. manipulation [IS86]. Mappings [Fea82]. Mark [Neu84]. Martin [Kil89]. Master [BD83b]. matrix [BK85a]. maturation [Rid84a]. may [Ger81c]. MCC [BR85]. McGraw-Hill [Con86]. means [DMP82]. Measurements [Ske86]. mechanical [DiV81]. mechanism [BH83, Gar85]. mechanization [Koo89b]. meeting [Sch85b]. meets [Gal87]. Mellish [Smo84]. membership [Zav84c]. mentor [DGKLM84]. merging [Fea89]. Message [Sch81c, HLC82, RC89]. message-based [RC89]. metamodel [Wil86]. Metamodels [Kok89]. metaphor [Ham80b, Nel85]. metaphors [Nel85]. Method [Bin89, Ber85, DS89, LH89, MG88, New89, RVAR85, Sto85, WPM89]. methodologies [ABB^{+}82, FWH84, Gla80a, Man87, Par83a, Sol84]. Methodology [Bal83, Hen88, Dut84, Gra83, Kin82, LST84, MRY86, Paz86, Put85a, RBB^{+}82, Sch86a, Sol84, Whi85, Yel81, vdBEF^{+}82]. methods [Ard86, Bol89, Ger89, Ham89, Per86a, Rin87, SB89, Win85, Zav86b]. metric [Leh82]. metrics [And82, Coo82, CN87, GGH^{+}82, Kok88, Sch81a]. Michael [Rei85]. micro [You89b]. micro-Gypsy [You89b]. microcode [MCL85a]. microprocessor [Hil83]. migratability [Ske86]. military [HLC82]. Millen [Den82]. Miller [Kim81]. Minimum [Gla82b]. minus [Rid84a]. mistakes [Mey83]. MITRE [Mil81]. mixup’d [Ham81]. mixware [Ham80b, Ham80c]. mockups [Gre84]. model [BL88, BK85a, BB89, Boe86, Chr89b, Gil85, HK88, Hit89, IMOL83, Iiv87, JG86, Lan85b, Ncu88, PF89, PL83, Sch81b, WBMN88]. Modeling [Kai89b, BJKW88]. Modelling [Sca89a, ARSV89, Kai89a, Kel89, Wil89]. Models [Chr88, Sta82, Bab88, Blu82, DP89, Huf89, Krz89, Leh82, Not89, Per89b, Phi89, Rob89a, Sch86b, Sut89, Tan80a, Tan81, Tul89, Wei82, Whi86]. modification [CL83]. Modula [GC88, Jor89]. Modula-2 [GC88, Jor89]. modular [EGK84]. modularity [DGKLM84]. Module [HB89, DMP82, Lev87]. module-testing [Lev87]. monitor [Rey85]. Monitoring [Sno84, Tan80b]. monolithic [Web84]. months [Bal83]. mortem [Fai88]. much [Lan85a]. Muller [Bin89]. Multi [VF83, MOT81]. multi-display [MOT81]. Multi-level [VF83]. Multilingual [Car83a, Bea83]. Multiparty [FF89, For89]. multiple [Rob89c]. multiprocessor [RC89]. must [Gil86b]. mutants [CMP89]. My [Bro84, Bal83]. Myers [AG83]. Myron [Smo81]. myths [Tra88].

N [Miy82]. name [Par84]. named [Smo81]. natural [YES86]. nature [Tul89]. need [Dru82a, Whi85, Zim88]. needs [GMS89]. Nelson [Smo81, Smo83]. net [DS89, Yue89]. Netmail [Rei82]. Nets [Mir89a, GMMP89, TP86]. network [Ros81]. Neumann [Mok86]. never [Nel85]. news [Par85b]. next [OCG^{+}81]. NFE [Lan81]. nomenclature [Pat83]. non [Ung84, War85]. non-disruptive [Ung84]. non-UNIX [War85]. North [Smo81].

North-Holland [Smo81]. notation

[Rob89b]. notations [Jon89, Tri88]. note [Den80, GMW80]. notebook [Bal83]. notes [Par87b]. NRL [LGM^{+}86]. nuenen [Par81]. number [BS89b, Rid84a].

OBJ [GM82]. Object [AB89, DHPT89, Jam84, Lad88, LH88, MG88, Rin87, SM89, WPM89, WY89, WWFT88, Ole85, PPT88]. Object-oriented [AB89, Lad88, MG88, Rin87, SM89, WPM89, WY89]. OBJECT/EVENT [Ole85]. OBJECT/EVENT-ANALYSIS [Ole85]. objectives [Gil84, Gil87]. objects [Gus89, RLL89]. observation [BD83a]. observations [Dun84, Gro86, JW89, SHD84, Tul89]. obsolete [Tar88]. off [Gla82a]. often [Loy85]. Old [Ano85]. Old-fashioned [Ano85]. one [Buc86b]. ongoing [Kem81]. OOSD [Bin89]. open [Mey83, Lun89]. operating [Ano80b, IS86]. Operational [BGW82, LS89]. optimized [Zel83]. optimizing [OO83]. orbital [Gar81]. order [BK85b, KN85, Yue89]. oriented [AB89, CH80, Dub89, Gar84, Jam84, Lad88, LCDK89, MG88, Rin87, SM89, WPM89, WY89]. Orthogonal [Rei89]. other [Neu85]. othra [CMP89]. our [Cha83]. Overview [Coo85, CG85, Kem85, LW85, BGMT88, HW87, MOPT88, SZB82a, SZB82b, Zav84b].

P [Bin89, Kil83, CMP89]. packet [SSDG81]. PACT [Sim89, Tho88]. Page [Neu89]. pages [Con86, Neu89]. PAISLey [Zav84b]. Pamela [Rei85]. panel [Wei85]. paper [Ano80a, Bal86, Bro84, Dij88, Dow86b, Min86, Ole85, Per86b, SL88, Smi85, Tay85]. parable [MB85]. paradigm [ACR^{+}88, Flo88, Koo89b, Loy88, Mac89]. paradox [Blu85]. parallel [Brä89, SM86]. part [Mil82]. PARTIAL [Rey85]. partition [JW89]. partitioning [Loy85]. PASCAL [Gar84, KT84, Gio80].

Pascal-like [Gio80]. passing [Bal89]. path [BH83, Dru82a]. paths [BS89b, YM89]. PCTE [BGMT88, Kem89, Tho88]. PDS [KK82]. PECAN [Rei84]. penny [Par85c]. performance [Ung84]. perils [Zah88a]. persistent [WWFT88]. person [Buc87]. perspective [AF89, Jam84, WMAJ89, Zav86b]. Pete [Sch80a]. Peter [Sch80b]. Petri [GMMP89, Mir89a, TP86, Yue89]. Petri-net-like [Yue89]. Pgraphite [WWFT88]. phase [Bal89, Bro84, Dut84, Loy85, Rob84]. Philip [Kim81]. philosophers [Kok87a]. Phoan [DeT84]. Piazza [Rei85]. pick [Par83a]. Pircher [Bin89]. place [SB89]. plan [FWH84, Gel85, HL88, Paz86]. plan-based [HL88]. Planning [WF87, AF89, Mus83]. plateau [Mar84]. plea [Gol80]. please [Nel85]. plight [Dij81]. pluralistic [NMS89]. plus [Rid84a, CY84]. PML [Rob89a]. Poe [FJM^{+}84]. poetry [Zav86a]. points [Gil86a]. portable [BD83a]. Position [Dij88, Dow86b, Smi85, Bal86, Min86, Per86b, SL88, Tay85]. positions [Yel81]. possibility [Lem86]. Post [Fai88]. Post-mortem [Fai88]. potential [Dru82b]. power [Jan84, Mac82b]. practical [Cle88, KWD81, Liv88, Mor89, WA85]. practice [Joh87]. practitioner [Bol89, Gro86]. Pragmatic [DH84]. Preconditions [Ric86b]. predicate [KN85]. predication [Ncu88]. Predicting [GGH^{+}82]. prediction [Tan80a, Tan81]. Preliminary [EGK84, Fri84, Ger89, BW83, Ben81, Car83b, CHR86, GH83, MRY86, RZ89, STJ83]. premature [Not89]. Prentice [Ger82]. Prentice-Hall [Ger82]. presentation [DO85, Man87]. presented [Smi85]. Press [Con86, Kil89, Miy82]. Press/McGraw [Con86]. Press/McGraw-Hill [Con86]. pressure [Cos84]. previous [Neu89]. Principles [Mag84a, Ano80b, Gol87, TD85, Rei85]. printing [Jan84]. PRISMA [NMS89]. prize [BM83]. Probing [Huf89]. Problem [KK88, Mur88a, AF89, BKL87, Gar81, Ham80c, KS82, Mar84, Pat83]. Problems [Per89b, DH84, EP86, Gio80, Mac85]. Proceedings [Sof85d]. Process [ARSV89, Bal89, Dow86b, OO89a, Rid84b, Blu86, BB89, BEH82, Chr89b, Dow86a, EP86, Fer89, FP89, Hag87, Hit89, Huf89, Hum89, Iiv87, Kai89a, Kat89, Kel89, LL89, Leh82, Leh86, Leh89, Mok86, Not88, Not89, Per89b, Phi89, RAO89, Rob89a, Sch85a, Sut89, Tul89, Wil86, WCW89, Wil89, Yia84]. process-driven [Fer89]. processes [Krz89, Ost89, RR89, Rom89, RLL89]. processing [DRAP82, Mit82, vI89]. processor [Pro85]. production [Cro85]. productive [Rei85]. Productivity [Baz85b, Bob87, BCK^{+}82]. products [HN89]. professor [Mey83]. Program [Bal83, Bal86, BM83, Con86, GRD84, Hay81, Kin81, LvH81, Ano83, BR85, CCR83, CY84, DH84, Fos80, Gor89, Jan84, Jon89, KM89, Lan85a, McL85b, Mik83, Nar89, OO84, Par85c, PC89, Rei84, Rob89b, Rze83, Sch81b, Tan80a, Tan81, Tho89b, Tri88]. Programmable [WO85]. programmer [BCK^{+}82, Bal83]. programmer/analyst [Bal83]. programmers [MB85, Sol86]. Programming [Jos84, Pec80, BCK^{+}82, Bal89, Bar82, BSSS84, BB89, DMS84, Dij81, Dru82a, EMS89, FM82, GF84, GM84, Ger82, Hen85, Hen86, Kis83, Leh82, Leh89, Lew88, Lok89, MOT81, NLY81, OO89a, Par85d, Per80, Per86a, RR85, RAO89, Sha82, SHG89, ST84, Tho89b, Wer83, ZAA^{+}81, Ger82, Ger82, Smo84]. programming-in-the-small [RR85]. programs [Che89, EGK84, Gar85, GJ84, HMH80, Hoo85, How89, Kok87b, LH89, Lin84, Lyo83, OM89, RS82, Ram89, RG89, Red88, RB89, Rey85, SM86, TSP83, Tul89, Web83b, Zel83]. progress [ABF88, Lem86, Mus83]. Project [Spi83, Bol89, CK84, Chr89a, DO85, FJM^{+}84, HLC82, Hek87, Leh86, Lem86, LH88, Mag84b, Sca89b, Zav84b, WA85]. projection [SA88]. projects [Dun84, Kem81, LF83, SJ82]. Prolog [Ger85, LH89, Ncu88, OO89a, Smo84]. proof [Ber85, MSS82]. propagation [Per89a]. Properties [Woo89, LT87, Win85]. Proposal [ADE^{+}82, Leh82]. Proposals [RBB^{+}82]. proposed [AF89, MOPT88]. Protocol [DiV81, Phi89]. protocols [Ros81]. Prototype [CN87, BK85b, RS82, ST84, Zel84b]. prototypes [CCAJ86, FM82, Gre84, HLC82, MM82]. Prototyping [Dun82, Keu82, Pat83, SZB82a, SZB82b, SJ82, BGW82, Bar82, BH82, Blu83, BK84, DGGK87, Dav82, DMP82, DRAP82, Fea82, Geh82, GLR^{+}82, GM82, Gom83, Har87, Hek87, HH82, KK82, KR89, LPRS88, Mac82a, MCB82, Mit82, Miy87, TD85, TS82, WS82, Wei82, WCW89]. prover [BK85b, Owr85]. provers [WO85]. Proving [Coo85, KN85, KS85, LGM^{+}86]. Psychosocial [Neu82]. public [Cha83, Neu86a, Neu86b, Neu86c, Neu86d, Neu87c, Neu87a, Neu87b, Neu88b, Neu88a, NC88]. Publishers [Neu84]. Punchcard [Wil88]. pursuing [Ger81c]. put [GMMP89, Zav86b].

Quality [Yia84, And82, Bow86, Buc86a, Buc86b, Cro85, Gil86b, Lam85, Liv88, Loy88, Per86a, Sch81a]. quantitative [Krz89]. Questionnaire [Sof85e]. questions [vV89]. quick [HLC82].

R [Bin89, Neu84]. R. [Per80]. RADC [Rze83]. Raddle87 [For89]. Radical [Hil89]. Raphael [Neu84]. Rapid [Bar82, BH82, Dav82, DMP82, GM82, KK82, WS82, BGW82, Blu83, Fea82, Gom83, MM82, TD85, TS82, Wei82, SZB82a, SZB82b]. ratios [CHR86]. RC2000 [Law85a]. reactive [GGW89, Hen89]. readings [Par85a]. Real [LF83, RG89, BH88, Hil83, JM83, LS89, Mac85, Nel85, Pet89, SHG89, Zuc89b]. Real-life [LF83]. Real-time [RG89, BH88, Hil83, SHG89, Zuc89b]. realistic [LM89]. reality [Mil82]. really [Par84]. realtime [GMS89]. reasonable [Keu82]. Reasoning [SHG89, vH85, RG89]. reclamation [Ung84]. Recommended [Gla82b]. redoing [Vit84]. Reducing [YM89]. reduction [Kin81]. references [SP89, Tan81]. refinement [DH84, Rey85]. Reflections [Weg82]. Reflective [WY89]. reification [LH89]. related [Neu87c, Neu88b]. Relating [Ear86]. relation [Web89]. Relational [CCR83, Saa85, Lin84, RVAR85, Sno84]. relationship [Not88]. relevant [Kem81]. reliability [Gla80a, IMOL83, Lub86a, Smo81, Smo81]. Reliable [Ano83, GGW89]. Replacing [Cle89]. reply [CR84, Fos84b]. Report [CKCM89, Rid84b, Zav84c, Blu87, KK88, Kra87, Mus83, PPT88, RW86, RZ89, Rze83]. Reportage [Bal82]. reporting [Mur88a]. reports [NTB^{+}86]. Representation [Kel89, Nar89, vI89]. Representing [Yue89]. request [LL89]. require [Sha89]. required [BS89b]. Requirements [Bat83, Dut84, Kar89, Win86, Cro85, Dav82, Dub89, Ear86, Gom83, Hag87, HH82, KK82, KS89, MB89, MLJ89, Pix85, RW89, Rom89, ST83, Whi86, Zuc89b]. Requirements/design [Bat83]. Research [Ger81c, Leh82, EP86, OCG^{+}81, Par86a]. reservations [Leh89]. Resource [Dav85]. Response [Kim81, Lee85, Pol80, Rob84]. responses [Sof85e]. restricted [Pro85]. Retrospective [Neu81]. reusability [Lub86a, Lub86b]. reusable [DS89, LM89]. Reuse [Bol89, Kar89, RR89, Ric86b, Tra88, Rat87]. reusing [Mus85]. Review [CS86, Den82, Kil83, Neu84, Neu85, Ree87, Rei85, Smo83, Smo84, Som87, Bin89, BK83, Buc86a, Con86, Ger82, Kil89, Miy82, Smo81]. Reviews [Miy82, Gri82]. Revision [Fos85, Wes89, MOPT88]. revisited [Blu87, HL85]. Revolution [Miy82]. Rewriting [Der85, DP85, KS85]. Reynolds [Ger82]. Richard [Neu84]. rigorously [Ost89]. Rinehart [CS86]. ring [Dun85]. Risks [Neu86a, Neu86b, Neu86c, Neu86d, Neu87c, Neu87a, Neu87b, Neu88b, Neu88a, NC88]. road [Buc86b]. Robert [Gla81]. role [Baz85b, Ben86, Gre86, Har87, Leh82, Mül83]. roles [Buc87]. Rooting [GP89]. ’round [Gar81]. Row [Neu84]. RRL [KS85]. RT [LS89]. RT-CDL [LS89]. Rule [Kai89a, Fos85]. Rule-based [Kai89a].

S [Neu84, Smo84]. Sachertorte [Neu85]. safe [Gol87]. Safety [Mac82b, Lev82]. SAGA [CK84]. satisfy [BS89b]. Scale [Wei82, Per89b, Sha89, SJ82]. scary [Lev87]. Scavenging [Ung84]. Scenario [HH82, RW89]. Scenario-based [HH82]. scheduling [CMP89]. schema [Mil89]. schemes [RS82]. Schorer [Sch80b]. science [Ger82, Lyo83, Zah88b]. sciences [Bar82]. scope [Hum89, Loy89]. Screen [Gar84]. Screen-oriented [Gar84]. scripts [BHO89]. SDE [Sti89a]. SDI [Moo85, Poh86, Tic86]. SDVS [MCL85a]. SE [Clu89, Nel85]. search [Buc86b, Ham81]. second [GBM85, Ham80c]. secretary [Sch81c]. secretary/treasurer [Sch81c]. secure [Goo81a, HLC82]. security [BK85a, Kem81, Lan85b, Web89, Den82]. security-relevant [Kem81]. selected [Par85a]. Selecting [Zuc89a]. Selection [BS89a, MRY86]. semantic [Mor89]. semantically [Cha83, Eji83]. semantics [LS89, Saa85]. SEN [Bin89, Neu89]. Sensitive [Fos84a, CR83, CR84, Fos83, Fos85]. sentences [vH85]. Separate [EHW89]. Separating [Put85a]. September [Den82]. sequencing [OO89b]. Sermon [Ham80a]. SESAW [BT85]. session [Bat83, Gra83, Gro83, LBC^{+}83, Web83a]. set [Con81, LPRS88, MOPT88]. sets [LM89]. several [Zuc89b]. shared [Ric85, RC89]. SHD [Gar84]. shift [AF89]. Shore [Neu85]. Shorer [Sch80a]. Short [GJ84, Jan84, Ole85]. should [Ger81c, Min86]. shuttle [Gar81]. SIFT [MSS82]. significance [Dun85]. significantly [Eji83]. SIGSOFT [Gan83, RBB^{+}82, Zav84c]. SIMOB [BD83a]. simplifying [RW82]. simplistic [Lyo83]. SIMULA [BD83a]. single [DMS84]. sixteen [Bal83]. sketching [MM82]. skills [Per80]. SLCSE [Str88]. SM [Par86c]. small [Jos84, Lub86b, RR85, SJ82, Zel84a]. social [Dun85]. Softool [Gal87, GY87, You88]. Software [Agr81, BD83b, Bre87, CM88a, Cha84, CDF^{+}84, Coo82, CKCM89, Cos84, DP89, Dow86b, Gil84, Gil86b, Gla82b, Har88, Hen85, Hen86, Hen87, Kok88, LC84, Lev82, Miy85, Moo85, OCG^{+}81, Par85b, Par85c, Par87b, Par85e, PPT88, Rat87, RR89, Sal89, Sch86a, Spi83, Str88, Tan80b, Tra88, Tul89, WB87, WMAJ89, Win88, Ano84, Aoy87, ABF88, ARSV89, AWD89, BT79, BL88, BCE^{+}82, BR85, Ben81, Blu85, Blu86, Boe86, Bol89, Bow86, BKL87, Buc86a, Buc86b, Buc88, CK84, CH80, CS89, Chr88, Chr89b, CCAJ86, Cle88, CN87, CGGS^{+}82, DS89, Dow86a, Dru82b, Dun84, Dut84, Ear86, EP86, Fai88, Fer89, FFNS89, Flo88, Fra82, FWH84, FP89, Gar81, Gel85, GMS89, Gil81b, Gil86a, Gla80a, Gol80, Goo81b, Gre86, Gus89, GHW86, Ham80b, Ham80c, Ham81]. software [Har87, Hek87, Hen89, HW87, HN89, HCI^{+}81, Hua87, HL88, Hum89, IMOL83, Iiv87, JG86, Kai89a, Kar89, Kat89, Kel89, KK88, Koe88, Kok87b, KM89, Krz89, Lam85, Law85a, Law85b, LF83, LST84, Leh89, Lew88, LCDK89, LH88, Liv87, Liv88, Lub86a, Mac89, Mag84a, ML88, Mar88, MRY86, Mil82, MB84, Min86, MR88, Miy82, Miy87, Mül83, Mus83, Neu82, NTB^{+}86, NMS89, Not88, Not89, Ost89, OO84, Par83a, Par84, Par85a, Par86b, Par86a, Par86c, Per86b, PC89, Poh86, Pra87, Pre81, Raj82, Rei85, RBB^{+}82, Rid84a, Rom89, RLL89, Sca89a, Sch85a, SAP89, SHD84, Sim89, Sno84, SGCZ87, Sof80, Sof81, Sof82a, Sof82b, Sof83a, Sof83b, Sof83c, Sof84a, Sof84b, Sof84c, Sof85a, Sof85b, Sof85c, Sti89b, SJ82, Sut89, Tho89a, Tic86, WF87, WBMN88, War85, Web84]. software [Weg82, Wei85, Wer87, Wes89, Wes88, Whe81, Wil86, Wil89, Yia84, YTFB89, Zah88b, Zav86a, Zim88, vdBEF^{+}82, Ano83, BK83, Gro86, Rid84b, RW86, Smo81, CS86, Smo81]. software-evolution [MB84]. solving [KS82]. Some [Ano80b, Cra85, Gio80, Gol87, JW89, Joh87, Lan85b, Leh89, Mey83, Sto85, Gil86a, Gla80a, Mac85]. something [Ham80b]. Sons [Rei85]. sound [Eji83]. Source [Mur88b, JK83, MCP89]. Southern [BD83b]. SP [Mur88a, Mur88b]. Specification [Con86, Har88, Mac82a, ACR^{+}88, AF89, Ard86, BGW82, BEH82, Brä89, DO85, DS89, Fea89, FF89, Gil81b, Gil86b, GM85, Gog81, GM82, Gum82, Hag87, KK82, KS89, Koo89a, KSJ89, McL85b, MSR85, MH89, Mus85, NMS89, Put85a, ROT89, Rom89, TMO89, Web89]. specification-based [ROT89]. specifications [Ant89, BHO89, BH88, Dav82, DMP82, EK85, How89, JvB83, Rob89c, RL89, Smo82, Tho89b, VF83, Woo89, Yue89]. specified [MLJ89]. specify [Win85]. Specifying [KsKGC89, LM89, Mar89, DdG89, Gom83]. Sperry [Lev87]. spiral [Boe86, Iiv87, Spi83]. sponsored [RBB^{+}82]. Spreadsheet [Kok87b]. Spring [Bal82]. Springer [Ger82, Smo81, Smo84]. Springer-Verlag [Ger82, Smo81, Smo84]. SQA [Buc87, Loy89]. SREM [Rze83]. SRI [LN81]. Stallman [Neu84]. stand [Nel85]. Standard [Gla82b, Sch85a]. standards [Buc86a, Buc88]. Stanford [LvH81, Rei86]. STARS [Ano83, MRY86]. State [Phi89, Ard86, Hen89]. state-machine [Ard86]. state-transition [Hen89]. statecharts [MLJ89]. Static [TSP83, HMH80, OO89b]. Status [Boy87, Kem81, Rze83, SC85]. STD [Coa87, MOPT88]. Steele [Neu84]. step [DH84, Ham83, Koo89b]. step-wise [DH84, Ham83]. stepwise [Rey85]. Still [Blu83, Sch80b]. Stimulating [Mus83]. Stop [Gla82a]. storage [Ung84]. story [Law85a]. straitjacket [Mok86]. strange [Jan84]. Strategic [Wei85, Par85e]. strategies [CR83, CR84, Fos83, RZ89]. strategy [Ano83]. strength [Koe88]. Structural [OM89]. Structure [PMG85, Sha82, BCK^{+}82, DGKLM84, Dow86a, GM84, GGH^{+}82, SAP89]. structured [BH88, Jam84, Pet89, Ric86a, Sha82, TP86, WPM89]. structures [vI89]. Structuring [BK84, CH80, Mus85, Rin87, Rit87]. student [Jam84]. Studies [Sol86]. study [CHR86, DG88, Dun82, FWH84, Geh82, Leh82, Lei89, Rid84a]. stuffed [Bro84]. styles [NLY81]. submitted [Ano80b]. subsystems [BK84, KT84]. success [Law85a]. successful [Ben86]. Summary [BT85, Joh83, Sch85b, Bat83, Ger81a, Ger89, Gra83, Gro83, Kra87, LBC^{+}83, MRY86, Web83a]. superfluous [Eji83]. supervisors [RL89]. Support [Str88, BL88, Boy87, Chr89a, CRZ88, Dru82a, Hor85, KMS89, LPRS88, Leh86, LH88, Ost89, SBB^{+}81]. supported [GLR^{+}82]. supporting [Dub89]. supports [HL88]. Survey [Hua87, Wer87]. Surveys [Den82]. Susan [Gri82]. SW [Fal89]. SWAT [Car83a]. symbolic [Bea83, CSB82, JK83, Kin81]. symposium [Ano80b, Ree87, RBB^{+}82]. Syntax [HL85, Zel84a]. Syntax-directed [HL85]. synthesis [DeT84]. synthesizer [RT84]. SySL [TS89]. System [GRD84, Gil81b, Loy85, MM82, SCC^{+}83, WA85, BCD^{+}88, CK84, Cle88, CL83, DeT84, EGK84, Fri84, GF84, IS86, KWD81, Keu82, KK88, Kok89, KM89, KMS89, KSJ89, LPRS88, Liv87, Mac85, MCP89, Man87, MOT81, MCL85a, Mas85, MSR85, PMG85, Pir86, Put85b, Red88, Ric85, RZ89, Sca89b, SB89, Tar88, TD85, Web84, Whi86, Win86, Zah83, NP88, Uni87, Lem86]. Systematic [Ant89, Fri83]. Systems [Ano83, AL85, Ano80b, Aoy87, AWD89, BW83, BH82, BJKW88, CCR83, CG85, Cro85, DGGK87, Ger85, Ger89, GGW89, GM85, Gol87, Goo81a, Hag87, HLC82, KsKGC89, KSJ89, MCB82, MM82, MR88, Moo81, Neu86a, Neu86b, Neu86c, Neu86d, Neu87c, Neu88b, Pak89, Par85e, Pet89, RR85, Rei84, Rit87, Sha89, SBB^{+}81, Str88, WS82, WY89, Wil88, WMAJ89, Win85, Yel81]. Sytek [Owr85].

tale [Lev87]. tape [Dun85]. target [KT84, Tar88]. targets [Wei81]. tasking [Ncu88, RL89]. taxonomy [Zel84b]. team [BCK^{+}82, CRZ88]. TEC [Sof85d]. technique [Hil83, SHD84]. Techniques [HM84, Zuc89b, CGGS^{+}82, Gol87, KS85, RAO89, TS82, Kil89]. Technology [Ano83, BR85, Ben85, Cer81, FB81, Ger85, OO83, Rid84a, Weg82, Wil85]. Ted [Smo83]. temporal [Koo89a]. terminal [MOT81]. Test [Paz86, BHO89, BS89b, DG88, Fos84a, Fos85, HB89, RZ89, Wei89]. Testimony [Goo83, Par83b]. Testing [Fos89, Gro86, HN89, LGM^{+}86, RL89, AG83, Ben81, BS89b, CY84, CR83, CR84, CRZ88, EK85, Fos83, Ham89, HS89, JvB83, JW89, KM89, Lev87, OM89, PC89, ROT89, RAO89, Tai89, Whi85, YM89]. Texas [Hay81]. text [DRAP82, Nar89]. text-based [Nar89]. text-processing [DRAP82]. Thayer [Smo81]. their [Blu89, Tul89]. them [Kok88]. Theorem [Coo85, BK85b, KN85, KS85, Owr85, WO85]. theorem-prover [BK85b]. theorems [Mus85]. Theoretical [Ham89]. theories [Ger81b]. theory [Con81, Pra87]. Thierry [CS86]. Thomas [Smo81]. those [Ano80b]. Thoughts [Blu86, Cra85, Gog81, TS82]. Three [Mas85, vV89, Rid84a, RZ89]. time [BH88, GMMP89, Hil83, LS89, Pet89, RG89, SHG89, Zuc89b]. timeshared [GF84]. Timing [Pet89]. tolerance [Web89]. Tool [Dar89, DGGK87, DMS84, MCB82, Miy87, Rey85, RBB^{+}82, Zuc89a]. toolbox [BD83a]. toolkit [Sno86]. Tools [Per86b, BS89a, Baz85b, BL88, Bob87, CN87, Dut84, Gil84, Kar89, LPRS88, Mar88, MN85, Miy87, Pak89, Per86a, Raj82]. Toolset [Gla82b, ML88]. top [Hil89, Zah88a]. top-down [Hil89, Zah88a]. trace [DMP82]. traces [VF83]. tracking [Mur88a]. transaction [MM82]. transactions [Kai89b]. transform [HK88]. transformation [KM89]. Transformational [Red88]. transformations [Mir89a, Tho89b]. transition [Hen89]. transportability [Ske86]. treasurer [Sch81c]. treatment [vI89]. TRIAD [RR85]. trip [Buc86b]. Tripp [Ano80a]. trusted [Yel81]. trustworthy [Ger89]. TSL [RL89]. Two [Dun84, Ger82, McL85b, Per80, Smo81, Whi86, Gri82, LGM^{+}86, MB85, Miy82]. typed [WWFT88].

uit [Par81]. UK [CM88a, CKCM89, Sto85]. unconventional [Gil86a]. underlying [Gum82]. Understanding [Blu85]. unifying [WBMN88]. unit [AG83]. UNITY [GP89]. Universal [Wal83]. UNIX [Rei86, War85]. unsound [Cha83]. untitled [Bon81]. update [CS85]. US&R [Vit84]. USC [Sca89b]. Use [BEH82, RS82, Gal87, HLC82, MB89, Neu82, RVAR85, RW82]. User [BL88, Gom83, Mag84b, KS89]. user-definable [Mag84b]. users [GY87, You88]. Using [CSB82, DP89, Dav82, Gil84, Kin81, KK82, KM89, LM89, MN85, Mur88a, Mur88b, Red88, Rob89c, TS89].

Validating [How89]. validation [Dav85, DHPT89, WF87]. variant [Sch89]. variants [Nar89]. varying [BCK^{+}82]. VAX [Bea83]. VCG [Moo81]. Verification [Ano80c, Ben85, Smi85, Tar81, WA85, AL85, Ano85, BM83, Cer81, CG85, Cro85, DO85, Der85, DiV81, Di 85, FB81, Fos80, Ger81c, Ger85, Gio80, Gog81, GBM85, Hay81, KWD81, Lan85a, LN81, LvH81, MN85, MSR85, Mil81, Pro85, Put85b, SC85, WF87, Wei81, Whi85, Wil85, Yel81]. Verified [You89b, Goo81a, Ram89, SSDG81]. Verifying [Hoo85, Mac85, MCL85a, Den82]. VERking [Con81]. VERkshop [Ano80c, Tay85, Smi85, Neu81]. Verlag [Ger82, Smo81, Smo84]. Version [CSG^{+}88, Tho89a, Win88, BKL87, Cle89, Gor89, MCP89, Rei89]. version-control [Cle89]. versioning [LCDK89]. versions [RB89]. versus [Gil85, Lub86b]. VERUS^{T M} [Put85b]. very [Jan84, Ole85, WMAJ89]. veteran [Bal83]. via [OO83]. view [Cer81, CM88b, Gil86a]. Viewing [DMS84, AF89]. Viewpoint [Lei89]. Views [Sti89a, Sti89b, Lin84, MRY86, SM86, Tar81]. Viking [Neu85]. visible [Lem86]. visual [EMS89]. visualizing [SAP89]. VM [Mur88a, Mur88b]. VM/SP [Mur88a, Mur88b]. volume [MRY86]. voting [Wil88]. vs [Gre84, Jam84, Zah88b]. Vulnerabilities [Ros81].

W [BK83, Rei85, SL88, Smo84]. Wang [Fai88]. want [Gla82a]. ware [Gel85]. Wasserman [Bin89]. waterfall [Gil85]. Watsonville [Smi85]. way [GMMP89]. Weinberg [Kil83]. we’re [Ham80c]. West [SA85]. WG [Sch85b]. where [Dru82a]. which [Gar81]. widespread [Ric86b]. Wild [SA85]. Wild-West [SA85]. Wiley [Rei85]. WILI [SA85]. Will [Nel85]. winner [Par83a]. Winston [CS86]. wise [DH84, Ham83, Par85c]. within [BT79, Lem86, Mül83]. without [How89, Mil82]. wizards [Neu84]. Woods [Neu84]. work [LN81, Mil81]. Workshop [Dow86b, LGM^{+}86, Rat87, Cle88, PPT88, Smi85, Sof85d, Ano80c, CM88a, CKCM89, Gro86, Har88, Rid84b, RW86, Sol86, SZB82a, SZB82b, Win88]. workstation [LC84, Win86]. workstations [AL85]. world [Gar81, Neu84, Par85a, Par86b, Par86a, Par86c]. worse [Mey83]. write [Dun85]. write-enable [Dun85]. writing [Moo81]. Writings [Miy82]. written [DMP82].

year [Zav84c]. years [OCG^{+}81]. yellow [Buc86b]. Yourdon [Miy82, Miy82].

Z [LM89, Woo89]. Zen [Yac82, Neu82].

References

Ambriola:1989:OOC

[AB89] V. Ambriola and L. Bendix. Object-oriented configuration control. ACM SIGSOFT Software Engineering Notes, 14(7): 133–136, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arisawa:1982:EDM

[ABB^{+}82] Makoto Arisawa, G. David Bergland, Earle C. Bowers, John N.

Buxton, Robert A. Kelley, Norman L. Kerth, Sabina H. Saib, Peter D. Ting, Douglas A. Troy, and Stuart H. Zweben. Evaluation of design methodologies. ACM SIGSOFT Software Engineering Notes, 7(1):56–69, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arango:1988:FIP

[ABF88] Guillermo Arango, Ira Baxter, and Peter Freeman. A framework for incremental progress in the application of AI to software engineering. ACM SIGSOFT Software Engineering Notes, 13 (1):46–50, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alpern:1988:GAS

[ACR^{+}88] Bowen Alpern, Alan Carle, Barry Rosen, Peter Sweeney, and Kenneth Zadeck. Graph attribution as a specification paradigm. ACM SIGSOFT Software Engineering Notes, 13(5): 121–129, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alford:1982:PEE

[ADE^{+}82] Mack Alford, Larry E. Druffel, Ben Elliott, Richard Fairley, Ray Houghton, J. Ramanathan, Samuel T. Redwine, P. G. Richards, Alan R. Segal, Leon Stucki, and Jack C. Wileden. Proposal: evaluation of environments. ACM SIGSOFT Software Engineering Notes, 7 (1):36–45, January 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arango:1988:AAI

[AF88] Guillermo Arango and Peter Freeman. Application of artificial intelligence. ACM SIGSOFT Software Engineering Notes, 13 (1):32–38, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anderson:1989:PPS

[AF89] J. S. Anderson and S. Fickas. A proposed perspective shift: viewing specification design as a planning problem. ACM SIGSOFT Software Engineering Notes, 14(3):177–184, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Adler:1983:FMC

[AG83] Mike Adler and Michael A. Gray. A formalization of Myers causeeffect graphs for unit testing. ACM SIGSOFT Software Engineering Notes, 8(5):24–32, October 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:1985:MDF

[Agg85]  Rajeev Aggarwal.  Manage-

ment of dormant files. ACM SIGSOFT Software Engineering Notes, 10(5):60–61, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agresti:1981:SEI

[Agr81] William W. Agresti. Software engineering as industrial engineering. ACM SIGSOFT Software Engineering Notes, 6(5): 11–13, October 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Abdali:1985:EWD

[AL85] S. Kamal Abdali and Ralph L.

London. Exploiting workstations and displays in verification systems. ACM SIGSOFT Software Engineering Notes, 10(4):35–36, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Andreu:1982:CQM

[And82] Rafael Andreu. On comparisons of quality metrics. ACM SIGSOFT Software Engineering Notes, 7(1):76, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:1980:CBT

[Ano80a] Anonymous. Correction to Babb and Tripp paper. ACM SIGSOFT Software Engineering Notes, 5(1):29, January 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [BT79].

Anonymous:1980:SCA

[Ano80b] Anonymous. Some computer anecdotes from among those submitted in connection with the 7th symposium on operating systems principles. ACM SIGSOFT Software Engineering Notes, 5 (1):30–33, January 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:1980:WFV

[Ano80c]  Anonymous.  Workshop on

Formal Verification: (VERkshop). ACM SIGSOFT Software Engineering Notes, 5(3):4–47, July 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:1983:STA

[Ano83] Anonymous. Software Technology for Adaptable, Reliable Systems (STARS) program strategy. ACM SIGSOFT Software Engineering Notes, 8(2):56–108, April 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:1984:ASE

[Ano84] Anonymous. Abstracts in software engineering. ACM SIG-SOFT Software Engineering Notes, 9(1):52–63, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:1985:OFL

[Ano85] Anonymous. Old-fashioned logic for verification. ACM SIGSOFT Software Engineering Notes, 10 (4):87–89, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Antoy:1989:SDA

[Ant89] S. Antoy. Systematic design of algebraic specifications. ACM SIGSOFT Software Engineering Notes, 14(3):278–280, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aoyama:1987:CDS

[Aoy87] M. Aoyama. Concurrent development of software systems. ACM SIGSOFT Software Engineering Notes, 12(3):20–24, July 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arango:1989:DAA

[Ara89] G. Arango. Domain analysis: from art form to engineering discipline. ACM SIGSOFT Software Engineering Notes, 14 (3):152–159, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:1986:CAS

[Ard86] Mark A. Ardis. Comparison of algebraic and state-machine specification methods. ACM SIGSOFT Software Engineering Notes, 11(4):54–56, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ashok:1989:PMS

[ARSV89] V. Ashok, J. Ramanathan, S. Sarkar, and V. Venugopal. Process modelling in software environments. ACM SIGSOFT Software Engineering Notes, 14 (4):36–42, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agrawal:1989:BDB

[AS89] H. Agrawal and E. H. Spafford. Bibliography on debugging and backtracking. ACM SIGSOFT Software Engineering Notes, 14 (2):49–56, April 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Avrunin:1989:EAA

[AWD89] G. Avrunin, J. Wileden, and L. Dillon. Experiments in automated analysis of concurrent software systems. ACM SIGSOFT Software Engineering Notes, 14(8):124–130, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Babb:1988:LM

[Bab88] Robert G. Babb II. Language and models. ACM SIGSOFT Software Engineering Notes, 13 (1):43–45, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baldwin:1982:RSI

[Bal82] Roger R. Baldwin. Reportage on spring 1982 IEEE COMPCON conference. ACM SIGSOFT Software Engineering Notes, 7 (2):13–20, April 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baldwin:1983:LNV

[Bal83] Roger R. Baldwin. Leaves from the notebook of a veteran programmer/analyst: my sixteen months of experience with Jackson program design methodology. ACM SIGSOFT Software Engineering Notes, 8(1):16, January 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balzer:1986:PEP

[Bal86] Robert Balzer. Program enhancement: a position paper. ACM SIGSOFT Software Engineering Notes, 11(4):66–68, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balzer:1989:PPP

[Bal89] R. Balzer. Process programming: passing into a new phase. ACM SIGSOFT Software Engineering Notes, 14(4):43–45, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barstow:1982:RPA

[Bar82] David Barstow. Rapid prototyping, automatic programming, and experimental sciences. ACM SIGSOFT Software Engineering Notes, 7(5):33–34, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bates:1983:RDD

[Bat83] Peter Bates. Requirements/ design debugging: session summary. ACM SIGSOFT Software Engineering Notes, 8(4):32–33, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bazelmans:1985:ECM

[Baz85a] Rudy Bazelmans. Evolution of configuration management. ACM SIGSOFT Software Engineering Notes, 10(5):37–46, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bazelmans:1985:PRT

[Baz85b] Rudy Bazelmans. Productivity: the role of the tools group. ACM SIGSOFT Software Engineering Notes, 10(3):63–75, July 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boeh:1989:APP

[BB89] B. Boeh and F. Belz. Applying process programming to the model. ACM SIGSOFT Software Engineering Notes, 14 (4):46–56, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Borras:1988:CS

[BCD^{+}88] P. Borras, D. Clement, Th. Despeyroux, J. Incerpi, G. Kahn, B. Lang, and V. Pascual. Centaur: the system. ACM SIGSOFT Software Engineering Notes, 13(5):14–24, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Belady:1982:IHE

[BCE^{+}82] Les Belady, George R. Cannon, Bryan Edwards, Elizabeth Kruesi, Tim Lindquist, and Mark Weiser. Investigation of human engineering: for software environments. ACM SIGSOFT Software Engineering Notes, 7 (1):69–72, January 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bailey:1982:EEI

[BCK^{+}82] John W. Bailey, Ken Christensen, Helmut Krcmar, Jean-Louis Lassez, Vincent Y. Shen, and Scott N. Woodfield. Evaluating the effect of interprogrammer communication: levels on productivity by varying programming team structure. ACM SIGSOFT Software Engineering Notes, 7(1):10–20, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barra:1983:SPT

[BD83a] Knut Barra and Hans Petter Dahle. SIMOB: a portable toolbox for observation of SIMULA executions. ACM SIGSOFT Software Engineering Notes, 8 (4):121–122, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bolognani:1983:MSE

[BD83b] Mario Bolognani and Sergio De Julio. A Master on Software Engineering Degree in Southern Italy. ACM SIGSOFT Software Engineering Notes, 8(5):

37–38, October 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baiardi:1983:DDC

[BDM^{+}83] F. Baiardi, N. De Francesco, E. Matteoli, S. Stefanini, and G. Vaglini. Development of a debugger for a concurrent language. ACM SIGSOFT Software Engineering Notes, 8(4):98–106, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beander:1983:VDI

[Bea83] Bert Beander. VAX DEBUG: an interactive, symbolic, multilingual debugger. ACM SIGSOFT Software Engineering Notes, 8 (4):173–179, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bradshaw:1982:UDA

[BEH82] Franklyn T. Bradshaw, George W. Ernst, and Raymond J. Hookway. Use of data abstraction in process specification. ACM SIGSOFT Software Engineering Notes, 7(3):16–21, July 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benson:1981:PEA

[Ben81] Jeoff Benson. A preliminary experiment in automated software testing. ACM SIGSOFT Software Engineering Notes, 6 (3):68–75, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benzel:1985:VTA

[Ben85] Terry C. Vickers Benzel. Verification technology and the A1 criteria. ACM SIGSOFT Software Engineering Notes, 10(4): 108–109, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benson:1986:RFS

[Ben86] David B. Benson. The role of failure in successful design. ACM SIGSOFT Software Engineering Notes, 11(1):15, January 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berry:1985:IJP

[Ber85] Daniel M. Berry. An ina Jo(R)

proof manager for the formal development method. ACM SIGSOFT Software Engineering Notes, 10(4):19–25, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boudier:1988:OPP

[BGMT88] Gerard Boudier, Ferdinando Gallo, Regis Minot, and Ian Thomas. An overview of PCTE and PCTE+. ACM SIGSOFT Software Engineering Notes, 13 (5):248–257, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balzer:1982:OSB

[BGW82] Robert M. Balzer, Neil M. Goldman, and David S. Wile. Operational specification as the basis for rapid prototyping. ACM SIGSOFT Software Engineering Notes, 7(5):3–16, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blum:1982:RPI

[BH82] Bruce I. Blum and Raymond C. Houghton. Rapid prototyping of information management systems. ACM SIGSOFT Software Engineering Notes, 7(5): 35–38, December 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bruegge:1983:GPE

[BH83] Bernd Bruegge and Peter Hibbard. Generalized path expressions: a high level debugging mechanism. ACM SIGSOFT Software Engineering Notes, 8 (4):34–44, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blumofe:1988:ERT

[BH88] R. Blumofe and A. Hecht. Executing real-time structured analysis specifications. ACM SIGSOFT Software Engineering Notes, 13(3):32–40, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balcer:1989:AGT

[BHO89] M. Balcer, W. Hasling, and T. Ostrand. Automatic generation of test scripts from formal test specifications. ACM SIGSOFT Software Engineering Notes, 14(8):210–218, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Binder:1989:ARO

[Bin89] R. V. Binder. Article review:

An OOSD Method for Code Generation by A. I. Wasserman, P. A. Pircher, R. J. Muller (SEN 14 1, January 1989). ACM SIGSOFT Software Engineering Notes, 14(2):79, April 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bruyn:1988:ESM

[BJKW88] William Bruyn, Randall Jense, Dinesh Keskar, and Paul Ward. An extended systems modeling language (ESML). ACM SIGSOFT Software Engineering Notes, 13(1):58–67, January 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bryant:1983:BWB

[BK83] A. Bryant and J. A. Kirkham.

B. W. Boehm, Software Engineering Economics: a review essay. ACM SIGSOFT Software Engineering Notes, 8(3):44–60, July 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bonet:1984:SSE

[BK84] Rafael Bonet and Antonio Kung.

Structuring into subsystems: the experience of a prototyping approach. ACM SIGSOFT Software Engineering Notes, 9(5): 23–27, October 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boebert:1985:EAM

[BK85a] W. E. Boebert and R. Y. Kain. The extended access matrix model of computer security. ACM SIGSOFT Software Engineering Notes, 10(4):119–125, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyer:1985:PTP

[BK85b] Bob Boyer and Matt Kaufmann. A prototype theorem-prover for a higher-order functional language. ACM SIGSOFT Software Engineering Notes, 10(4):72–74, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brilliant:1987:CCP

[BKL87] S. S. Brilliant, J. C. Knight, and N. G. Leveson. The consistent comparison problem in N-version software. ACM SIGSOFT Software Engineering Notes, 12(1):29–34, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beaudouin-Lafon:1988:UIS

[BL88] Michel Beaudouin-Lafon. User interface support for the integration of software tools: an iconic model of interaction. ACM SIGSOFT Software Engineering Notes, 13(5):143–152, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blum:1982:LCD

[Blu82] Bruce I. Blum. The life cycle:

a debate over alternate models. ACM SIGSOFT Software Engineering Notes, 7(4):18–20, October 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blum:1983:SMA

[Blu83] Bruce I. Blum. Still more about rapid prototyping. ACM SIGSOFT Software Engineering Notes, 8(3):9–11, July 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blum:1985:USP

[Blu85] Bruce Blum. Understanding the software paradox. ACM SIGSOFT Software Engineering Notes, 10(1):43–46, January 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blum:1986:TSP

[Blu86] B. Blum. Thoughts on the software process. ACM SIGSOFT Software Engineering Notes, 11 (4):25–26, August 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blum:1987:GRF

[Blu87]  B. I. Blum.  GAO report

FGMSD-80-4 revisited. ACM SIGSOFT Software Engineering Notes, 12(1):49–50, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blum:1989:DTA

[Blu89] B. I. Blum. On data and their analysis. ACM SIGSOFT Software Engineering Notes, 14 (5):24–34, July 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyer:1983:PVP

[BM83] Robert S. Boyer and J. Strother Moore. Program verification prize. ACM SIGSOFT Software Engineering Notes, 8(3):64, July 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bobbie:1987:PTA

[Bob87] P. O. Bobbie. Productivity through automated tools. ACM SIGSOFT Software Engineering Notes, 12(2):30–31, April 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boehm:1986:SMS

[Boe86] B. Boehm. A spiral model of software development and enhancement. ACM SIGSOFT Software Engineering Notes, 11 (4):14–24, August 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boldyreff:1989:RSC

[Bol89] C. Boldyreff. Reuse, software concepts, descriptive methods, and the practitioner project. ACM SIGSOFT Software Engineering Notes, 14(2):25–31, April 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bonyun:1981:UCG

[Bon81] David Bonyun. An untitled Canada goose. ACM SIGSOFT Software Engineering Notes, 6 (3):60–61, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Botting:1986:FDI

[Bot86] Richard J. Botting. Into the fourth dimension—an introduction to dynamic analysis and design. ACM SIGSOFT Software Engineering Notes, 11(2):36–48, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowen:1986:CSQ

[Bow86] John B. Bowen. Comments on software quality evaluation guidebook. ACM SIGSOFT Software Engineering Notes, 11 (1):19, January 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boyd:1987:SDD

[Boy87] S. Boyd. Status of DAPSE distributed Ada support. ACM SIGSOFT Software Engineering Notes, 12(2):56, April 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Belady:1985:MST

[BR85] Les Belady and Charles Richter. The MCC software technology program. ACM SIGSOFT Software Engineering Notes, 10(3): 33–36, July 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Braunl:1989:SLP

[Brä89] T. Bräunl. A specification language for parallel architectures and algorithms. ACM SIGSOFT Software Engineering Notes, 14 (3):49–51, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brereton:1987:SEE

[Bre87] P. Brereton. 1987 software engineering environments conference. ACM SIGSOFT Software Engineering Notes, 12(4):35, October 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brown:1984:MAS

[Bro84] Dave Brown. My accordian’s stuffed full of paper: or why we did so badly in the design phase. ACM SIGSOFT Software Engineering Notes, 9(4):58–60, July 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See response [Rob84].

Baram:1989:SCA

[BS89a] G. Baram and G. Steinberg.

Selection criteria for analysis and design CASE tools. ACM SIGSOFT Software Engineering Notes, 14(6):73–80, October 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bieman:1989:ENT

[BS89b] J. Bieman and J. Schultz. Estimating the number of test cases required to satisfy the all-du-paths testing criterion. ACM SIGSOFT Software Engineering Notes, 14(8):179–186, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barstow:1984:IPE

[BSSS84] David R. Barstow, Howard E.

Shrobe, Erik Sandewall, and Stephen W. Smoliar. Interactive programming environments. ACM SIGSOFT Software Engineering Notes, 9(4):56–58, July 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Babb:1979:ADA

[BT79] Robert G. Babb and Leonard L.

Tripp. An approach to defining areas within the field of software engineering. ACM SIGSOFT Software Engineering Notes, 4 (4):9–17, October 1979. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See correction [Ano80a].

Branstad:1985:SSI

[BT85] Martha Branstad and Leonard L. Tripp. Summary of SESAW III. ACM SIGSOFT Software Engineering Notes, 10(5):33–36, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Buckley:1986:RQA

[Buc86a] F. Buckley. A review of the quality aspects of the approved software engineering standards. ACM SIGSOFT Software Engineering Notes, 11(1):60–63, January 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Buckley:1986:SSQ

[Buc86b] Fletcher J. Buckley. The search for software quality, or one more trip down the yellow brick road. ACM SIGSOFT Software Engineering Notes, 11(1):16–18, January 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Buckley:1987:RSP

[Buc87] F. J. Buckley. The roles of a SQA person. ACM SIGSOFT Software Engineering Notes, 12 (3):42–44, July 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Buckley:1988:BES

[Buc88] Fletcher J. Buckley. Balloting experiment with software engineering standards. ACM SIGSOFT Software Engineering Notes, 13(1):51, January 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baalbergen:1989:DAC

[BVT89] E. H. Baalbergen, K. Verstoep, and A. S. Tanenbaum. On the design of the Amoeba configuration manager. ACM SIGSOFT Software Engineering Notes, 14 (7):15–22, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bates:1983:AHL

[BW83] Peter Bates and Jack C. Wileden. An approach to high-level debugging of distributed systems: preliminary draft. ACM SIGSOFT Software Engineering Notes, 8(4):107–111, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cardell:1983:MDS

[Car83a] James R. Cardell. Multilingual debugging with the SWAT highlevel debugger. ACM SIGSOFT Software Engineering Notes, 8 (4):180–189, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cargill:1983:BDP

[Car83b] Thomas A. Cargill. The Blit debugger: preliminary draft. ACM SIGSOFT Software Engineering Notes, 8(4):190–200, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Church:1986:AAS

[CCAJ86] VE Church, D. N. Card, W. W. Agresti, and Q. L. Jordan. An approach for assessing software prototypes. ACM SIGSOFT Software Engineering Notes, 11 (3):65–76, July 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ceri:1983:RDB

[CCR83] S. Ceri and S. Crespi-Reghizzi. Relational data bases in the design of program construction systems. ACM SIGSOFT Software Engineering Notes, 8(3):17–29, July 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chesi:1984:IIS

[CDF^{+}84] M. Chesi, E. Dameri, M. P. Franceschi, M. G. Gatti, and C. Simonelli. ISDE: an interactive software development environment. ACM SIGSOFT Software Engineering Notes, 9 (3):81–88, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cerf:1981:VVT

[Cer81] Vinton G. Cerf. A view of verification technology. ACM SIGSOFT Software Engineering Notes, 6(3):3–4, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Craigen:1985:OVS

[CG85] D. Craigen and D. Good. Overview of verification systems.

ACM SIGSOFT Software Engineering Notes, 10(4):1, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Curtis:1982:AFF

[CGGS^{+}82] Bill Curtis, John Gannon, Shirley Gloss-Soler, Amrit Goel, Isao Miyamoto, David Panzl, and Roger Warburton. The application of fractional factorial techniques: to the design of software experiments, or, confounding the issues. ACM SIGSOFT Software Engineering Notes, 7 (1):20–30, January 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cashman:1980:COA

[CH80] Paul M. Cashman and Anatol W. Holt. A communicationoriented approach to structuring the software maintenance environment. ACM SIGSOFT Software Engineering Notes, 5 (1):4–17, January 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chambers:1983:EMB

[Cha83] John M. Chambers. Evolution and maintenance are both semantically unsound, and also hazardous to the health of our public image. ACM SIGSOFT Software Engineering Notes, 8 (3):70–71, July 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See comments [Leh83].

Chapin:1984:SMF

[Cha84] Ned Chapin. Software maintenance with fourth-generation languages. ACM SIGSOFT Software Engineering Notes, 9(1): 41–42, January 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheatham:1989:ACP

[Che89] T. E. Cheatham, Jr. Activity coordination programs. ACM SIGSOFT Software Engineering Notes, 14(4):57–60, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chelini:1986:PSA

[CHR86] James V. Chelini, Edmund B. Hudson, and Stephen M. Reidy. A preliminary study of Ada expansion ratios. ACM SIGSOFT Software Engineering Notes, 11 (1):35–46, January 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chroust:1988:MIC

[Chr88] G. Chroust. Models, instances, and concurrency in software development. ACM SIGSOFT Software Engineering Notes, 13 (3):41–42, July 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chroust:1989:ADP

[Chr89a] G. Chroust. Application development project support. ACM SIGSOFT Software Engineering Notes, 14(5):83–104, July 1989.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chroust:1989:DIE

[Chr89b] G. Chroust. Duplicate instances of elements of a software process model. ACM SIGSOFT Software Engineering Notes, 14 (4):61–64, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Campbell:1984:SPS

[CK84] Roy H. Campbell and Peter A.

Kirslis. The SAGA project: a system for software development. ACM SIGSOFT Software Engineering Notes, 9(3):73–80, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:1989:RUS

[CKCM89] S. D. Cooper, R. J. Kenning, F. W. Calliss, and M. Munro. Report on 2nd UK Software Maintenance Workshop. ACM SIGSOFT Software Engineering Notes, 14(5):124–127, July 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cook:1983:DDM

[CL83] Robert P. Cook and Insup Lee.

DYMOS: a dynamic modification system. ACM SIGSOFT Software Engineering Notes, 8 (4):201–202, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clemm:1988:WSP

[Cle88]  Geoffrey M. Clemm.  The

workshop system: a practical knowledge-based software environment. ACM SIGSOFT Software Engineering Notes, 13(5): 55–64, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clemm:1989:RVC

[Cle89]  G. M. Clemm.  Replacing

version-control with job-control. ACM SIGSOFT Software Engineering Notes, 14(7):162–169, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clune:1989:SFG

[Clu89] T. R. Clune. SE: the field and the ground. ACM SIGSOFT Software Engineering Notes, 14 (6):50, October 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Calliss:1988:FUS

[CM88a] F. Calliss and M. Munro. The First UK Software Maintenance Workshop. ACM SIGSOFT Software Engineering Notes, 13 (3):52–53, July 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chandy:1988:AVF

[CM88b] K. Mani Chandy and J. Misra. Another view on “fairness”. ACM SIGSOFT Software Engineering Notes, 13(3):20, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij88, SL88].

Choi:1989:PSM

[CMP89] B. Choi, A. Mathur, and B. Pattison. P^{M} othra: scheduling mutants for execution on a hypercube. ACM SIGSOFT Software Engineering Notes, 14(8): 58–65, December 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cook:1987:PSC

[CN87] C. C. Cook and M. Nanja. Prototype software complexity metrics tools. ACM SIGSOFT Software Engineering Notes, 12(3):58–60, July 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Coad:1987:DAD

[Coa87] P. Coad, Jr. Design approaches and DoD-STD-2167. ACM SIGSOFT Software Engineering Notes, 12(1):50, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Constable:1981:VCS

[Con81] Robert L. Constable. VERking in constructive set theory. ACM SIGSOFT Software Engineering Notes, 6(3):58–60, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Conde:1986:BRA

[Con86] Dan Conde. Book review: Abstraction and Specification in Program Development by Barbara Liskov and John Guttag (MIT Press/McGraw-Hill, 1988, 469 pages, ISBN 0-262-12112-3). ACM SIGSOFT Software Engineering Notes, 11(3):79–81, July 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cook:1982:SMI

[Coo82] Michael L. Cook. Software metrics: an introduction and annotated bibliography. ACM SIGSOFT Software Engineering Notes, 7(2):41–60, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:1985:OTP

[Coo85] D. Cooper. Overview of theorem proving. ACM SIGSOFT Software Engineering Notes, 10 (4):53–54, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Costello:1984:SEU

[Cos84] Scott H. Costello. Software engineering under deadline pressure. ACM SIGSOFT Software Engineering Notes, 9(5):15–19, October 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clarke:1983:AES

[CR83] Lori A. Clarke and Debra J. Richardson. The application of error-sensitive testing strategies to debugging. ACM SIGSOFT Software Engineering Notes, 8 (4):45–52, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See comment [Fos83] and replies [CR84, Fos84b].

Clarke:1984:RFC

[CR84] Lori A. Clarke and Debra J.

Richardson. A reply to Foster’s “Comment on ’The application of error-sensitive testing strategies to debugging”’. ACM SIGSOFT Software Engineering Notes, 9(1):24–28, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [CR83, Fos83, Fos84b].

Craigen:1985:STA

[Cra85] Dan Craigen. Some thoughts arising from a language design effort. ACM SIGSOFT Software Engineering Notes, 10(4):80–81, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Crocker:1985:ERP

[Cro85] Stephen D. Crocker. Engineering requirements for production quality verification systems. ACM SIGSOFT Software Engineering Notes, 10(4):15–16, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clarke:1988:TSE

[CRZ88] Lori A. Clarke, Debra J.

Richardson, and Steven J. Zeil. TEAM: a support environment for testing, evaluation, and analysis. ACM SIGSOFT Software Engineering Notes, 13(5): 153–162, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Craigen:1985:EU

[CS85] Dan Craigen and Mark Saaltink. An EVES update. ACM SIGSOFT Software Engineering Notes, 10(4):33–34, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chapman:1986:BRS

[CS86] Gary Chapman and Marvin Schaefer. Book review: Software Engineering by Thierry Breton and Denis Beneich (Holt, Rinehart Winston). ACM SIGSOFT Software Engineering Notes, 11 (3):77–78, July 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Choi:1989:ACC

[CS89] S. C. Choi and W. S. Scacchi. Assuring the correctness of configured software descriptions. ACM SIGSOFT Software Engineering Notes, 14(7): 66–75, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cohen:1982:USE

[CSB82] Donald Cohen, William Swartout, and Robert Balzer. Using symbolic execution to characterize behavior. ACM SIGSOFT Software Engineering Notes, 7(5):

25–32, December 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cohen:1988:VMG

[CSG^{+}88] Ellis S. Cohen, Dilip A. Soni, Raimund Gluecker, William M. Hasling, Robert W. Schwanke, and Michael E. Wagner. Version management in gypsy. ACM SIGSOFT Software Engineering Notes, 13(5):201–215, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Currey:1986:LEL

[Cur86] Richard Currey. Letter on ERTS and LEGO. ACM SIGSOFT Software Engineering Notes, 11 (5):51, October 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chan:1984:PTA

[CY84] Mee-Yee Chan and Siu-Yuen Yam. A program testing assistant for BASIC-PLUS. ACM SIGSOFT Software Engineering Notes, 9(2):89–103, April 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dart:1989:TCA

[Dar89] S. Dart. Tool configuration assistant. ACM SIGSOFT Software Engineering Notes, 14(7): 110–113, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:1982:RPU

[Dav82] Alan M. Davis. Rapid prototyping using executable requirements specifications. ACM SIGSOFT Software Engineering Notes, 7(5):39–44, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:1985:RAV

[Dav85] Daniel Davis. Resource abstraction and validation. ACM SIGSOFT Software Engineering Notes, 10(4):110, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Delisle:1989:FSE

[DdG89] N. Delisle and d. Garlan. Formally specifying electronic instruments. ACM SIGSOFT Software Engineering Notes, 14 (3):242–248, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Denning:1980:ENI

[Den80] Dorothy E. Denning. Embellishments to the note on information flow into arrays. ACM SIGSOFT Software Engineering Notes, 5(2):15–16, April 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Denning:1982:RVS

[Den82] Dorothy Denning. Review of Verifying security by M. H. Cheheyl, M. Gasser, G. A. Huff, and J. K. Millen. Computing Surveys 13, 3 (September 1981). ACM SIGSOFT Software Engineering Notes, 7(1):80–81, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dershowitz:1985:RV

[Der85] Nachum Dershowitz. Rewriting and verification. ACM SIGSOFT Software Engineering Notes, 10(4):60, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DeTreville:1984:PIS

[DeT84] John DeTreville. Phoan: an intelligent system for distributed control synthesis. ACM SIG-SOFT Software Engineering Notes, 9(3):96–103, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Donnelly:1988:CST

[DG88] K. F. Donnelly and K. A. Gluck.

A case study in test environment evolution. ACM SIGSOFT Software Engineering Notes, 13 (1):22–28, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dahler:1987:GTD

[DGGK87] J. Dähler, P. Gerber, H.-P.

Gisiger, and A. Kündig. A graphical tool for the design and prototyping of distributed systems. ACM SIGSOFT Software Engineering Notes, 12(3):25–36, July 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Donzeau-Gouge:1984:DSM

[DGKLM84] V. Donzeau-Gouge, G. Kahn, B. Lang, and B. Mélèse. Document structure and modularity in mentor. ACM SIGSOFT Software Engineering Notes, 9 (3):141–148, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Diaz-Herrera:1984:PPS

[DH84] Jorge L. Diaz-Herrera. Pragmatic problems with step-wise refinement program development. ACM SIGSOFT Software Engineering Notes, 9(2):80–88, April 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dinsart:1989:ODV

[DHPT89] A. Dinsart, B. Hucq, N. Piperakis, and P. J. Thollembeck. Object derivation and validation from a data base definition. ACM SIGSOFT Software Engineering Notes, 14(7): 170–178, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DiVito:1985:TDB

[Di 85] Ben Di Vito. Towards a definition of “Beyond AI” verification. ACM SIGSOFT Software Engineering Notes, 10(4):101–102, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dijkstra:1981:APP

[Dij81] Edsger W. Dijkstra. American programming’s plight. ACM SIGSOFT Software Engineering Notes, 6(1):5, January 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dijkstra:1988:PPF

[Dij88] Edsger W. Dijkstra. Position paper on “fairness”. ACM SIGSOFT Software Engineering Notes, 13(2):18–20, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See comments [SL88, CM88b].

DiVito:1981:MVA

[DiV81] Ben DiVito. A mechanical verification of the alternating bit protocol. ACM SIGSOFT Software Engineering Notes, 6(3):8–13, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dixon:1982:RPM

[DMP82] John K. Dixon, John McLean, and David L. Parnas. Rapid prototyping by means of abstract module specifications written as trace axioms. ACM SIGSOFT Software Engineering Notes, 7 (5):45–49, December 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Delisle:1984:VPE

[DMS84] Norman M. Delisle, David E.

Menicosy, and Mayer D. Schwartz.

Viewing a programming environment as a single tool. ACM SIGSOFT Software Engineering Notes, 9(3):49–56, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dahl:1985:PSV

[DO85] Ole-Johan Dahl and Olaf Owe. A presentation of the specification and verification project “ABEL”. ACM SIGSOFT Software Engineering Notes, 10(4): 28–32, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dowson:1986:SSP

[Dow86a] M. Dowson. The structure of the software process. ACM SIGSOFT Software Engineering Notes, 11(4):6–8, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dowson:1986:PPI

[Dow86b] Mark Dowson. Position paper: International Workshop on the Software Process and Software Environments. ACM SIGSOFT Software Engineering Notes, 11 (4):26–27, August 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dershowitz:1985:CR

[DP85] Nachum Dershowitz and David A. Plaisted. Conditional rewriting. ACM SIGSOFT Software Engineering Notes, 10(4):55–59, August 1985. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

DIppolito:1989:SDU

[DP89] R. S. D’Ippolito and C. P.

Plinta. Software development using models. ACM SIGSOFT Software Engineering Notes, 14 (3):140–142, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dodd:1982:PLT

[DRAP82] W. P. Dodd, P. Ramsay, T. H.

Axford, and D. G. Parkyn. A prototyping language for textprocessing applications. ACM SIGSOFT Software Engineering Notes, 7(5):50–53, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Druffel:1982:NPD

[Dru82a] Larry E. Druffel. The need for a programming discipline to support the APSE: where does the APSE path lead? ACM SIGSOFT Software Engineering Notes, 7(3):12–13, July 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Druffel:1982:PEA

[Dru82b] Larry E. Druffel. The potential effect of Ada on software engineering in the 1980’s. ACM SIGSOFT Software Engineering Notes, 7(3):5–11, July 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dhama:1989:NMS

[DS89] H. S. Dhama and D. Shtern. A net method for specification of reusable software. ACM SIGSOFT Software Engineering Notes, 14(3):137–139, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubois:1989:LAS

[Dub89] M. S. Dubois. A logic of action for supporting goal-oriented elaborations of requirements. ACM SIGSOFT Software Engineering Notes, 14(3):160–168, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Duncan:1982:PAC

[Dun82] Arthur G. Duncan. Prototyping in ADA: a case study. ACM SIGSOFT Software Engineering Notes, 7(5):54–60, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dunn:1984:TOL

[Dun84] Dick Dunn. Two observations on large software projects. ACM SIGSOFT Software Engineering Notes, 9(5):8–10, October 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dunn:1985:ESS

[Dun85] Dick Dunn. The emerging social significance of the magnetictape write-enable ring. ACM SIGSOFT Software Engineering Notes, 10(2):47–50, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dutton:1984:RTP

[Dut84] Clifton C. Dutton. Requirements tools: phase i in a software development methodology. ACM SIGSOFT Software Engineering Notes, 9(1):19–23, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Early:1986:RSR

[Ear86] M. Early. Relating software requirements to software design. ACM SIGSOFT Software Engineering Notes, 11(3):37–39, July 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Estublier:1984:PEC

[EGK84] J. Estublier, S. Ghoul, and S. Krakowiak. Preliminary experience with a configuration control system for modular programs. ACM SIGSOFT Software Engineering Notes, 9(3): 149–156, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eidnes:1989:SCC

[EHW89] H. Eidnes, D. O. Hallsteinsen, and D. H. Wanvik. Separate compilation in CHIPSY. ACM SIGSOFT Software Engineering Notes, 14(7):42–45, November 1989. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Ejiogu:1983:ESS

[Eji83] Lem O. Ejiogu. Evolution is significantly superfluous; maintenance is semantically sound. ACM SIGSOFT Software Engineering Notes, 8(2):13, April 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eckmann:1985:IIE

[EK85] Steven T. Eckmann and Richard A. Kemmerer. INATEST: an interactive environment for testing formal specifications. ACM SIGSOFT Software Engineering Notes, 10(4):17–18, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eisenbach:1989:DFD

[EMS89] S. Eisenbach, L. McLoughlin, and C. Sadler. Data-flow design as a visual programming language. ACM SIGSOFT Software Engineering Notes, 14(3): 281–283, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Evangelist:1986:FPS

[EP86] M. Evangelist and J. Pellegrin. Foundational problems in software process research. ACM SIGSOFT Software Engineering Notes, 11(4):12–13, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fairley:1988:PMA

[Fai88] Richard E. Fairley. Post-mortem analysis of software engineering at Wang institute. ACM SIGSOFT Software Engineering Notes, 13(2):41–47, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Falkerngerg:1989:CML

[Fal89] B. Falkerngerg. Configuration management for a large (SW) development. ACM SIGSOFT Software Engineering Notes, 14 (7):34–37, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feiertag:1981:AED

[FB81] R. Feiertag and T. Berson. An avenue for exploitation and development of verification technology. ACM SIGSOFT Software Engineering Notes, 6(3): 45–46, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feather:1982:MRP

[Fea82] Martin S. Feather. Mappings for rapid prototyping. ACM SIGSOFT Software Engineering Notes, 7(5):17–24, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feather:1989:DIW

[Fea89] M. S. Feather. Detecting interference when merging specification evolutions. ACM SIGSOFT Software Engineering Notes, 14 (3):169–176, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ferrante:1983:HLL

[Fer83] Jeanne Ferrante. High level language debugging with a compiler. ACM SIGSOFT Software Engineering Notes, 8(4): 123–129, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fernstrom:1989:DCP

[Fer89] C. Fernstrom. Design considerations for process-driven software environments. ACM SIGSOFT Software Engineering Notes, 14 (4):65–67, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkelstein:1989:MS

[FF89] A. Finkelstein and H. Fuks. Multiparty specification. ACM SIGSOFT Software Engineering Notes, 14(3):185–195, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkelstein:1989:CDF

[FFNS89] A. Finkelstein, H. Fuks, C. Niskier, and M. Sadler. Constructing a dialogic framework for software development. ACM SIGSOFT Software Engineering Notes, 14 (4):68–72, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fischer:1984:PLB

[FJM^{+}84] C. N. Fischer, Gregory F. Johnson, Jon Mauney, Anil Pal, and Daniel L. Stock. The Poe language-based editor project. ACM SIGSOFT Software Engineering Notes, 9(3):21–29, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Floyd:1988:PCS

[Flo88] Christiane Floyd. A paradigm change in software engineering. ACM SIGSOFT Software Engineering Notes, 13(2):25–38, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ford:1982:IPD

[FM82] Ray Ford and Chris Marlin.

Implementation prototypes in the development of programming language features. ACM SIGSOFT Software Engineering Notes, 7(5):61–66, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Forman:1989:DDM

[For89] I. R. Forman. Design by decomposition of multiparty interactions in raddle87. ACM SIGSOFT Software Engineering Notes, 14(3):2–10, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Foster:1980:CEA

[Fos80] Ken Foster. A comment on: “An exercise in automatic program verification”. ACM SIGSOFT Software Engineering Notes, 5 (1):36–37, January 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Foster:1983:CAE

[Fos83] Kenneth A. Foster. Comment on “The application of errorsensitive testing strategies to debugging”. ACM SIGSOFT Software Engineering Notes, 8(5): 40–42, October 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [CR83, CR84, Fos84b].

Foster:1984:STD

[Fos84a] K. A. Foster. Sensitive test data for logic expressions. ACM SIGSOFT Software Engineering Notes, 9(2):120–125, April 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Foster:1984:RR

[Fos84b] Kenneth A. Foster. A reply to a reply. ACM SIGSOFT Software Engineering Notes, 9 (2):126, April 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [CR84, Fos83].

Foster:1985:RES

[Fos85] K. A. Foster. Revision of an error sensitive test rule. ACM SIGSOFT Software Engineering Notes, 10(1):62–67, January 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Foster:1989:TI

[Fos89] K. A. Foster. Testing by identifying. ACM SIGSOFT Software Engineering Notes, 14(5):64–65, July 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fritsch:1989:MCS

[FP89] C. A. Fritsch and D. L. Perry. A manager/controller for the software development process. ACM SIGSOFT Software Engineering Notes, 14(4):73–75, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frankel:1982:CCS

[Fra82] Eric C. Frankel. A contracts course for software engineers. ACM SIGSOFT Software Engineering Notes, 7(4):34–38, October 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fritzson:1983:SAA

[Fri83] Peter Fritzson. A systematic approach to advanced debugging: incremental compilation. ACM SIGSOFT Software Engineering Notes, 8(4):130–139, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fritzson:1984:PED

[Fri84]  Peter Fritzson.  Preliminary

experience from the dice system a distributed incremental compiling environment. ACM SIGSOFT Software Engineering Notes, 9(3):113–123, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Freeman:1984:CSD

[FWH84] Peter Freeman, Anthony I. Wasserman, and Raymond C. Houghton. Comparing software development methodologies for Ada: a study plan. ACM SIGSOFT Software Engineering Notes, 9(4):22–55, July 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Galbraith:1987:SUGa

[Gal87]  J. Galbraith.  Softool uses’

group meets. ACM SIGSOFT Software Engineering Notes, 12 (1):48, January 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gannon:1983:SB

[Gan83] John Gannon. SIGSOFT budget 1983. ACM SIGSOFT Software Engineering Notes, 8(2):6, April 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garman:1981:BHR

[Gar81] John R. Garman. The “BUG” heard ’round the world: discussion of the software problem which delayed the first shuttle orbital flight. ACM SIGSOFT Software Engineering Notes, 6 (5):3–10, October 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garg:1984:SOH

[Gar84] Vijay Kumar Garg. Screenoriented highlevel debugger(SHD) for PASCAL. ACM SIGSOFT Software Engineering Notes, 9 (2):117–119, April 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garg:1985:FCM

[Gar85]  Vijay Kumar Garg.  Func-

tional composition: a mechanism to compose programs. ACM SIGSOFT Software Engineering Notes, 10(3):37–39, July 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Good:1985:SGV

[GBM85] Donald I. Good, Robert S.

Boyer, and J. Strother Moore. A second generation verification environment. ACM SIGSOFT Software Engineering Notes, 10 (4):48, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guoliang:1988:KMI

[GC88] Z. Guoliang and Z. Chengxiang. The kernel of Modula-2 integrated environment. ACM SIGSOFT Software Engineering Notes, 13(4):57–60, October 1988. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Gehani:1982:SP

[Geh82] N. H. Gehani. A study in prototyping. ACM SIGSOFT Software Engineering Notes, 7(5): 71–74, December 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Geller:1985:BWC

[Gel85] Dennis P. Geller. B-ware: contradictions in a software development plan. ACM SIGSOFT Software Engineering Notes, 10 (1):48–51, January 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerhart:1981:AS

[Ger81a] Susan L. Gerhart. AFFIRM summary. ACM SIGSOFT Software Engineering Notes, 6 (3):16–24, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerhart:1981:HLT

[Ger81b] Susan L. Gerhart. High-level theories. ACM SIGSOFT Software Engineering Notes, 6(3): 24–25, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerhart:1981:RAV

[Ger81c] Susan L. Gerhart. Research avenues verification is not pursuing, but may be should be. ACM SIGSOFT Software Engineering Notes, 6(3):24, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerhart:1982:TRB

[Ger82] Susan L. Gerhart. Two recent books on programming: review of The science of programming by David Gries. Springer-Verlag 1981. and review of The craft of programming by John Reynolds. Prentice-Hall 1981. ACM SIGSOFT Software Engineering Notes, 7(2):63–64, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See comments [Gri82, Mey83].

Gerhart:1985:PTB

[Ger85] Susan L. Gerhart. Prolog technology as a basis for verification systems. ACM SIGSOFT Software Engineering Notes, 10 (4):37–40, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerhart:1989:PSF

[Ger89] S. Gerhart. Preliminary summary: FM89 assessment of formal methods for trustworthy computer systems. ACM SIGSOFT Software Engineering Notes, 14(8):152–155, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gabriel:1984:PET

[GF84] Richard P. Gabriel and Martin E. Frost. A programming environment for a timeshared system. ACM SIGSOFT Software Engineering Notes, 9 (3):185–192, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

George:1982:PCC

[GGH^{+}82] Dennis George, David Gustafson, Sallie Henry, David Hutchens, Dennis Kafura, and John Sayler. Predicting cost-of-change: from design structure metrics. ACM SIGSOFT Software Engineering Notes, 7(1):30–35, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilham:1989:TRR

[GGW89] L.-M. Gilham, A. Goldberg, and T. C. Wang. Toward reliable reactive systems. ACM SIGSOFT Software Engineering Notes, 14 (3):68–74, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gentleman:1983:HAH

[GH83] W. Morven Gentleman and Henry Hoeksma. Hardware assisted high level debugging: preliminary draft. ACM SIGSOFT Software Engineering Notes, 8 (4):140–144, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guttag:1986:ASE

[GHW86] J. V. Guttag, J. J. Horning, and J. M. Wing. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 11(1):103–110, January 1986. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilb:1981:ED

[Gil81a] Tom Gilb. Evolutionary development. ACM SIGSOFT Software Engineering Notes, 6(2):17, April 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilb:1981:SAS

[Gil81b] Tom Gilb. System attribute specification: a cornerstone of software engineering. ACM SIGSOFT Software Engineering Notes, 6(3):78–79, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilb:1984:SEU

[Gil84] Tom Gilb. Software engineering: using “design by objectives” tools (DBO). ACM SIGSOFT Software Engineering Notes, 9 (2):104–113, April 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilb:1985:EDV

[Gil85] Tom Gilb. Evolutionary delivery versus the “waterfall model”. ACM SIGSOFT Software Engineering Notes, 10(3):49–61, July 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilb:1986:ESA

[Gil86a] Tom Gilb. Estimating software attributes: some unconventional points of view. ACM SIGSOFT Software Engineering Notes, 11 (1):49–59, January 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilb:1986:SSD

[Gil86b] Tom Gilb. Software specification and design must “engineer” quality and cost iteratively. ACM SIGSOFT Software Engineering Notes, 11(1): 47–48, January 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilb:1987:DO

[Gil87] T. Gilb. Design by objectives. ACM SIGSOFT Software Engineering Notes, 12(2):42–49, April 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Giordano:1980:SVP

[Gio80] Joseph V. Giordano. Some verification problems in Pascallike languages. ACM SIGSOFT Software Engineering Notes, 5 (1):18–27, January 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gustafson:1984:SGP

[GJ84] G. G. Gustafson and T. A. Johnson. Short “good” programs. ACM SIGSOFT Software Engineering Notes, 9(2): 129, April 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glass:1980:BAS

[Gla80a] Robert L. Glass. A benefit analysis of some software reliability methodologies. ACM SIGSOFT Software Engineering Notes, 5 (2):26–33, April 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glass:1980:II

[Gla80b] Robert L. Glass. The importance of the individual. ACM SIGSOFT Software Engineering Notes, 5(3):48–50, July 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glass:1981:LRG

[Gla81] Robert L. Glass. Letter from Robert L. Glass. ACM SIG-SOFT Software Engineering Notes, 6(5):28–29, October 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gladden:1982:SLC

[Gla82a] G. R. Gladden. Stop the lifecycle, i want to get off. ACM SIGSOFT Software Engineering Notes, 7(2):35–39, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glass:1982:RMS

[Gla82b] Robert L. Glass. Recommended:

a minimum standard software toolset. ACM SIGSOFT Software Engineering Notes, 7(4):

3–13, October 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:1982:ECS

[GLR^{+}82] Hans Gill, Roland Lindvall, Olle Rosin, Erik Sandewall, Henrik Sörensen, and Ove Wigertz. Experience from computer supported prototyping for information flow in hospitals. ACM SIGSOFT Software Engineering Notes, 7(5):67–70, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goguen:1982:RPO

[GM82] Joseph Goguen and Jose Meseguer. Rapid prototyping: in the OBJ executable specification language. ACM SIGSOFT Software Engineering Notes, 7(5): 75–84, December 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1984:GIP

[GM84] David B. Garlan and Philip L. Miller. GNOME: an introductory programming environment based on a family of structure editors. ACM SIGSOFT Software Engineering Notes, 9 (3):65–72, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glasgow:1985:LSL

[GM85] Jexice I. Glasgow and Glena H. MacEwe. LUCID: a specification language for distributed systems.

ACM SIGSOFT Software Engineering Notes, 10(4):76–79, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ghezzi:1989:GWP

[GMMP89] C. Ghezzi, D. Mandrioli, S. Morasca, and M. Pezze. A general way to put time in Petri nets. ACM SIGSOFT Software Engineering Notes, 14(3):60–67, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gentleman:1989:CRS

[GMS89] W. M. Gentleman, A. MacKay, and D. A. Stewart. Commercial realtime software needs different configuration management. ACM SIGSOFT Software Engineering Notes, 14(7): 152–161, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gasser:1980:NIF

[GMW80] M. Gasser, J. K. Millen, and W. F. Wilson. A note on information flow into arrays. ACM SIGSOFT Software Engineering Notes, 5(1):28–29, January 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goguen:1981:MTS

[Gog81] J. A. Goguen. More thoughts on specification and verification. ACM SIGSOFT Software Engineering Notes, 6(3):38–41, July 1981. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Golden:1980:PFS

[Gol80] Donald Golden. A plea for friendly software. ACM SIG-SOFT Software Engineering Notes, 5(4):4–5, October 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goldberg:1987:SPT

[Gol87] Jack Goldberg. Some principles and techniques for designing safe systems. ACM SIGSOFT Software Engineering Notes, 12 (3):17–19, July 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gomaa:1983:IRP

[Gom83] Hassan Gomaa. The impact of rapid prototyping on specifying user requirements. ACM SIGSOFT Software Engineering Notes, 8(2):17–27, April 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Good:1981:TBV

[Goo81a] Donald I. Good. Toward building verified, secure systems. ACM SIGSOFT Software Engineering Notes, 6(3):4–8, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goodman:1981:ESE

[Goo81b] S. E. Goodman. An experiment in software engineering. ACM SIGSOFT Software Engineering Notes, 6(1):15, January 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goodfellow:1983:T

[Goo83] Geoffrey S. Goodfellow. Testimony. ACM SIGSOFT Software Engineering Notes, 8(5): 18–22, October 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gordon:1989:CVC

[Gor89] M. Gordon. Combining version control with automatic program building. ACM SIGSOFT Software Engineering Notes, 14 (6):25–31, October 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerth:1989:RU

[GP89] R. Gerth and A. Pnueli. Rooting UNITY. ACM SIGSOFT Software Engineering Notes, 14 (3):11–19, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gramlich:1983:DMS

[Gra83] Wayne C. Gramlich. Debugging methodology: session summary. ACM SIGSOFT Software Engineering Notes, 8(4):1–3, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Giacalone:1984:III

[GRD84] Alessandro Giacalone, Martin C.

Rinard, and Thomas W. Doeppner, Jr. IDEOSY: an ideographic and interactive program description system. ACM SIGSOFT Software Engineering Notes, 9 (3):15–20, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gregory:1984:PVM

[Gre84] S. T. Gregory. On prototypes vs. mockups. ACM SIGSOFT Software Engineering Notes, 9 (5):13, October 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Greenspan:1986:RDK

[Gre86] S. Greenspan. On the role of domain knowledge-based approaches to software development. ACM SIGSOFT Software Engineering Notes, 11(4):34–36, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gries:1980:EC

[Gri80] David Gries. Eliminating the chaff. ACM SIGSOFT Software Engineering Notes, 5(2):17–21, April 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gries:1982:LDS

[Gri82] David Gries. Letter to Dr. Susan Gerhart: comments on two book reviews. ACM SIGSOFT Software Engineering Notes, 7 (3):30–32, July 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Ger82, Mey83].

Gross:1983:DDS

[Gro83] Thomas Gross. Distributed debugging: session summary. ACM SIGSOFT Software Engineering Notes, 8(4):96–97, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grochow:1986:POS

[Gro86] Jerrold M. Grochow. A practitioner’s observations on the Software Testing Workshop. ACM SIGSOFT Software Engineering Notes, 11(5):36–37, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gumb:1982:ULS

[Gum82] Raymond D. Gumb. On the underlying logics of specification languages. ACM SIGSOFT Software Engineering Notes, 7 (4):21–23, October 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gustavsson:1989:MES

[Gus89] A. Gustavsson. Maintaining the evolution of software objects in an integrated environment. ACM SIGSOFT Software Engineering Notes, 14(7): 114–117, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Galbraith:1987:SUGb

[GY87] J. Galbraith and J. C. Young.

Softool users group. ACM SIGSOFT Software Engineering Notes, 12(3):62–63, July 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hagemann:1987:FRS

[Hag87]  M. Hagemann.  Formal re-

quirements specification of process control systems. ACM SIGSOFT Software Engineering Notes, 12(4):36–42, October 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hall:1982:DLC

[Hal82] Patrick A. V. Hall. In defence of life cycles. ACM SIGSOFT Software Engineering Notes, 7(3):23, July 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamilton:1980:CDG

[Ham80a] Dennis Hamilton. A comment on David Gries’ “Sermon”. ACM SIGSOFT Software Engineering Notes, 5(3):52, July 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamilton:1980:DLB

[Ham80b] Dennis E. Hamilton. Don’t look back, something’s gaining on us: more mixware software engineering metaphor. ACM SIGSOFT Software Engineering Notes, 5 (4):5–8, October 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamilton:1980:IWF

[Ham80c] Dennis E. Hamilton. If we’re on first, what’s on second?: a mixware software engineering discussion problem. ACM SIGSOFT Software Engineering Notes, 5(3):51–52, July 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamilton:1981:WME

[Ham81] Dennis E. Hamilton. We have met the enemy and she is us: a mixup’d search for software engineering. ACM SIGSOFT Software Engineering Notes, 6 (2):18–22, April 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:1983:DLS

[Ham83] Dick Hamlet. Debugging “level”:

step-wise debugging. ACM SIGSOFT Software Engineering Notes, 8(4):4–8, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:1989:TCT

[Ham89] R. Hamlet. Theoretical comparison of testing methods. ACM SIGSOFT Software Engineering Notes, 14(8):28–37, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harbaugh:1985:BB

[Har85] Sam Harbaugh. Beware of beware. ACM SIGSOFT Software Engineering Notes, 10(2):46, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harwood:1987:PRS

[Har87] K. Harwood. On prototyping and the role of the software engineer. ACM SIGSOFT Software Engineering Notes, 12(4):34, October 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harandi:1988:WSS

[Har88]  Mehdi T. Harandi.  Work-

shop on Software Specification and Design. ACM SIGSOFT Software Engineering Notes, 13 (1):29, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Haynes:1981:PVT

[Hay81]  Gregory A. Haynes.  Pro-

gram verification at Texas instruments. ACM SIGSOFT Software Engineering Notes, 6 (3):55–56, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hoffman:1989:MTC

[HB89] D. Hoffman and C. Brealey. Module test case generation. ACM SIGSOFT Software Engineering Notes, 14(8):97–102, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Howden:1981:CSD

[HCI^{+}81] William Howden, Paul Cohen, Al Irvine, James King, Patricia Powell, William Riddle, Leon Stucki, and Leonard Tripp. Contemporary software development environments. ACM SIGSOFT Software Engineering Notes, 6 (4):6–15, August 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hekmatpour:1987:EEP

[Hek87] S. Hekmatpour. Experience with evolutionary prototyping in a large software project. ACM SIGSOFT Software Engineering Notes, 12(1):38–41, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:1985:SDP

[Hen85]  Peter B. Henderson.  Soft-

ware development/programming environments. ACM SIGSOFT Software Engineering Notes, 10 (2):60–62, April 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:1986:SDP

[Hen86]  Peter B. Henderson.  Soft-

ware development/programming environments. ACM SIGSOFT Software Engineering Notes, 11 (2):29–31, April 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:1987:SDE

[Hen87] P. B. Henderson. Software development environments. ACM SIGSOFT Software Engineering Notes, 12(1):51–52, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson:1988:ME

[Hen88] Peter B. Henderson. Methodology and environments. ACM SIGSOFT Software Engineering Notes, 13(1):39–40, January 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hendricksen:1989:AST

[Hen89] C. S. Hendricksen. Augmented state-transition diagrams for reactive software. ACM SIGSOFT Software Engineering Notes, 14 (6):61–67, October 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hooper:1982:SBP

[HH82] James W. Hooper and Pei Hsia. Scenario-based prototyping for requirements identification. ACM SIGSOFT Software Engineering Notes, 7(5): 88–93, December 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hill:1983:RTM

[Hil83] Charles R. Hill. A real-time microprocessor debugging technique. ACM SIGSOFT Software Engineering Notes, 8(4): 145–148, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hill:1989:RCT

[Hil89] G. Hill. Radical and conservative top-down development. ACM SIGSOFT Software Engineering Notes, 14(2):32–38, April 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hitchcock:1989:PMA

[Hit89] P. Hitchcock. The process model of the aspect IPSE. ACM SIGSOFT Software Engineering Notes, 14(4):76–78, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heimbigner:1988:GTM

[HK88] Dennis Heimbigner and Steven Krane. A graph transform model for configuration management environments. ACM SIGSOFT Software Engineering Notes, 13(5):216–225, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heyman:1985:SDE

[HL85] Jerrold Heyman and William M.

Lively. Syntax-directed editing revisited. ACM SIGSOFT Software Engineering Notes, 10 (3):24–27, July 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huff:1988:PBI

[HL88] Karen E. Huff and Victor R.

Lesser. A plan-based intelligent assistant that supports the software development. ACM SIGSOFT Software Engineering Notes, 13(5):97–106, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heitmeyer:1982:UQP

[HLC82] C. Heitmeyer, C. Landwehr, and M. Cornwell. The use of quick prototypes in the secure military message systems project. ACM SIGSOFT Software Engineering Notes, 7(5):85–87, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Horgan:1984:TIL

[HM84] J. R. Horgan and D. J. Moore. Techniques for improving language-based editors. ACM SIGSOFT Software Engineering Notes, 9(3):7–14, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hennell:1980:SAC

[HMH80] M. A. Hennell, W. M. McNicol, and J. Hawkins. The static analysis of Cobol programs. ACM SIGSOFT Software Engineering Notes, 5(4):17–25, October 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

House:1989:TLS

[HN89] D. E. House and W. F. Newman. Testing large software products. ACM SIGSOFT Software Engineering Notes, 14(2):71–77, April 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hookway:1985:VAP

[Hoo85] Raymond J. Hookway. Verifying Ada programs. ACM SIGSOFT Software Engineering Notes, 10 (4):51–52, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Horning:1985:CSB

[Hor85] Jim Horning. Computing in support of battle management. ACM SIGSOFT Software Engineering Notes, 10(5):24–27, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Howden:1989:VPS

[How89] W. Howden. Validating programs without specifications. ACM SIGSOFT Software Engineering Notes, 14(8):2–9, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrold:1989:IDF

[HS89] Mary Jean Harrold and Mary Lou Soffa. Interprocedual data flow testing. ACM SIGSOFT Software Engineering Notes, 14(8): 158–167, December 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huang:1987:SSD

[Hua87] P. Huang. Survey in software development. ACM SIGSOFT Software Engineering Notes, 12 (3):51–54, July 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huff:1989:PLA

[Huf89] K. E. Huff. Probing limits to automation: towards deeper process models. ACM SIGSOFT Software Engineering Notes, 14 (4):79–81, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Humphrey:1989:SEP

[Hum89] W. S. Humphrey. The software engineering process: definition and scope. ACM SIGSOFT Software Engineering Notes, 14 (4):82–83, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Houghton:1987:CFS

[HW87] R. Houghton and D. Wallace. Characteristics and functions of software environments: an overview. ACM SIGSOFT Software Engineering Notes, 12(1): 64–84, January 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Iivari:1987:HSM

[Iiv87] J. Iivari. A hierarchical spiral model for the software process. ACM SIGSOFT Software Engineering Notes, 12(1):35–37, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Iannino:1983:CSR

[IMOL83] A. Iannino, J. D. Musa, K. Okumoto, and B. Littlewood. Criteria for software reliability model comparisons. ACM SIGSOFT Software Engineering Notes, 8 (3):12–16, July 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Iseki:1986:ADM

[IS86] Osamu Iseki and Ben Shneiderman. Applying direct manipulation concepts: direct manipulation dik operating system (DM-DOS). ACM SIGSOFT Software Engineering Notes, 11(2):22–26, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jamsa:1984:OOD

[Jam84] Kris A. Jamsa. Object oriented design vs structured design: a student’s perspective. ACM SIGSOFT Software Engineering Notes, 9(1):43–49, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Janssens:1984:SPF

[Jan84] Hans Janssens. The strange power of FORTRAN: a very short calendar printing program. ACM SIGSOFT Software Engineering Notes, 9(1):50–51, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jahnichen:1986:TAM

[JG86] S. Jähnichen and G. Goos. Towards an alternative model for software developments. ACM SIGSOFT Software Engineering Notes, 11(4):36–38, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:1983:IIS

[JK83] John D. Johnson and Gary W. Kenney. Implementation issues for a source level symbolic debugger. ACM SIGSOFT Software Engineering Notes, 8(4): 149–151, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

James:1983:RAC

[JM83] Carol L. James and Duncan E. Morrill. The real Ada, countess of Lovelace. ACM SIGSOFT Software Engineering Notes, 8 (1):30–31, January 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:1983:SEC

[Joh83] Mark Scott Johnson. Summary of evaluation and comments. ACM SIGSOFT Software Engineering Notes, 8(4):206, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:1987:SCC

[Joh87] W. L. Johnson. Some comments on coding practice. ACM SIGSOFT Software Engineering Notes, 12(2):32–35, April 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jonsson:1989:GPN

[Jon89] D. Jonsson. Graphical program notations. ACM SIGSOFT Software Engineering Notes, 14 (5):78–79, July 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jordan:1989:ECM

[Jor89] M. Jordan. Experiences in configuration management for Modula-2. ACM SIGSOFT Software Engineering Notes, 14(7): 126–128, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Joseph:1984:PSB

[Jos84] Mark K. Joseph. Programming with small blocks. ACM SIGSOFT Software Engineering Notes, 9(5):28–42, October 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jard:1983:ATS

[JvB83] C. Jard and G. v. Bochmann.

An approach to testing specifications. ACM SIGSOFT Software Engineering Notes, 8(4):53–59, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jeng:1989:SOP

[JW89] B. Jeng and E. Weyuker. Some observations on partition testing. ACM SIGSOFT Software Engineering Notes, 14(8): 38–47, December 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kenniston:1983:III

[KA83] Michael S. Kenniston and Russ Atkinson. Implementation issues II. ACM SIGSOFT Software Engineering Notes, 8(4): 117–120, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaiser:1989:RBM

[Kai89a] G. E. Kaiser. Rule-based modelling of the software development process. ACM SIGSOFT Software Engineering Notes, 14 (4):84–86, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaiser:1989:MCT

[Kai89b] G. W. Kaiser. Modeling configuration as transactions. ACM SIGSOFT Software Engineering Notes, 14(7):133–136, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Karakostas:1989:RCT

[Kar89] V. Karakostas. Requirements for CASE tools in early software reuse. ACM SIGSOFT Software Engineering Notes, 14(2):39–41, April 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Katayama:1989:HFA

[Kat89] T. Katayama. A hierarchical and functional approach to software process description. ACM SIGSOFT Software Engineering Notes, 14(4):87–92, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kellner:1989:RFS

[Kel89]  M. I. Kellner.  Representa-

tion formalisms for software process modelling. ACM SIGSOFT Software Engineering Notes, 14 (4):93–96, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kemmerer:1981:SOS

[Kem81] Richard A. Kemmerer. Status of ongoing security-relevant projects. ACM SIGSOFT Software Engineering Notes, 6(3): 42–43, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kemmerer:1985:OF

[Kem85] R. A. Kemmerer. Overview of foundations. ACM SIGSOFT Software Engineering Notes, 10 (4):75, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kempe:1989:HEP

[Kem89] M. Kempe. Hyperbook: an experiment with PCTE. ACM SIGSOFT Software Engineering Notes, 14(5):55–61, July 1989.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Keus:1982:PMR

[Keu82] Hans E. Keus. Prototyping: a more reasonable approach to system development. ACM SIGSOFT Software Engineering Notes, 7(5):94–95, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kilov:1981:LHK

[Kil81] Haim Kilov. Letter from haim kilov. ACM SIGSOFT Software Engineering Notes, 6(5):28, October 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kilov:1983:RER

[Kil83] Haim Kilov. Review of Ethnotechnical review handbook by Daniel P. Freedman and Gerald M. Weinberg. Ethnotech, Inc. 1979. ACM SIGSOFT Software Engineering Notes, 8(1): 26–29, January 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kilov:1989:BRA

[Kil89] Haim Kilov. Book review: Advanced database techniques by Daniel Martin (MIT Press 1986). ACM SIGSOFT Software Engineering Notes, 14(2):57, April 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kimball:1981:RPM

[Kim81] Edward M. Kimball. Response to Philip Miller’s letter. ACM SIGSOFT Software Engineering Notes, 6(3):65, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

King:1981:PRU

[Kin81] James C. King. Program reduction using symbolic execution. ACM SIGSOFT Software Engineering Notes, 6(1):9–14, January 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

King:1982:JDM

[Kin82] David King. Jackson design methodology. ACM SIGSOFT Software Engineering Notes, 7 (2):33–34, April 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kishimoto:1983:EDL

[Kis83] Zen Kishimoto. An experimental debugger in a limited programming environment. ACM SIGSOFT Software Engineering Notes, 8(4):63–66, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Klausner:1982:RPR

[KK82] A. Klausner and T. E. Konchan.

Rapid prototyping and requirements specification using PDS. ACM SIGSOFT Software Engineering Notes, 7(5):96–105, December 1982. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Kienle:1988:PRF

[KK88] S. C. Kienle and W. A. Keller. Problem report forms: a system for software configuration management. ACM SIGSOFT Software Engineering Notes, 13 (4):79–84, October 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kotik:1989:ASA

[KM89] G. Kotik and L. Markosian. Automating software analysis and testing using a program transformation system. ACM SIGSOFT Software Engineering Notes, 14 (8):75–84, December 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kramer:1989:CSS

[KMS89] J. Kramer, J. Magee, and M. Sloman. Configuration support for system description, construction and evolution. ACM SIGSOFT Software Engineering Notes, 14 (3):28–33, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kapur:1985:EAT

[KN85] Deepak Kapur and Paliath Narendran. An equational approach to theorem proving in first-order predicate calculus. ACM SIGSOFT Software Engineering Notes, 10(4):63–66, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koenig:1988:III

[Koe88] Shaye Koenig. (ISEF): an integrated industrial-strength software engineering framework. ACM SIGSOFT Software Engineering Notes, 13(5):45–54, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1987:DPJ

[Kok87a] P. Kokol. Dining philosophers in JSD. ACM SIGSOFT Software Engineering Notes, 12(4): 27–33, October 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1987:SPS

[Kok87b] P. Kokol. Spreadsheet programs in software engineering. ACM SIGSOFT Software Engineering Notes, 12(3):45–50, July 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1988:SEM

[Kok88] P. Kokol. Software effort metrics: how to join them. ACM SIGSOFT Software Engineering Notes, 13(2):55–57, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1989:MSD

[Kok89] P. Kokol. Metamodels for system development. ACM SIGSOFT Software Engineering Notes, 14 (5):118–123, July 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kooij:1989:IST

[Koo89a] M. Kooij. Interface specification with temporal logic. ACM SIGSOFT Software Engineering Notes, 14(3):104–110, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koomen:1989:LMD

[Koo89b] C. J. Koomen. Limits to the mechanization of the detailing step paradigm. ACM SIGSOFT Software Engineering Notes, 14 (4):97–102, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krista:1989:CAP

[KR89] R. Krista and I. Rozman. Computer aided prototyping. ACM SIGSOFT Software Engineering Notes, 14(6):68–72, October 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krasner:1987:CCS

[Kra87] Herb Krasner. CSCW’86 conference summary report. ACM SIGSOFT Software Engineering Notes, 12(2):50–51, April 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krzanik:1989:EMQ

[Krz89] L. Krzanik. Enactable models for quantitative evolutionary software processes. ACM SIGSOFT Software Engineering Notes, 14(4):103–110, June 1989.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kowalski:1982:LPS

[KS82] Robert Kowalski and Steve Smoliar. Logic for problem solving. ACM SIGSOFT Software Engineering Notes, 7(2):61–62, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kapur:1985:RTP

[KS85] Deepak Kapur and G. Sivakumar. RRL: theorem proving environment based on rewriting techniques. ACM SIGSOFT Software Engineering Notes, 10 (4):67–68, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Knoll:1989:GLB

[KS89] H. D. Knoll and W. Suk. A graphic language for business applications … user requirements specification. ACM SIGSOFT Software Engineering Notes, 14 (6):58–60, October 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kurki-Suonio:1989:ASA

[KSJ89] R. Kurki-Suonio and H.-M.

Järvinen. Action system approach to the specification and design of distributed systems. ACM SIGSOFT Software Engineering Notes, 14(3):34–40, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaplan:1989:SCS

[KsKGC89] S. M. Kaplan, s. K. Goering, and R. H. Campbell. Specifying concurrent systems with Δ-grammars. ACM SIGSOFT Software Engineering Notes, 14 (3):20–27, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kung:1984:AGH

[KT84] Antonio Kung and Michael Ian Thomas. Applying a generalized host/target concept for developing subsystems in PASCAL. ACM SIGSOFT Software Engineering Notes, 9(5):20–22, October 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Keeton-Williams:1981:PVS

[KWD81] Jim Keeton-Williams and Anne-Marie G. Discepolo. A practical verification system. ACM SIGSOFT Software Engineering Notes, 6(3):50–55, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ladden:1988:OOD

[Lad88] R. M. Ladden. Object-oriented development for Ada. ACM SIGSOFT Software Engineering Notes, 13(3):24–31, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lamb:1985:MSQ

[Lam85] Rebecca Ann Lamb. More on software quality assurance. ACM SIGSOFT Software Engineering Notes, 10(2):45, April 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Landauer:1981:AFK

[Lan81] J. R. Landauer. Applications of FDM to KVM and COS/ NFE. ACM SIGSOFT Software Engineering Notes, 6(3):43–45, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Landwehr:1985:DPV

[Lan85a] Carl E. Landwehr. Does program verification help?: how much? ACM SIGSOFT Software Engineering Notes, 10 (4):107, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Landwehr:1985:SLF

[Lan85b]  Carl E. Landwehr.  Some

lessons from formalizing a security model. ACM SIGSOFT Software Engineering Notes, 10 (4):111–112, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lawrence:1985:RSS

[Law85a] J. L. Lawrence. The RC2000:

a software success story. ACM SIGSOFT Software Engineering Notes, 10(1):31–42, January 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lawrence:1985:WSA

[Law85b] J. L. Lawrence. Why is software always late? ACM SIGSOFT Software Engineering Notes, 10 (1):19–30, January 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Linton:1983:KBD

[LBC^{+}83] Mark A. Linton, Robert Balzer, R. J. Cunningham, Cordell Green, Robert Sedlmeyer, Elliott Solloway, William Swartout, and Pamela Zave. Knowledgebased debugging: session summary. ACM SIGSOFT Software Engineering Notes, 8(4):23–24, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leblang:1984:CAS

[LC84] David B. Leblang and Robert P. Chase, Jr. Computer-aided software engineering in a distributed workstation environment. ACM SIGSOFT Software Engineering Notes, 9(3):104–112, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lie:1989:COV

[LCDK89] A. Lie, R. Conradi, T. M. Didriksen, and E.-A. Karlsson. Change oriented versioning in a software engineering database. ACM SIGSOFT Software Engineering Notes, 14(7):56–65, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:1983:IE

[Lee83] Insup Lee. Integrated environments. ACM SIGSOFT Software Engineering Notes, 8(4):60–62, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:1985:RJH

[Lee85] Kyu Y. Lee. Response to Jim Horning’s letter. ACM SIGSOFT Software Engineering Notes, 10(5):30, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lehman:1982:RPS

[Leh82] M. M. Lehman. Research proposal to study: the role of executable metric models in the programming process. ACM SIGSOFT Software Engineering Notes, 7(5):106–111, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lehman:1983:MEM

[Leh83] M. M. Lehman. More on “evolution” and “maintenance”. ACM SIGSOFT Software Engineering Notes, 8(5):39, October 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Cha83].

Lehman:1986:ADD

[Leh86] M. Lehman. Approach to a disciplined development process — the ISTAR integrated project support environment. ACM SIGSOFT Software Engineering Notes, 11(4):28–33, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lehman:1989:SRS

[Leh89] M. M. Lehman. Some reservations on software process programming. ACM SIGSOFT Software Engineering Notes, 14 (4):111–112, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leite:1989:VAC

[Lei89] J. C. S. P. Leite. Viewpoint analysis: a case study. ACM SIGSOFT Software Engineering Notes, 14(3):111–119, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lempp:1986:PMP

[Lem86] P. Lempp. A possibility to make project progress visible: the progress diagram within the EPOS—system. ACM SIGSOFT Software Engineering Notes, 11(2):32–35, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leveson:1982:SS

[Lev82] Nancy G. Leveson. Software safety. ACM SIGSOFT Software Engineering Notes, 7(2):21–24, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leveson:1987:STS

[Lev87] N. Leveson. A scary tale — Sperry avionics module-testing bites the dust? ACM SIGSOFT Software Engineering Notes, 12 (2):23–25, April 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lewerentz:1988:EPL

[Lew88] Claus Lewerentz. Extended programming in the large in a software development environment. ACM SIGSOFT Software Engineering Notes, 13(5): 173–182, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:1983:RLS

[LF83] Kyu Y. Lee and Eric C. Frankel.

Real-life software projects as software engineering laboratory exercises. ACM SIGSOFT Software Engineering Notes, 8(3): 39–43, July 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Landwehr:1986:NIW

[LGM^{+}86] Carl E. Landwehr, Susan L. Gerhart, John McLean, Donald I. Good, and Nancy Leveson. NRL Invitational Workshop on Testing and Proving: two approaches to assurance. ACM SIGSOFT Software Engineering Notes, 11 (5):63–85, October 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:1988:ODS

[LH88] Lung-Chun Liu and Ellis Horowitz. Object database support for a software project management environment. ACM SIGSOFT Software Engineering Notes, 13 (5):85–96, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lin:1989:LRF

[LH89] F. Lin and F. E. Hunt. LCDreification: a formal method for developing Prolog programs. ACM SIGSOFT Software Engineering Notes, 14(3):249–256, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Linton:1984:IRV

[Lin84] Mark A. Linton. Implementing relational views of programs. ACM SIGSOFT Software Engineering Notes, 9(3):132–140, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Livson:1987:FSD

[Liv87] B. Livson. Future software development management system concepts. ACM SIGSOFT Software Engineering Notes, 12 (3):37–41, July 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Livson:1988:PAS

[Liv88] B. U. Livson. A practical approach to software quality assurance. ACM SIGSOFT Software Engineering Notes, 13(3):45–48, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lacroix:1989:CRP

[LL89] M. Lacroix and P. Lavency. The change request process. ACM SIGSOFT Software Engineering Notes, 14(7):122–125, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

London:1989:SRC

[LM89] R. L. London and K. R. Milsted. Specifying reusable components using z: realistic sets and dictionaries. ACM SIGSOFT Software Engineering Notes, 14 (3):120–127, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Levitt:1981:RSW

[LN81] Karl N. Levitt and Peter G.

Neumann. Recent SRI work in verification. ACM SIGSOFT Software Engineering Notes, 6 (3):27–35, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loka:1989:PLE

[Lok89]  R. R. Loka.  A program-

ming language and its efficiency. ACM SIGSOFT Software Engineering Notes, 14(1):29–31, January 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loy:1985:SPC

[Loy85] Patrick H. Loy. System partitioning: a critical and often belittled development phase. ACM SIGSOFT Software Engineering Notes, 10(2):41–44, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loy:1988:TNP

[Loy88] P. Loy. Toward a new paradigm for quality assurance. ACM SIGSOFT Software Engineering Notes, 13(3):43–44, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loy:1989:ESS

[Loy89] P. H. Loy. Enlarging the scope of SQA. ACM SIGSOFT Software Engineering Notes, 14(1): 89–95, January 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:1988:ESS

[LPRS88] Peter Lee, Frank Pfenning, Gene Rollins, and William Scherlis. The ergo support system: an integrated set of tools for prototyping integrated environments. ACM SIGSOFT Software Engineering Notes, 13(5): 25–34, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:1989:OSR

[LS89] L. Y. H. Liu and R. K. Shyamasundar. An operational semantics of real time design language RT-CDL. ACM SIGSOFT Software Engineering Notes, 14 (3):75–82, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lehman:1984:ALS

[LST84] M. M. Lehman, V. Stenning, and W. M. Turski. Another look at software design methodology. ACM SIGSOFT Software Engineering Notes, 9(2):38–53, April 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lehman:1987:EPI

[LT87] M. Lehman and W. Turski.

Essential properties of IPSEs. ACM SIGSOFT Software Engineering Notes, 12(1):52–55, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lubars:1986:AHR

[Lub86a] Mitchell D. Lubars. Affording higher reliability through software reusability. ACM SIGSOFT Software Engineering Notes, 11(5):39–42, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lubars:1986:CRL

[Lub86b] Mitchell D. Lubars. Code reusability in the large versus code reusability in the small. ACM SIGSOFT Software Engineering Notes, 11(1):21–28, January 1986. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Lundholm:1989:DMB

[Lun89] P. Lundholm. Design management in base/OPEN. ACM SIGSOFT Software Engineering Notes, 14(7):38–41, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Luckham:1981:PVS

[LvH81] D. C. Luckham and F. W. von Henke. Program verification at Stanford. ACM SIGSOFT Software Engineering Notes, 6 (3):25–27, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Levitt:1985:OA

[LW85] K. N. Levitt and A. Whitehurst. Overview of applications. ACM SIGSOFT Software Engineering Notes, 10(4):95–96, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lyon:1983:MPL

[Lyo83] Gordon Lyon. Machines, programs, and languages: a simplistic introduction to computer science. ACM SIGSOFT Software Engineering Notes, 8(3):61–63, July 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

MacEwen:1982:SP

[Mac82a] Glenn H. MacEwen. Specification prototyping. ACM SIGSOFT Software Engineering Notes, 7(5):112–119, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Macrakis:1982:SP

[Mac82b] Stavros M. Macrakis. Safety and power. ACM SIGSOFT Software Engineering Notes, 7(2):25–26, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Macdonald:1985:VRS

[Mac85] Ruaridh Macdonald. Verifying a real system design: some of the problems. ACM SIGSOFT Software Engineering Notes, 10 (4):126–129, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

MacLean:1989:FPS

[Mac89]  R. MacLean.  A functional

paradigm for software development. ACM SIGSOFT Software Engineering Notes, 14(4): 113–115, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Magel:1984:PSE

[Mag84a] Kenneth Magel. Principles for software environments. ACM SIGSOFT Software Engineering Notes, 9(1):32–35, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Magel:1984:UDI

[Mag84b] Kenneth Magel. The userdefinable interface project. ACM SIGSOFT Software Engineering Notes, 9(1):36–40, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mannino:1987:PCF

[Man87] P. Mannino. A presentation and comparison of four information system development methodologies. ACM SIGSOFT Software Engineering Notes, 12(2):26–29, April 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Martin:1984:DGP

[Mar84] Alain J. Martin. On David Gries’s plateau problem. ACM SIGSOFT Software Engineering Notes, 9(1):29–30, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Martin:1988:ECS

[Mar88] R. Martin. Evaluation of current software costing tools. ACM SIGSOFT Software Engineering Notes, 13(3):49–51, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marca:1989:SCG

[Mar89] D. A. Marca. Specifying coordinators: guidelines for groupware developers. ACM SIGSOFT Software Engineering Notes, 14 (3):235–237, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mastro:1985:TDS

[Mas85] Vincent A. Mastro. Three dimensional system development. ACM SIGSOFT Software Engineering Notes, 10(5):47–59, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Minsky:1984:DSE

[MB84] N. Minsky and A. Borgida. The Darwin software-evolution environment. ACM SIGSOFT Software Engineering Notes, 9 (3):89–95, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maurer:1985:PTP

[MB85] W. D. Maurer and Tim E. Barrios. The parable of the two programmers, continued. ACM SIGSOFT Software Engineering Notes, 10(2):19–22, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maarek:1989:ULA

[MB89] Y. S. Maarek and D. M. Berry.

The use of lexical affinities in requirements extraction. ACM SIGSOFT Software Engineering Notes, 14(3):196–202, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mason:1982:ATI

[MCB82] R. E. A. Mason, T. T. Carey, and A. Benjamin. ACT/1: a tool for information systems prototyping. ACM SIGSOFT Software Engineering Notes, 7(5): 120–126, December 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McFarland:1986:BBD

[McF86] Gregory McFarland. The benefits of bottom-up design. ACM SIGSOFT Software Engineering Notes, 11(5):43–51, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marcus:1985:SSV

[MCL85a] Leo Marcus, Stephen D. Crocker, and Jaisook R. Landauer. SDVS: a system for verifying microcode correctness. ACM SIGSOFT Software Engineering Notes, 10 (4):7–14, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McLean:1985:TDP

[McL85b] John McLean. Two dogmas of program specification. ACM SIGSOFT Software Engineering Notes, 10(4):85–87, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mack-Crane:1989:CMS

[MCP89] B. Mack-Crane and A. Pal. Conflict management in a source version management system. ACM SIGSOFT Software Engineering Notes, 14(7):149–151, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meyer:1983:SMW

[Mey83] Bertrand Meyer. Some mistakes are worse than others: an open letter to professor David Gries. ACM SIGSOFT Software Engineering Notes, 8(5):43–48, October 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Ger82, Gri82].

Masiero:1988:JOO

[MG88] P. Masiero and F. S. R. Germano. JSD as an object-oriented design method. ACM SIGSOFT Software Engineering Notes, 13 (3):22–23, July 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miriyala:1989:AAS

[MH89] K. Miriyala and M. T. Harandi.

Analogical approach to specification derivation. ACM SIGSOFT Software Engineering Notes, 14 (3):203–210, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mikelsons:1983:IPE

[Mik83] Martin Mikelsons. Interactive program execution in lispedit. ACM SIGSOFT Software Engineering Notes, 8(4):71–80, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Millen:1981:RVW

[Mil81] J. K. Millen. Recent verification work at MITRE. ACM SIGSOFT Software Engineering Notes, 6(3):55, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mili:1982:DSB

[Mil82] Ali Mili. Designing software without backtracking: the part of dream and the part of reality. ACM SIGSOFT Software Engineering Notes, 7(1):77, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miller:1989:SCM

[Mil89] T. C. Miller. A schema for configuration management. ACM SIGSOFT Software Engineering Notes, 14(7):26–29, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Minsky:1986:WSW

[Min86]  Naftaly H. Minsky.  What

should we do about the evolution of software: a position paper. ACM SIGSOFT Software Engineering Notes, 11(4):68–70, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miranda:1989:CTT

[Mir89a] E. L. Miranda. Control transformations through Petri nets. ACM SIGSOFT Software Engineering Notes, 14(2):45–48, April 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miranda:1989:LEL

[Mir89b] E. L. Miranda. Looking for the event list. ACM SIGSOFT Software Engineering Notes, 14 (5):80–82, July 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mittermeir:1982:HLF

[Mit82] Roland T. Mittermeir. HIBOL:

a language for fast prototyping in data processing environments. ACM SIGSOFT Software Engineering Notes, 7(5):133–140, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miya:1982:RTB

[Miy82] E. N. Miya. Reviews of two books by E. N. Yourdon: review of “Classics in software engineering by E. N. Yourdon. Yourdon Press 1979. and Writings of the Revolution by E. N. Yourdon. Yourdon Press 1982. ACM SIGSOFT Software Engineering Notes, 7(3):24–29, July 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miya:1985:SE

[Miy85] Eugene Miya. On “Software engineering”. ACM SIGSOFT Software Engineering Notes, 10 (1):12, January 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miyamoto:1987:PTG

[Miy87] I. Miyamoto. A prototyping tool for graphical software engineering tools. ACM SIGSOFT Software Engineering Notes, 12 (4):45–51, October 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McCracken:1982:LCC

[MJ82] Daniel D. McCracken and Michael A. Jackson. Life cycle concept considered harmful. ACM SIGSOFT Software Engineering Notes, 7(2):29–32, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mahler:1988:ITE

[ML88] Axel Mahler and Andreas Lampen. An integrated toolset for engineering software configurations. ACM SIGSOFT Software Engineering Notes, 13(5): 191–200, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Melhart:1989:ACR

[MLJ89] B. E. Melhart, N. G. Leveson, and M. S. Jaffe. Analysis capabilities for requirements specified in statecharts. ACM SIGSOFT Software Engineering Notes, 14 (3):100–103, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McCoyd:1982:SSG

[MM82] Gerard C. McCoyd and John R. Mitchell. System sketching: the generation of rapid prototypes for transaction based systems. ACM SIGSOFT Software Engineering Notes, 7(5):127–132, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McHugh:1985:AVU

[MN85] John McHugh and Karl Nyberg.

Ada verification using existing tools. ACM SIGSOFT Software Engineering Notes, 10(4): 104–106, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mok:1986:NSP

[Mok86] Aloysius K. Mok. The von Neumann straitjacket: the process construct. ACM SIGSOFT Software Engineering Notes, 11 (4):60–65, August 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Moore:1981:EWV

[Moo81] J. Moore. Experience in writing VCG systems. ACM SIGSOFT Software Engineering Notes, 6 (3):35–38, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Moore:1985:SES

[Moo85] Martin J. Moore. Software engineering and SDI. ACM SIGSOFT Software Engineering Notes, 10(5):28, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Munck:1988:ODS

[MOPT88] Robert Munck, Patricia Oberndorf, Erhard Ploedereder, and Richard Thall. An overview of DOD-STD-1838A (proposed) the common APSE interface set: revision. ACM SIGSOFT Software Engineering Notes, 13(5): 235–247, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Moriconi:1989:PAS

[Mor89]  M. Moriconi.  A practical

approach to semantic configuration management. ACM SIGSOFT Software Engineering Notes, 14(8):103–113, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mano:1981:IMD

[MOT81] Yoshihisa Mano, Kazuhito Omaki, and Koji Torii. An intelligent multi-display terminal system towards: a better programming environment. ACM SIGSOFT Software Engineering Notes, 6(2):8–14, April 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Minsky:1988:SDE

[MR88] Naftaly H. Minsky and David Rozenshtein. A software development environment for lawgoverned systems. ACM SIG-SOFT Software Engineering Notes, 13(5):65–75, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McDonald:1986:SMA

[MRY86] Catherine W. McDonald, William Riddle, and Christine Youngblut. Stars methodology area summary: volume II: preliminary views on the software life cycle and methodology selection. ACM SIGSOFT Software Engineering Notes, 11(2):58–85, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miller:1989:IAC

[MSK89] D. B. Miller, R. G. Stockton, and C. W. Krueger. An inverted approach to configuration management. ACM SIGSOFT Software Engineering Notes, 14 (7):1–4, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Melliar-Smith:1985:EHS

[MSR85] Michael Melliar-Smith and John Rushby. The enhanced HDM system for specification and verification. ACM SIGSOFT Software Engineering Notes, 10(4): 41–43, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Melliar-Smith:1982:PS

[MSS82] P. M. Melliar-Smith and Richard L.

Schwartz. The proof of SIFT. ACM SIGSOFT Software Engineering Notes, 7(1):2–5, January 1982. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Mullerburg:1983:RDW

[Mül83] Monika A. F. Müllerburg. The role of debugging within software engineering environments. ACM SIGSOFT Software Engineering Notes, 8(4):81–90, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Murray:1988:PRD

[Mur88a] J. Murray. Problem reporting and development tracking using VM/SP. ACM SIGSOFT Software Engineering Notes, 13 (4):74–78, October 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Murray:1988:SCU

[Mur88b] John Murray. Source control using VM/SP and CMS. ACM SIGSOFT Software Engineering Notes, 13(2):51–54, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Musa:1983:SSE

[Mus83] John D. Musa. Stimulating software engineering progress: a report of the software engineering planning group. ACM SIGSOFT Software Engineering Notes, 8 (2):29–54, April 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Musser:1985:AHS

[Mus85] David R. Musser. AIDS to hierarchical specification structuring and reusing theorems in AFFIRM-85. ACM SIGSOFT Software Engineering Notes, 10 (4):2–4, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Narayanaswamy:1989:TBR

[Nar89] K. Narayanaswamy. A textbased representation for program variants. ACM SIGSOFT Software Engineering Notes, 14 (7):30–33, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1988:RPb

[NC88] P. Neumann and Other Contributors. Risks to the public. ACM SIGSOFT Software Engineering Notes, 13(4):3–20, October 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ncube:1988:ATM

[Ncu88] Cathy C. Ncube. Ada tasking model and Prolog predication. ACM SIGSOFT Software Engineering Notes, 13(2):48–50, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nelson:1985:WRS

[Nel85] David A. Nelson. Will the real SE metaphors please stand up!: (or, I never metaphor I didn’t like!). ACM SIGSOFT Software Engineering Notes, 10 (5):31–32, October 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1981:RIV

[Neu81] Peter G. Neumann. Retrospective introduction to VERkshop II. ACM SIGSOFT Software Engineering Notes, 6(3):2, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1982:PIC

[Neu82] Peter G. Neumann. Psychosocial implications of computer software development and use: Zen and the art of computing. ACM SIGSOFT Software Engineering Notes, 7(2):3–12, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1984:RHD

[Neu84] Peter G. Neumann. Review of The hacker’s dictionary: a guide to the world of computer wizards by Guy L. Steele, Donald R. Woods, Raphael A. Finkel, Mark R. Crispin, Richard M. Stallman, and Geoffrey S. Goodfellow. Harper and Row Publishers 1983. ACM SIGSOFT Software Engineering Notes, 9 (1):12–15, January 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1985:RSA

[Neu85] Peter G. Neumann. Review of The Sachertorte algorithm and other antidotes to computer anxiety by John Shore. Viking 1985. ACM SIGSOFT Software Engineering Notes, 10(2):23–27, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1986:RPCa

[Neu86a] Peter G. Neumann. Risks to the public in computer systems. ACM SIGSOFT Software Engineering Notes, 11(1):2–14, January 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1986:RPCb

[Neu86b] Peter G. Neumann. Risks to the public in computer systems. ACM SIGSOFT Software Engineering Notes, 11(2):1, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1986:RPCc

[Neu86c] Peter G. Neumann. Risks to the public in computer systems. ACM SIGSOFT Software Engineering Notes, 11(2):2–17, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1986:RPCd

[Neu86d] Peter G. Neumann. Risks to the public in computer systems. ACM SIGSOFT Software Engineering Notes, 11(5):3–28, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1987:RPa

[Neu87a] P. G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 12 (3):2–17, July 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1987:RPb

[Neu87b] P. G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 12 (4):2–18, October 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1987:RPC

[Neu87c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 12 (1):3–22, January 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1988:RPa

[Neu88a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 13 (1):3–16, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1988:RPR

[Neu88b] Peter G. Neumann. Risks to the public and related systems. ACM SIGSOFT Software Engineering Notes, 13(2):5–17, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1989:EIP

[Neu89] Peter G. Neumann. Errata: Interchanged pages in the previous issue: SEN 14-1: Page 31 and page 46. ACM SIGSOFT Software Engineering Notes, 14 (2):78, April 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Newbery:1989:ECM

[New89] F. J. Newbery. Edge concentration: a method for clustering directed graphs. ACM SIGSOFT Software Engineering Notes, 14 (7):76–85, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Newsted:1981:IPS

[NLY81] Peter R. Newsted, Wing-Keen Leong, and Joanna Yeung. The impact of programming styles on debugging efficiency. ACM SIGSOFT Software Engineering Notes, 6(5):14–18, October 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Niskier:1989:LTP

[NMS89] C. Niskier, T. Maibaum, and D. Schwabe. A look through PRISMA: towards pluralistic knowledge-based environments for software specification acquisition. ACM SIGSOFT Software Engineering Notes, 14(3): 128–136, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:1988:RBS

[Not88] David Notkin. The relationship between software development environments and the software process. ACM SIGSOFT Software Engineering Notes, 13(5): 107–109, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:1989:ASP

[Not89]  D. Notkin.  Applying soft-

ware process models to the full lifecycle is premature. ACM SIGSOFT Software Engineering Notes, 14(4):116–117, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nord:1988:EAS

[NP88] Robert L. Nord and Frank Pfenning. The Ergo Attribute System. ACM SIGSOFT Software Engineering Notes, 13(5): 110–120, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Newcastle:1986:ASE

[NTB^{+}86] U. Newcastle, U. Texas, R. S. Boyer, J. S. Moore, W. A. Hunt, R. M. Cohen, U. Toronto, and R. C. Holt. Abstracts in software engineering — reports. ACM SIGSOFT Software Engineering Notes, 11(5):92–94, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:1981:SER

[OCG^{+}81] Leon Osterweil, Lori Clarke, Donald Good, Raymond Houghton, Thomas Love, Patricia Santoni, Daniel Teichroew, and Anthony Wasserman. Software environment research: the next five years. ACM SIGSOFT Software Engineering Notes, 6(4):15–25, August 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Offutt:1989:CEF

[Off89] A. Offutt. The coupling effect: fact or fiction. ACM SIGSOFT Software Engineering Notes, 14(8):131–140, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Olenfeldt:1985:VSP

[Ole85] Lars Olenfeldt. A very short paper on the OBJECT/EVENT-ANALYSIS. ACM SIGSOFT Software Engineering Notes, 10 (1):52–57, January 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Omar:1989:STP

[OM89] A. A. Omar and F. A. Mohammed. Structural testing of programs. ACM SIGSOFT Software Engineering Notes, 14 (2):62–70, April 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ottenstein:1983:HLD

[OO83] Karl J. Ottenstein and Linda M.

Ottenstein. High-level debugging assistance via optimizing compiler technology. ACM SIGSOFT Software Engineering Notes, 8(4):152–154, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ottenstein:1984:PDG

[OO84] Karl J. Ottenstein and Linda M. Ottenstein. The program dependence graph in a software development environment. ACM SIGSOFT Software Engineering Notes, 9(3):177–184, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ohki:1989:PPP

[OO89a] A. Ohki and K. Ochimizu. Process programming with Prolog. ACM SIGSOFT Software Engineering Notes, 14(4):118–121, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Olender:1989:CSS

[OO89b] K. Olender and L. Osterweil. Cesar: a static sequencing constraint analyzer. ACM SIG-SOFT Software Engineering Notes, 14(8):66–74, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:1989:ASE

[Ost89] L. Osterweil. Automated support for the enactment of rigorously described software processes. ACM SIGSOFT Software Engineering Notes, 14(4): 122–125, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Owre:1985:STP

[Owr85] Sam Owre. The sytek theorem prover. ACM SIGSOFT Software Engineering Notes, 10 (4):69, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pakstas:1989:ALC

[Pak89]  A. Pakstas.  Aladdin/lamp:

Configuration management tools for distributed computer control systems. ACM SIGSOFT Software Engineering Notes, 14(7): 141–144, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1981:LGU

[Par81] D. L. Parnas. Lastig gepreek uit nuenen. ACM SIGSOFT Software Engineering Notes, 6 (3):64–65, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1983:HPW

[Par83a] Girish Parikh. How to pick a winner: a look at software methodologies. ACM SIGSOFT Software Engineering Notes, 8 (2):11–12, April 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parker:1983:T

[Par83b] Donn B. Parker. Testimony.

ACM SIGSOFT Software Engineering Notes, 8(5):10–17, October 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1984:WSM

[Par84] Girish Parikh. What is software maintenance really?: what is in a name? ACM SIGSOFT Software Engineering Notes, 9 (2):114–116, April 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1985:DWS

[Par85a]  Girish Parikh.  Discovering

the world of software maintenance: selected readings. ACM SIGSOFT Software Engineering Notes, 10(5):79–88, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1985:SMN

[Par85b] Girish Parikh. Software maintenance news. ACM SIGSOFT Software Engineering Notes, 10 (2):58–59, April 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1985:SMP

[Par85c] Girish Parikh. Software maintenance: penny wise, program foolish. ACM SIGSOFT Software Engineering Notes, 10(5): 89–98, October 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parker:1985:TMI

[Par85d] Jeff Parker. Towards more intelligent programming environments. ACM SIGSOFT Software Engineering Notes, 10(3):28–32, July 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1985:SAS

[Par85e] David Lorge Parnas. Software aspects of strategic defense systems. ACM SIGSOFT Software Engineering Notes, 10(5): 15–23, October 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1986:EWSb

[Par86a] Girish Parikh. Exploring the world of software maintenance III: research and development — future directions in software maintenance. ACM SIGSOFT Software Engineering Notes, 11 (5):86–91, October 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1986:EWSa

[Par86b] Girish Parikh. Exploring the world of software maintenance: what is software maintenance? ACM SIGSOFT Software Engineering Notes, 11(2):49–52, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1986:NWA

[Par86c] Girish Parikh. The new world of “AI-SM”: (artificial intelligence in software maintenance). ACM SIGSOFT Software Engineering Notes, 11(2):53–56, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1987:EE

[Par87a] G. Parikh. “experter” and “expertation”. ACM SIGSOFT Software Engineering Notes, 12 (3):61, July 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parikh:1987:SMN

[Par87b] G. Parikh. Software maintenance notes. ACM SIGSOFT Software Engineering Notes, 12 (2):51–53, April 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Patton:1983:PNP

[Pat83] Bob Patton. Prototyping: a nomenclature problem. ACM SIGSOFT Software Engineering Notes, 8(2):14–16, April 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pazirandeh:1986:TPM

[Paz86] Reza Pazirandeh. Test plan methodology. ACM SIGSOFT Software Engineering Notes, 11 (1):64–72, January 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Podgurski:1989:IPD

[PC89] A. Podgurski and L. Clarke. The implications of program dependencies for software testing, debugging, and maintenance. ACM SIGSOFT Software Engineering Notes, 14(8):168–178, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pechura:1980:PEI

[Pec80] Michael Pechura. Programming as engineering: insights and comparisons. ACM SIGSOFT Software Engineering Notes, 5 (2):22–25, April 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perlman:1980:THL

[Per80] Gary Perlman. Two high-level skills for programming: a comment on R. L. Glass’ The importance of the individual. ACM SIGSOFT Software Engineering Notes, 5(4):9, October 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perlman:1986:CQT

[Per86a] G. Perlman. Coding quality and tools in programming methods. ACM SIGSOFT Software Engineering Notes, 11(3):44–50, July 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perry:1986:TES

[Per86b] Dewayne E. Perry. Tools for evolving software: a position paper. ACM SIGSOFT Software Engineering Notes, 11(4):70–71, August 1986. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Perry:1989:LPI

[Per89a] D. Perry. The logic of propagation in the Inscape Environment. ACM SIGSOFT Software Engineering Notes, 14(8): 114–121, December 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perry:1989:PSP

[Per89b] D. E. Perry. Problems of scale and process models. ACM SIGSOFT Software Engineering Notes, 14(4):126–128, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Peters:1989:TES

[Pet89] L. Peters. Timing extensions to structured analysis for real time systems. ACM SIGSOFT Software Engineering Notes, 14 (3):83–90, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ploedereder:1989:DMC

[PF89] E. Ploedereder and A. Fergany. The data model of the configuration management assistant (CMA). ACM SIGSOFT Software Engineering Notes, 14(7): 5–14, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Phillips:1989:SCA

[Phi89] R. W. Phillips. State change architecture: a protocol for executable process models. ACM SIGSOFT Software Engineering Notes, 14(4):129–132, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pirie:1986:LS

[Pir86] I. W. Pirie. The LIFESPAN system. ACM SIGSOFT Software Engineering Notes, 11(2):27–28, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pixon:1985:EFE

[Pix85] Mark R. Pixon. Enhancing FDM for the expression of concurrency requirements. ACM SIGSOFT Software Engineering Notes, 10 (4):92–94, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Powell:1983:DMD

[PL83] Michael L. Powell and Mark A.

Linton. A database model of debugging. ACM SIGSOFT Software Engineering Notes, 8 (4):67–70, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Penteado:1985:SCJ

[PMG85] R. D. Penteado, P. C. Masiero, and F. S. Germano. Structure charts for Jackson system development. ACM SIGSOFT Software Engineering Notes, 10 (3):40–46, July 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pohl:1986:SSA

[Poh86] Ira Pohl. SDI software: AI is not the answer. ACM SIGSOFT Software Engineering Notes, 11 (2):18–19, April 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Polak:1980:RKF

[Pol80] Wolfgang Polak. Response to K. Foster’s letter. ACM SIGSOFT Software Engineering Notes, 5 (1):37–38, January 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Fos80].

Penedo:1988:OMI

[PPT88] Maria H. Penedo, Erhard Ploedereder, and Ian Thomas. Object Management Issues for Software Engineering Environments workshop report. ACM SIGSOFT Software Engineering Notes, 13 (5):226–234, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pratt:1987:CTS

[Pra87] S. J. Pratt. Catastrophe theory and software engineering. ACM SIGSOFT Software Engineering Notes, 12(3):55–57, July 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prentice:1981:ASD

[Pre81] Dan Prentice. An analysis of software development environments. ACM SIGSOFT Software Engineering Notes, 6(5):

19–27, October 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Proctor:1985:RAP

[Pro85] Norman Proctor. The restricted access processor: an example of formal verification. ACM SIGSOFT Software Engineering Notes, 10(4):116–118, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pett:1989:FFD

[PRS89] M. M. Pett, M. D. Rice, and S. B. Seidman. A formal framework for the design of development environments. ACM SIGSOFT Software Engineering Notes, 14(3):284–286, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Putnam:1985:SMS

[Put85a]  Dan Putnam.  Separating

methodology and specification constructs. ACM SIGSOFT Software Engineering Notes, 10 (4):90–91, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Putnam:1985:VDV

[Put85b] Dan Putnam. The VERUS^{T M} design verification system. ACM SIGSOFT Software Engineering Notes, 10(4):26–27, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rajaraman:1982:CSD

[Raj82] M. K. Rajaraman. A characterization of software design tools. ACM SIGSOFT Software Engineering Notes, 7(4):14–17, October 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramsey:1989:DFV

[Ram89] N. Ramsey. Developing formally verified Ada programs. ACM SIGSOFT Software Engineering Notes, 14(3):257–265, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richardson:1989:ITT

[RAO89] D. Richardson, S. Aha, and L. Osterweil. Integrating testing techniques through process programming. ACM SIGSOFT Software Engineering Notes, 14 (8):219–228, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ratcliffe:1987:HWS

[Rat87] M. Ratcliffe. Hereford Workshop on Software Reuse. ACM SIGSOFT Software Engineering Notes, 12(1):42–47, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reps:1989:III

[RB89] T. Reps and T. Bricker. Illustrating interference in interfering versions of programs. ACM SIGSOFT Software Engineering Notes, 14(7):46–55, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riddle:1982:PTM

[RBB^{+}82] Bill Riddle, Vic Basili, Les Belady, Carl Davis, Susan Gerhart, Bill Howden, John Musa, and Leon Stucki. Proposals for tool and methodology evaluation experiments: ACM SIG-SOFT sponsored software engineering symposium on tool and methodology evaluation. ACM SIGSOFT Software Engineering Notes, 7(1):6–10, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Roman:1989:ISD

[RC89] G.-C. Roman and K. C. Cox.

Implementing a shared dataspace language on a messagebased multiprocessor. ACM SIGSOFT Software Engineering Notes, 14(3):41–48, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reddy:1988:TDP

[Red88] Uday S. Reddy. Transformational derivation of programs using the focus system. ACM SIGSOFT Software Engineering Notes, 13(5):163–172, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reed:1987:RES

[Ree87] K. Reed. Review of the environments symposium. ACM SIGSOFT Software Engineering Notes, 12(1):56–62, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reid:1982:NM

[Rei82] Brain Reid. Netmail manhandled. ACM SIGSOFT Software Engineering Notes, 7(2):27–28, April 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reiss:1984:GPD

[Rei84] Steven P. Reiss. Graphical program development with PECAN program development systems. ACM SIGSOFT Software Engineering Notes, 9(3):30–41, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reifer:1985:RPP

[Rei85] Donald J. Reifer. Review of Principles of productive software management by Michael W. Evans, Pamela Piazza, and James B. Dolkes. John Wiley and Sons 1983. ACM SIGSOFT Software Engineering Notes, 10 (1):58, January 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reid:1986:LUB

[Rei86]  Brian Reid.  Lessons from

the UNIX breakins at Stanford.

ACM SIGSOFT Software Engineering Notes, 11(5):29–35, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reichenberger:1989:OVM

[Rei89] C. Reichenberger. Orthogonal version management. ACM SIGSOFT Software Engineering Notes, 14(7):137–140, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reynolds:1985:PTM

[Rey85] Robert G. Reynolds. PARTIAL:

a tool to monitor the stepwise refinement of Ada programs. ACM SIGSOFT Software Engineering Notes, 10(3):76–94, July 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Razouk:1989:RTI

[RG89] R. Razouk and M. Gorlick. Realtime interval logic for reasoning about executions of real-time programs. ACM SIGSOFT Software Engineering Notes, 14(8): 10–19, December 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rickert:1985:SMS

[Ric85] Neil W. Rickert. A system for managing shared files. ACM SIGSOFT Software Engineering Notes, 10(2):35–40, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richter:1986:ASA

[Ric86a] Charles A. Richter. An assessment of structured analysis and structured design. ACM SIGSOFT Software Engineering Notes, 11(4):41–45, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rickert:1986:PWR

[Ric86b] Neil W. Rickert. Preconditions for widespread reuse of code. ACM SIGSOFT Software Engineering Notes, 11(2):21, April 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riddle:1984:MNE

[Rid84a] William E. Riddle. The magic number eighteen plus or minus three: a study of software technology maturation. ACM SIGSOFT Software Engineering Notes, 9(2):21–37, April 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riddle:1984:RSP

[Rid84b] William E. Riddle. Report on the Software Process Workshop. ACM SIGSOFT Software Engineering Notes, 9(2):13–20, April 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rine:1987:CEO

[Rin87] D. C. Rine. A common error in the object structuring of object-oriented design methods. ACM SIGSOFT Software Engineering Notes, 12(4):42–44, October 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ritter:1987:HNA

[Rit87] A. Ritter. Highlights of the new approach for structuring complex systems. ACM SIGSOFT Software Engineering Notes, 12 (2):38–42, April 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosenblum:1989:TCT

[RL89] D. Rosenblum and D. Luckham.

Testing the correctness of tasking supervisors with TSL specifications. ACM SIGSOFT Software Engineering Notes, 14(8): 187–196, December 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rueher:1989:CSP

[RLL89] M. Rueher, D. Ladret, and B. Legeard. Capturing software processes through the generated objects. ACM SIGSOFT Software Engineering Notes, 14 (4):148–152, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robinson:1984:RDB

[Rob84] David Robinson. A response to Dave Brown’s “Why we did so badly in the design phase”. ACM SIGSOFT Software Engineering Notes, 9(5):13, October 1984. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic). See [Bro84].

Roberts:1989:DAP

[Rob89a] C. Roberts. Describing and acting process models with PML. ACM SIGSOFT Software Engineering Notes, 14(4):136–141, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robillard:1989:EGN

[Rob89b] P.-N. Robillard. On the evolution of graphical notation for program design. ACM SIGSOFT Software Engineering Notes, 14(1):84–88, January 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robinson:1989:IMS

[Rob89c] W. N. Robinson. Integrating multiple specifications using domain goals. ACM SIGSOFT Software Engineering Notes, 14 (3):219–226, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rombach:1989:SFS

[Rom89]  D. Rombach.  A specifica-

tion framework for software processes: formal specification and derivation of information base requirements. ACM SIGSOFT Software Engineering Notes, 14 (4):142–147, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosen:1981:VNC

[Ros81] Eric C. Rosen. Vulnerabilities of network control protocols: an example. ACM SIGSOFT Software Engineering Notes, 6 (1):6–8, January 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richardson:1989:ASB

[ROT89] D. Richardson, O. O’Malley, and C. Tittle. Approaches to specification-based testing. ACM SIGSOFT Software Engineering Notes, 14(8):86–96, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramanathan:1985:TBI

[RR85] Jayashree Ramanathan and John M. Rose. TRIAD: beyond isolated systems for development-in-the-large and programming-in-the-small. ACM SIGSOFT Software Engineering Notes, 10(5):62–72, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Redwine:1989:SRP

[RR89] S. S. Redwine, Jr. and W. E.

Riddle. Software reuse processes. ACM SIGSOFT Software Engineering Notes, 14(4):133–135, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramanathan:1982:UAS

[RS82] J. Ramanathan and C. J.

Shubra. Use of annotated schemes for developing prototype programs. ACM SIGSOFT Software Engineering Notes, 7 (5):141–149, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rodden:1988:IAI

[RSS88] Thomas Rodden, Pete Sawyer, and Ian Sommerville. Interacting with an active, integrated environment. ACM SIGSOFT Software Engineering Notes, 13 (5):76–84, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reps:1984:SG

[RT84] Thomas Reps and Tim Teitelbaum. The synthesizer generator. ACM SIGSOFT Software Engineering Notes, 9(3):42–48, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramanathan:1985:URD

[RVAR85] Jayashree Ramanathan, Robert Vermilyer, Thorbjorn Anderson, and John M. Rose. The use of a relational data base in method driven environments. ACM SIGSOFT Software Engineering Notes, 10(5):73–78, October 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rich:1982:DUS

[RW82] Charles Rich and Richard C. Waters. The disciplined use of simplifying assumptions. ACM SIGSOFT Software Engineering Notes, 7(5):150–154, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riddle:1986:SEW

[RW86] William E. Riddle and Lloyd G.

Williams. Software Environments Workshop report. ACM SIGSOFT Software Engineering Notes, 11(1):73–102, January 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reubenstein:1989:RAI

[RW89] H. B. Reubenstein and R. C. Waters. The requirements apprentice: an initial scenario. ACM SIGSOFT Software Engineering Notes, 14(3):211–218, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rowland:1989:ECT

[RZ89] J. Rowland and Y. Zuyuan. Experimental comparison of three system test strategies preliminary report. ACM SIGSOFT Software Engineering Notes, 14 (8):141–149, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rzepka:1983:RSE

[Rze83] William E. Rzepka. RADC SREM evaluation program: a status report. ACM SIGSOFT Software Engineering Notes, 8 (1):20–22, January 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Spinrad:1985:WWL

[SA85] Mark Spinrad and Curt Abraham. The Wild-West Lifecycle (WILI). ACM SIGSOFT Software Engineering Notes, 10 (3):47–48, July 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smestad:1988:IPI

[SA88] Tore Smestad and Kristian Andersen. Integrated projection illustrating. ACM SIGSOFT Software Engineering Notes, 13 (1):52–56, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saaltink:1985:RS

[Saa85] Mark Saaltink. Relational semantics. ACM SIGSOFT Software Engineering Notes, 10(4): 82–83, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Salenieks:1989:SDA

[Sal89] P. Salenieks. Software development: advice by analogy. ACM SIGSOFT Software Engineering Notes, 14(1):56–60, January 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schwanke:1989:DVC

[SAP89] R. W. Schwanke, R. Z. Altucher, and M. A. Platoff. Discovering, visualizing, and controlling software structure. ACM SIGSOFT Software Engineering Notes, 14 (3):147–154, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Summersgill:1989:HFP

[SB89] R. Summersgill and D. P.

Browne. Human factors: its place in system development methods. ACM SIGSOFT Software Engineering Notes, 14(3): 227–234, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Standish:1981:ADS

[SBB^{+}81] Thomas Standish, Robert Balzer, David Barstow, Meera Blattner, Martha Branstad, John Buxton, Adele Goldberg, Robert Morris, and Stephen Squires. Advanced development support systems. ACM SIGSOFT Software Engineering Notes, 6(4):25–35, August 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smith:1985:GVE

[SC85] Michael K. Smith and Richard M.

Cohen. Gypsy verification environment: status. ACM SIGSOFT Software Engineering Notes, 10(4):5–6, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scacchi:1989:MSE

[Sca89a] W. Scacchi. Modelling software evolution: a knowledge-based approach. ACM SIGSOFT Software Engineering Notes, 14(4): 153–155, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scacchi:1989:USF

[Sca89b] W. Scacchi. The USC system factory project. ACM SIGSOFT Software Engineering Notes, 14 (1):61–82, January 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smoliar:1983:SAD

[SCC^{+}83] Stephen W. Smoliar, William W. Cotterman, J. Daniel Couger, Norman L. Enger, and Frederick Harold. System analysis and design: a foundation for the 1980’s. ACM SIGSOFT Software Engineering Notes, 8 (3):65–69, July 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schorer:1980:MPS

[Sch80a] Pete Schorer. More from Pete shorer. ACM SIGSOFT Software Engineering Notes, 5(2): 38–39, April 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schorer:1980:SMP

[Sch80b] Peter Schorer. Still more from Peter Schorer. ACM SIGSOFT Software Engineering Notes, 5 (3):52, July 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scheffer:1981:CQM

[Sch81a] Paul A. Scheffer. Comparisons of quality metrics. ACM SIGSOFT Software Engineering Notes, 6 (3):76–77, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schilling:1981:PMD

[Sch81b] Peter E. Schilling. A program model for design. ACM SIGSOFT Software Engineering Notes, 6(2):15–16, April 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schneidewind:1981:MST

[Sch81c] Norman F. Schneidewind. Message from the secretary/treasurer. ACM SIGSOFT Software Engineering Notes, 6(1):4, January 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schultz:1985:SSL

[Sch85a] David J. Schultz. Standard for the software life-cycle process. ACM SIGSOFT Software Engineering Notes, 10(3):62, July 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schwartz:1985:SIC

[Sch85b] Devid P. Schwartz. Summary of IEEE CSM WG meeting #5, 23–25 Jan 85. ACM SIGSOFT Software Engineering Notes, 10 (2):51–57, April 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schilling:1986:SMA

[Sch86a] Pete Schilling. Software methodology: ASTM. ACM SIGSOFT Software Engineering Notes, 11 (5):52, October 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schindler:1986:AM

[Sch86b] Max J. Schindler. Algorithms and models. ACM SIGSOFT Software Engineering Notes, 11 (1):28, January 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schroeder:1989:IVC

[Sch89] U. Schroeder. Incremental variant control. ACM SIGSOFT Software Engineering Notes, 14 (7):145–148, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Soupos:1987:SDE

[SGCZ87] R. Soupos, S. Goutas, D. Christodoulakis, and C. Zaroliagis. A software development environment database. ACM SIGSOFT Software Engineering Notes, 12(1): 63, January 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaw:1982:SCJ

[Sha82] Chris Shaw. Structure charts for Jackson structured programming. ACM SIGSOFT Software Engineering Notes, 7(1): 78–80, January 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaw:1989:LSS

[Sha89] M. Shaw. Larger scale systems require higher-level abstractions. ACM SIGSOFT Software Engineering Notes, 14(3):143–146, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaw:1984:IFT

[SHD84] R. C. Shaw, P. N. Hudson, and N. W. Davis. Introduction of a formal technique into a software development environment: early observations. ACM SIGSOFT Software Engineering Notes, 9 (2):54–79, April 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shyamasundar:1989:RRT

[SHG89] R. K. Shyamasundar, J. Hooman, and R. Gerth. Reasoning of realtime distributed programming languages. ACM SIGSOFT Software Engineering Notes, 14 (3):91–99, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Simmonds:1989:CMP

[Sim89] I. Simmonds. Configuration management in the PACT software engineering environment. ACM SIGSOFT Software Engineering Notes, 14(7):118–121, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Strand:1982:PSS

[SJ82] Eugene M. Strand and Warren T. Jones. Prototyping and small scale software projects. ACM SIGSOFT Software Engineering Notes, 7(5):169–170, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Skelton:1986:MMT

[Ske86] Stan Skelton. Measurements of migratability and transportability. ACM SIGSOFT Software Engineering Notes, 11(1): 29–34, January 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schneider:1988:WDP

[SL88] F. B. Schneider and L. Lamport. On E. W. Dijkstra’s position paper on “fairness”. ACM SIGSOFT Software Engineering Notes, 13(3):18–19, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij88, CM88b].

Schwan:1986:GVP

[SM86] K. Schwan and J. Matthews. Graphical views of parallel programs. ACM SIGSOFT Software Engineering Notes, 11(3):51–64, July 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shlaer:1989:OOA

[SM89] S. Shlaer and S. J. Mellor. An object-oriented approach to domain analysis. ACM SIGSOFT Software Engineering Notes, 14 (5):66–77, July 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smith:1985:PPP

[Smi85] Brian T. Smith. Position paper to be presented at “Verification workshop III (verkshop III)” to be held in Watsonville, CA. Feb. 18–21, 1985. ACM SIGSOFT Software Engineering Notes, 10 (4):61–62, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smoliar:1981:TBN

[Smo81]  Steve Smoliar.  Two books

named Software reliability: review of Software reliability by Thomas A. Thayer, Myron Lipow, Eldred C. Nelson. North-Holland 1978. and Software reliability by Hermann Kopetz. Springer-Verlag 1980. ACM SIGSOFT Software Engineering Notes, 6(3):66–67, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smoliar:1982:AES

[Smo82]  Stephen W. Smoliar.  Ap-

proaches to executable specifications. ACM SIGSOFT Software Engineering Notes, 7(5): 155–159, December 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smoliar:1983:RLM

[Smo83] Stephen W. Smoliar. Review of Literary machines by Ted Nelson. 1983. ACM SIGSOFT Software Engineering Notes, 8 (5):34–36, October 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smoliar:1984:RPP

[Smo84] Stephen W. Smoliar. Review of Programming in PROLOG by W. F. Clocksin and C. S. Mellish. Springer-Verlag 1981. ACM SIGSOFT Software Engineering Notes, 9(1):16–18, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Snodgrass:1984:MSD

[Sno84] Richard Snodgrass. Monitoring in a software development environment: a relational approach. ACM SIGSOFT Software Engineering Notes, 9(3):124–131, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Snodgrass:1986:AIT

[Sno86] Richard Snodgrass. Announcement of the IDL toolkit. ACM SIGSOFT Software Engineering Notes, 11(1):111, January 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1980:ASE

[Sof80] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 5 (1):39–44, January 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1981:ASE

[Sof81] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 6 (3):80–83, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1982:ASEa

[Sof82a] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 7 (1):82–83, January 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1982:ASEb

[Sof82b] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 7 (3):33–36, July 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1983:ASEa

[Sof83a] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 8(1): 99–102, January 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1983:ASEb

[Sof83b] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 8 (3):75–79, July 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1983:ASEc

[Sof83c] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 8 (5):49–54, October 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1984:ASEa

[Sof84a] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 9 (2):145–146, April 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1984:ASEb

[Sof84b] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 9 (4):62–66, July 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1984:ASEc

[Sof84c] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 9 (5):90–94, October 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1985:ASEa

[Sof85a] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 10 (1):74–81, January 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1985:ASEb

[Sof85b] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 10 (2):118–125, April 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1985:ASEc

[Sof85c] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 10 (3):109–117, July 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1985:PAA

[Sof85d] Software Engineering Notes Staff. Proceedings of the ACM Ada TEC “Future Ada environment workshop”. ACM SIGSOFT Software Engineering Notes, 10(2):77–117, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1985:QR

[Sof85e] Software Engineering Notes Staff. Questionnaire responses. ACM SIGSOFT Software Engineering Notes, 10(2):12–15, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Soloway:1984:CBM

[Sol84] Elliot Soloway. A cognitivelybased methodology for designing languages/environments/ methodologies. ACM SIGSOFT Software Engineering Notes, 9 (3):193–196, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Soloway:1986:WES

[Sol86] Eliot Soloway. Workshop on Empirical Studies of Programmers. ACM SIGSOFT Software Engineering Notes, 11(5):38, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sommerville:1987:REC

[Som87] I. Sommerville. Review of 1986 environments conference. ACM SIGSOFT Software Engineering Notes, 12(2):54–55, April 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schwanke:1989:CRF

[SP89] R. W. Schwanke and M. A. Platoff. Cross references are features. ACM SIGSOFT Software Engineering Notes, 14(7): 86–95, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Spicer:1983:SAS

[Spi83] Joseph C. Spicer. A spiral approach to software engineering project management education. ACM SIGSOFT Software Engineering Notes, 8(3):30–38, July 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smith:1981:VEP

[SSDG81] Michael K. Smith, Ann E.

Siebert, Benedetto L. DiVito, and Donald I. Good. A verified encrypted packet interface. ACM SIGSOFT Software Engineering Notes, 6(3):13–16, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Seidner:1983:IDR

[ST83] Rich Seidner and Nick Tindall.

Interactive debug requirements. ACM SIGSOFT Software Engineering Notes, 8(4):9–22, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Standish:1984:APA

[ST84] Thomas A. Standish and Richard N.

Taylor. Arcturus: a prototype advanced Ada(R) programming environment. ACM SIGSOFT Software Engineering Notes, 9 (3):57–64, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stavely:1982:MED

[Sta82]  Allan M. Stavely.  Models

as executable designs. ACM SIGSOFT Software Engineering Notes, 7(5):167–168, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stinson:1989:VSA

[Sti89a] W. Stinson. Views of SDE: automation of engineering and engineering of automation. ACM SIGSOFT Software Engineering Notes, 14(6):32–41, October 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stinson:1989:VSD

[Sti89b] W. Stinson. Views of software development environments. ACM SIGSOFT Software Engineering Notes, 14(5):108–117, July 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sedlmeyer:1983:KBF

[STJ83] Robert L. Sedlmeyer, William B. Thompson, and Paul E. Johnson. Knowledge-based fault localization in debugging: preliminary draft. ACM SIGSOFT Software Engineering Notes, 8 (4):25–31, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stokes:1985:SFM

[Sto85] R. Stokes. Some formal method activities in UK industry. ACM SIGSOFT Software Engineering Notes, 10(4):129–130, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Strelich:1988:SLC

[Str88] Tom Strelich. The software life cycle support environment (SLCSE): a computer based framework for developing software systems. ACM SIGSOFT Software Engineering Notes, 13 (5):35–44, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sutton:1989:AMS

[Sut89] W. L. Sutton. Advanced models of the software process. ACM SIGSOFT Software Engineering Notes, 14(4):156–158, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Squires:1982:RPWa

[SZB82a] Stephen L. Squires, Marvin Zelkowitz, and Martha Branstad. Rapid Prototyping Workshop: an overview. ACM SIGSOFT Software Engineering Notes, 7(3):14–15, July 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Squires:1982:RPWb

[SZB82b] Stephen L. Squires, Marvin Zelkowitz, and Martha Branstad. Rapid Prototyping Workshop: overview. ACM SIGSOFT Software Engineering Notes, 7(5):2, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tai:1989:WDB

[Tai89] K. C. Tai. What to do beyond branch testing. ACM SIGSOFT Software Engineering Notes, 14 (2):58–61, April 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tanik:1980:CPC

[Tan80a] Murat M. Tanik. A comparison of program complexity prediction models. ACM SIGSOFT Software Engineering Notes, 5 (4):10–16, October 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See more references [Tan81].

Tanik:1980:SDM

[Tan80b] Murat M. Tanik. Software development monitoring graphs. ACM SIGSOFT Software Engineering Notes, 5(2):34–37, April 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tanik:1981:ARC

[Tan81] Murat M. Tanik. Additional references for: “A comparison of program complexity prediction models”. ACM SIGSOFT Software Engineering Notes, 6 (1):14, January 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Tan80a].

Tardo:1981:VIO

[Tar81]  Joseph J. Tardo.  Verifica-

tion: inside/outside views. ACM SIGSOFT Software Engineering Notes, 6(3):57–58, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tardy:1988:HTA

[Tar88] J. E. Tardy. Host/target approach to embedded system development is becoming obsolete. ACM SIGSOFT Software Engineering Notes, 13(4):45–51, October 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tavares:1982:MA

[Tav82] C. D. Tavares. More anecdotage.

ACM SIGSOFT Software Engineering Notes, 7(1):81–82, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:1985:VPP

[Tay85] Tad Taylor. Verkshop position paper. ACM SIGSOFT Software Engineering Notes, 10(4): 113–116, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:1988:FAE

[TBC^{+}88] Richard N. Taylor, Frank C.

Belz, Lori A. Clarke, Leon Osterweil, Richard W. Selby, Jack C. Wileden, Alexander L. Wolf, and Michael Young. Foundations for the Arcadia environment architecture. ACM SIGSOFT Software Engineering Notes, 13(5): 1–13, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tate:1985:RPS

[TD85] Graham Tate and Thomas W. G.

Docker. A rapid prototyping system based on data flow principles. ACM SIGSOFT Software Engineering Notes, 10(2):28–34, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thomas:1988:PEP

[Tho88] I. Thomas. PCTE Esprit and Pact. ACM SIGSOFT Software Engineering Notes, 13(4): 52–56, October 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thomas:1989:VCM

[Tho89a] I. Thomas. Version and configuration management on a software engineering database. ACM SIGSOFT Software Engineering Notes, 14(7):23–25, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thompson:1989:FPE

[Tho89b] S. Thompson. Functional programming: executable specifications and program transformations. ACM SIGSOFT Software Engineering Notes, 14(3): 287–290, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tichy:1986:SS

[Tic86]  Walter Tichy.  SDI soft-

ware. ACM SIGSOFT Software Engineering Notes, 11(1): 1:1–1:??, January 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terwilliger:1989:EFS

[TMO89] R. B. Terwilliger, M. J. Maybee, and L. J. Osterweil. An example of formal specification as an aid to design and development. ACM SIGSOFT Software Engineering Notes, 14(3):266–272, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tse:1986:APN

[TP86] T. H. Tse and L. Pong. An application of Petri nets in structured analysis. ACM SIGSOFT Software Engineering Notes, 11 (5):53–56, October 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1988:SRM

[Tra88] Will Tracz. Software reuse myths. ACM SIGSOFT Software Engineering Notes, 13(1): 17–21, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tripp:1988:GNP

[Tri88] L. L. Tripp. Graphical notations for program design. ACM SIGSOFT Software Engineering Notes, 13(4):39–44, October 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:1982:ITR

[TS82] Tamara Taylor and Thomas A.

Standish. Initial thoughts on rapid prototyping techniques.

ACM SIGSOFT Software Engineering Notes, 7(5):160–166, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thomson:1989:CMU

[TS89] R. Thomson and I. Sommerville. Configuration management using SySL. ACM SIGSOFT Software Engineering Notes, 14(7): 106–109, November 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tischler:1983:SAP

[TSP83] Ron Tischler, Robin Schaufler, and Charlotte Payne. Static analysis of programs as an aid to debugging. ACM SIGSOFT Software Engineering Notes, 8 (4):155–158, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tully:1989:SPM

[Tul89] C. J. Tully. Software process models and programs: observations on their nature and context. ACM SIGSOFT Software Engineering Notes, 14(4): 159–162, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ungar:1984:GSN

[Ung84] David Ungar. Generation scavenging: a non-disruptive high performance storage reclamation algorithm. ACM SIGSOFT Software Engineering Notes, 9 (3):157–167, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

ofLowell:1987:GKS

[Uni87] University of Lowell. Graphical Kernel System. ACM SIGSOFT Software Engineering Notes, 12(3):64, July 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vandenBosch:1982:ESD

[vdBEF^{+}82] Fred van den Bosch, John R. Ellis, Peter Freeman, Len Johnson, Carma L. McClure, Dick Robinson, Walt Scacchi, Ben Scheff, Arndt von Staa, and Leonard L. Tripp. Evaluation of software development life cycle: methodology implementation. ACM SIGSOFT Software Engineering Notes, 7(1):45–60, January 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Veloso:1983:MLS

[VF83] P. A. S. Veloso and A. L. Furtado. Multi-level specifications based on traces. ACM SIGSOFT Software Engineering Notes, 8 (1):17–19, January 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vonHenke:1985:RHS

[vH85] Friedrich W. von Henke. Reasoning with Hoare sentences. ACM SIGSOFT Software Engineering Notes, 10(4):84, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vonIssendorff:1989:ART

[vI89] H. von Issendorff. On the algebraic representation and treatment of processing structures. ACM SIGSOFT Software Engineering Notes, 14(3):52–55, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vitter:1984:UNF

[Vit84] Jeffrey Scott Vitter. US&R: a new framework for redoing (extended abstract). ACM SIGSOFT Software Engineering Notes, 9(3):168–176, May 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vanVleck:1989:TQA

[vV89] T. van Vleck. Three questions about each bug you find. ACM SIGSOFT Software Engineering Notes, 14(5):62–63, July 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Williams:1985:PVS

[WA85] James Williams and Charles Applebaum. The Practical Verification System Project. ACM SIGSOFT Software Engineering Notes, 10(4):44–47, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Walker:1981:IC

[Wal81] Steve Walker. Introductory comments. ACM SIGSOFT Software Engineering Notes, 6(3):3, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Walter:1983:DUD

[Wal83] Carol K. Walter. DELTA: the universal debugger for CP-6. ACM SIGSOFT Software Engineering Notes, 8(4):203–205, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ward:1985:UEN

[War85] Jean Renard Ward. UNIX as an environment for non-UNIX software development: a case history. ACM SIGSOFT Software Engineering Notes, 10 (3):95–108, July 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Waguespack:1987:SCA

[WB87] L. Waguespack, Jr. and S. Badiani. Software complexity assessment: annotated bibliography. ACM SIGSOFT Software Engineering Notes, 12(4):52–71, October 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Walpole:1988:UMC

[WBMN88] J. Walpole, G. S. Blair, J. Malik, and J. R. Nicol. A unifying model for consistent distributed software development environments. ACM SIGSOFT Software Engineering Notes, 13(5): 183–190, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wileden:1989:FPP

[WCW89] J. C. Wileden, L. A. Clarke, and Alexander L. Wolf. Facilitating process prototyping by controlling the impact of change. ACM SIGSOFT Software Engineering Notes, 14(4):163–166, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber:1983:IIS

[Web83a] Janice Cynthia Weber. Implementation issues: session summary. ACM SIGSOFT Software Engineering Notes, 8(4): 114–116, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber:1983:IDC

[Web83b] Jonice Cynthia Weber. Interactive debugging of concurrent programs. ACM SIGSOFT Software Engineering Notes, 8(4): 112–113, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber:1984:DDS

[Web84] Herbert Weber. The distributed development system: a monolithic software development environment. ACM SIGSOFT Software Engineering Notes, 9 (5):43–72, October 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weber:1989:FSF

[Web89] D. G. Weber. Formal specification of fault-tolerance and its relation to computer security. ACM SIGSOFT Software Engineering Notes, 14(3):273–277, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wegner:1980:ALE

[Weg80] Peter Wegner. The Ada language and environment. ACM SIGSOFT Software Engineering Notes, 5(2):8–14, April 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wegner:1982:RCI

[Weg82] Peter Wegner. Reflections on capital-intensive software technology. ACM SIGSOFT Software Engineering Notes, 7(4): 24–33, October 1982. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weissman:1981:VT

[Wei81] Clark Weissman. On verification targets. ACM SIGSOFT Software Engineering Notes, 6 (3):61–63, July 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weiser:1982:SMR

[Wei82] Mark Weiser. Scale models and rapid prototyping. ACM SIGSOFT Software Engineering Notes, 7(5):181–185, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weiss:1985:SDI

[Wei85] David Weiss. Strategic defense initiative software feasibility (panel): panel discussion at the 8th ICSE. ACM SIGSOFT Software Engineering Notes, 10 (5):29, October 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weiss:1989:CTD

[Wei89] S. N. Weiss. Comparing test data adequacy criteria. ACM SIGSOFT Software Engineering Notes, 14(6):42–49, October 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wertz:1983:ILP

[Wer83] Harald Wertz. An integrated Lisp programming environment. ACM SIGSOFT Software Engineering Notes, 8(4):91–95, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Werth:1987:SSE

[Wer87] L. H. Werth. Survey of software engineering education. ACM SIGSOFT Software Engineering Notes, 12(4):19–26, October 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Westphal:1988:KES

[Wes88] Christopher Westphal. Knowledge engineering for software design. ACM SIGSOFT Software Engineering Notes, 13(1):

57–58, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Westfechtek:1989:RCI

[Wes89] B. Westfechtek. Revision control in an integrated software development environment. ACM SIGSOFT Software Engineering Notes, 14(7):96–105, November 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wallace:1987:PSV

[WF87] D. R. Wallace and R. U. Fujii. Planning for software verification and validation. ACM SIGSOFT Software Engineering Notes, 12(2):37, April 1987. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Whelan:1981:ICS

[Whe81] J. J. Whelan. 5th international conference on software engineering. ACM SIGSOFT Software Engineering Notes, 6 (5):32–35, October 1981. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Whitehurst:1985:NID

[Whi85] R. Alan Whitehurst. The need for an integrated design, implementation, verification and testing methodology. ACM SIGSOFT Software Engineering Notes, 10(4):97–100, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

White:1986:TEC

[Whi86] Stephanie White. Two embedded computer system requirements models issues for investigation. ACM SIGSOFT Software Engineering Notes, 11(4): 56–59, August 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

William:1985:CVT

[Wil85] Jim William. Components of verification technology. ACM SIGSOFT Software Engineering Notes, 10(4):49–50, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wileden:1986:IMS

[Wil86] J. Wileden. This is IT: a metamodel of the software process. ACM SIGSOFT Software Engineering Notes, 11(4):9–11, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Williams:1988:PVS

[Wil88] B. Williams. Punchcard voting systems. ACM SIGSOFT Software Engineering Notes, 13 (3):21, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Williams:1989:BAS

[Wil89] L. G. Williams. A behavioral approach to software process modelling. ACM SIGSOFT Software Engineering Notes, 14 (4):167–170, June 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:1985:BFB

[Win85] Jeannette M. Wing. Beyond functional behavior: combining methods to specify different classes of properties of large systems. ACM SIGSOFT Software Engineering Notes, 10(4): 102–104, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Winters:1986:RCS

[Win86] Edward Winters. Requirements checklist for a system development workstation. ACM SIGSOFT Software Engineering Notes, 11(5):57–62, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Winkler:1988:WSV

[Win88] J. F. H. Winkler. Workshop on Software Version and Configuration Control. ACM SIGSOFT Software Engineering Notes, 13 (4):61–73, October 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Williams:1989:SDI

[WMAJ89] Gerald B. Williams, Chunka Mui, Vairam Alagappan, and Bruce B. Johnson. Software design issues: a very large information systems perspective. ACM SIGSOFT Software Engineering Notes, 14(3):238–240, May 1989.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wilson:1985:PHT

[WO85] William Wilson and Sam Owre. Programmable heuristics for theorem provers. ACM SIGSOFT Software Engineering Notes, 10 (4):70–71, August 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Woodcock:1988:F

[Woo88] J. C. P. Woodcock. Formalisms. ACM SIGSOFT Software Engineering Notes, 13(1):30–32, January 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Woodcock:1989:PZS

[Woo89] J. C. P. Woodcock. Properties of z specifications. ACM SIGSOFT Software Engineering Notes, 14(5):43–54, July 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wasserman:1989:OOS

[WPM89] A. I. Wasserman, Peter Pircher, and R. J. Muller. An object-oriented structured design method for code generation. ACM SIGSOFT Software Engineering Notes, 14(1):32–55, January 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wasserman:1982:RPI

[WS82] Anthony I. Wasserman and David T. Shewmake. Rapid prototyping of interactive information systems. ACM SIGSOFT Software Engineering Notes, 7 (5):171–180, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wileden:1988:PEP

[WWFT88] Jack C. Wileden, Alexander L.

Wolf, Charles D. Fisher, and Peri L. Tarr. Pgraphite: an experiment in persistent typed object management. ACM SIGSOFT Software Engineering Notes, 13(5):130–142, November 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Watanabe:1989:RCO

[WY89] T. Watanabe and A. Yonezawa.

Reflective computation in objectoriented concurrent systems and its applications. ACM SIGSOFT Software Engineering Notes, 14 (3):56–58, May 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yack:1982:CZA

[Yac82] Stan Yack. A comment on “Zen and the art of computing”. ACM SIGSOFT Software Engineering Notes, 7(4):39, October 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yelowitz:1981:FAP

[Yel81]  L. Yelowitz.  FACC activi-

ties/positions in methodology, verification, and trusted systems. ACM SIGSOFT Software Engineering Notes, 6(3):46–50, July 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yonezaki:1986:LEN

[YES86] Naoki Yonezaki, Hajime Enomoto, and Motoshi Saeki. Logical environment for natural language approach. ACM SIGSOFT Software Engineering Notes, 11 (4):45–48, August 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yiannakos:1984:QAQ

[Yia84] Andrew E. Yiannakos. Quality assurance and quality control in the software development process. ACM SIGSOFT Software Engineering Notes, 9 (2):130–132, April 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yates:1989:REI

[YM89] D. Yates and N. Malevris. Reducing the effects of infeasible paths in branch testing. ACM SIGSOFT Software Engineering Notes, 14(8):48–54, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Young:1988:SUG

[You88] Jenny C. Young. SofTool users group. ACM SIGSOFT Software Engineering Notes, 13(1): 68–70, January 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Young:1989:CCM

[You89a] J. Young. Change/configuration management gathering. ACM SIGSOFT Software Engineering Notes, 14(5):105–107, July 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Young:1989:VCM

[You89b] W. Young. Verified compilation in micro-Gypsy. ACM SIGSOFT Software Engineering Notes, 14(8):20–26, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Young:1989:ICA

[YTFB89] M. Young, R. Taylor, K. Forester, and D. Brodbeck. Integrated concurrency analysis in a software development environment. ACM SIGSOFT Software Engineering Notes, 14(8):200–209, December 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yue:1989:RFO

[Yue89]  K. Yue.  Representing first

order logic-based specifications in Petri-net-like graphs. ACM SIGSOFT Software Engineering Notes, 14(3):291–293, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zelkowitz:1981:HLL

[ZAA^{+}81] Marvin Zelkowitz, W. Richards Adrion, Alfred Aho, Daniel Bobrow, Thomas Cheatham, John Cherniavsky, Susan Gerhart, Gordon Lyon, John Nestor, and Terry Straeter. High level language programming environments. ACM SIGSOFT Software Engineering Notes, 6(4):36–51, August 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zahniser:1983:LAS

[Zah83]  Richard A. Zahniser.  Lev-

els of abstraction in the system life cycle. ACM SIGSOFT Software Engineering Notes, 8 (1):6–12, January 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zahniser:1988:PTD

[Zah88a] Richard A. Zahniser. The perils of top-down design. ACM SIGSOFT Software Engineering Notes, 13(2):22–24, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zahniser:1988:CSV

[Zah88b] Rivhard A. Zahniser. Computer science vs. software engineering. ACM SIGSOFT Software Engineering Notes, 13 (2):21, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zave:1984:C

[Zav84a] Pamela Zave. Correction. ACM SIGSOFT Software Engineering Notes, 9(5):10, October 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Zav84b].

Zave:1984:OPP

[Zav84b] Pamela Zave. An overview of the PAISLey project — 1984. ACM SIGSOFT Software Engineering Notes, 9(4):12–19, July 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See correction [Zav84a].

Zave:1984:RSM

[Zav84c] Pamela Zave. Report to SIG-SOFT membership fiscal year 1983 actual income and expenses. ACM SIGSOFT Software Engineering Notes, 9(1): 8, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zave:1986:FPS

[Zav86a] Pamela Zave. Found poetry of software engineering. ACM SIGSOFT Software Engineering Notes, 11(2):20, April 1986. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zave:1986:LPM

[Zav86b] Pamela Zave. Let’s put more emphasis on perspective methods. ACM SIGSOFT Software Engineering Notes, 11(4):52–53, August 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zave:1988:A

[Zav88] Pamela Zave. Assessment. ACM SIGSOFT Software Engineering Notes, 13(1):40–43, January 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zellweger:1983:IHL

[Zel83] Polle T. Zellweger. An interactive high-level debugger for control-flow optimized programs. ACM SIGSOFT Software Engineering Notes, 8(4): 159–172, August 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zelkowits:1984:SCE

[Zel84a] Marvin V. Zelkowits. A small contribution to editing with a syntax directed editor. ACM SIGSOFT Software Engineering Notes, 9(3):1–6, May 1984. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zelkowitz:1984:TPD

[Zel84b] Marvin V. Zelkowitz. A taxonomy of prototype designs. ACM SIGSOFT Software Engineering Notes, 9(5):11–12, October 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zimmer:1988:WND

[Zim88] J. A. Zimmer. We need a discipline of software composing. ACM SIGSOFT Software Engineering Notes, 13(2):39–40, April 1988. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Zucconi:1989:SCT

[Zuc89a] L. Zucconi. Selecting a CASE tool. ACM SIGSOFT Software Engineering Notes, 14(2):42–44, April 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zucconi:1989:TEC

[Zuc89b] L. Zucconi. Techniques and experiences capturing requirements for several real-time applications. ACM SIGSOFT Software Engineering Notes, 14(6): 51–55, October 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).